### v1

# A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization

Abstract
Gaussian graphical models (GGMs) are widely used for recovering the conditional independence structure among random variables. Recently, several key advances have been made to exploit an additional set of variables for better estimating the GGMs of the variables of interest. For example, in co-expression quantitative trait locus (eQTL) studies, both the mean expression level of genes as well as their pairwise conditional independence structure may be adjusted by genetic variants local to those genes. Existing methods to estimate covariate-adjusted GGMs either allow only the mean to depend on covariates or suffer from poor scaling assumptions due to the inherent non-convexity of simultaneously estimating the mean and precision matrix. In this paper, we propose a convex formulation that jointly estimates the covariate-adjusted mean and precision matrix by utilizing the natural parametrization of the multivariate Gaussian likelihood. This convexity yields theoretically better performance as the sparsity and dimension of the covariates grow large relative to the number of samples. We verify our theoretical results with numerical simulations and perform a reanalysis of an eQTL study of glioblastoma multiforme (GBM), an aggressive form of brain cancer.
1 Introduction
Graphical models are used to represent the distribution of a random vector $\bm{X}=(X_{1},\dotsc,X_{p})$ by relating its conditional independence structure to a graph. This correspondence is particularly salient when $\bm{X}$ is Gaussian. Letting $\boldsymbol{\mathbf{\Omega}}=\boldsymbol{\mathbf{\Sigma}}^{-1}$ be the precision matrix of a Gaussian random vector $\bm{X}$, for $i\neq j$, the components $X_{i}$ and $X_{j}$ are conditionally independent given $\bm{X}_{\{1,\dotsc,p\}\setminus{\{i,j\}}}$ if and only if $\Omega_{ij}=0$ [Lauritzen, 1996]. Models that estimate the conditional independence structure by imposing sparsity on $\boldsymbol{\mathbf{\Omega}}$ are known as Gaussian graphical models (GGMs).
There has been considerable research on GGMs and their estimation in the high-dimensional setting [Yuan and Lin, 2007, Friedman et al., 2007, Meinshausen and Bühlmann, 2006]. GGMs find major applications in genomics where the network structure of genes is of interest. Such a setting is typically high-dimensional such that the number of genes exceeds the number of samples [Schäfer and Strimmer, 2005]. Of particular interest in genomics are co-expression quantitative trait locus (eQTL) studies. In these studies, gene expression data are analyzed alongside information about external genetic markers that are known to confound the expression levels of genes. In order to use GGMs for eQTL studies, one must allow the GGM framework to incorporate covariate information.
Let $\bm{X}\in\mathbb{R}^{p}$ be the vector of responses and $\bm{U}\in\mathbb{R}^{q}$ the associated vector of covariates. We consider a general model
$\bm{X}\>|\>\bm{U}=\boldsymbol{\mathbf{u}}\sim N(\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}}),\boldsymbol{\mathbf{\Omega}}^{-1}(\boldsymbol{\mathbf{u}}))$ (1)
and a specification given by
$\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Omega}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}.$ (2)
Note that in this specification, the error covariance matrix does not depend on covariates. Multivariate regression models specify (2) with the goal to improve estimation of $\boldsymbol{\mathbf{\Gamma}}$ by also estimating $\boldsymbol{\mathbf{B}}_{0}$; see for example Yuan et al. [2007] and references therein for this well-studied setting. By contrast, the focus of the covariate-adjusted GGM framework is to estimate the sparsity pattern of the precision matrix while accounting for covariate information through $\boldsymbol{\mathbf{B}}_{0}$. Several methods exist in this framework that employ sparse penalization to the negative log-likelihood of (1) [Rothman et al., 2010, Yin and Li, 2011, Cai et al., 2013, Chen et al., 2016, Wang, 2015, Chen et al., 2018]. These models also are termed “conditional Gaussian graphical models” in Yin and Li [2011].
Estimating (1) with (2) is challenging because the negative log-likelihood of (1) is not jointly convex in $\boldsymbol{\mathbf{\Gamma}}$ and $\boldsymbol{\mathbf{B}}_{0}$. Therefore, covariate-adjusted GGM methods use either alternating [Rothman et al., 2010, Yin and Li, 2011, Chen et al., 2018] or two-stage [Cai et al., 2013, Chen et al., 2016] estimation algorithms. Despite these challenges, minimax-optimal error rates in jointly estimating (2) have been established [Chen et al., 2018, Lv et al., 2022].
Although there has been progress in allowing heterogeneous means in GGMs, there is little work on allowing the graph structure of $\boldsymbol{\mathbf{\Omega}}$ to also depend on covariates; indeed the aforementioned methods all assume that $\boldsymbol{\mathbf{B}}_{0}$ is fixed across subjects. Yet subject-specific graph structures are plausible in applications such as eQTL analyses. For example, gene C may mediate the co-expression of genes A and B, but only in the presence of a single-nucleotide polymorphism (SNP) local to gene C. In other words, the expression of genes A and B may be conditionally independent given the rest of the network unless a genetic variant is present near gene C [Fehrmann et al., 2011, Kolberg et al., 2020, Rockman and Kruglyak, 2006]. To capture this additional structure, Zhang and Li [2022] extends (2) to allow both the mean and the network structure of the responses to be adjusted by covariates. Their model specifies (1) by taking
$\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Omega}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}$ (3)
where $\boldsymbol{\mathbf{B}}_{h},h=0,1\dotsc,q$ are sparse, symmetric matrices. Like previous works in covariate-adjusted GGMs, the joint estimation of the parameters $(\boldsymbol{\mathbf{\Gamma}},\boldsymbol{\mathbf{B}}_{0},\boldsymbol{\mathbf{B}}_{1},\dotsc,\boldsymbol{\mathbf{B}}_{h})$ in (3) involves a non-jointly-convex objective. Therefore Zhang and Li [2022] use a two-stage estimation procedure like in preceding methods.
In this work, we contribute a jointly convex formulation of (1) such that both the mean and the graph structure may depend on covariates. Like Zhang and Li [2022], our method uses nodewise regression to estimate the covariate-adjusted graph structure [Meinshausen and Bühlmann, 2006]. However, we base our formulation upon a natural parametrization of the multivariate Gaussian likelihood such that each nodewise regression is a convex optimization problem.
1.1 Previous work
Our method is inspired by the penalized likelihood-based method of Zhang and Li [2022]. Bayesian approaches to subject-level precision matrix adjustment have been considered before. In Niu et al. [2024], a Bayesian method is developed where a discrete set of precision matrices $\boldsymbol{\mathbf{\Omega}}_{1},\dotsc,\boldsymbol{\mathbf{\Omega}}_{K}$ is determined for a partition of $n$ samples into $K$ sets, with the partition depending on external covariates. However, this approach does not immediately allow for a covariate-adjusted mean, nor is it easy to interpret how each covariate affects the overall graph structure through the $\boldsymbol{\mathbf{\Omega}}_{i}$. Wang et al. [2022] develops a Bayesian approach to solve similar nodewise regression problems as Zhang and Li [2022], utilizing sparsity-inducing priors to estimate the sparse component matrices. Their method also does not allow a covariate-adjusted mean, nor does it identify group structures.
Convex formulations of covariate-adjusted GGMs have been proposed before. Wang [2015] remarks on the non-convex nature of existing covariate-adjusted GGM methods and suggests a convex formulation of (1) with (2). However, their method achieves convexity by relying on an initial estimate of $\boldsymbol{\mathbf{\Gamma}}$, thereby placing it in the category of two-stage methods. A related line of work is the conditional Gaussian random field, which is also sometimes called the conditional Gaussian graphical model [Sohn and Kim, 2012, Yuan and Zhang, 2014]. This class of methods investigates estimating a sparse $\boldsymbol{\mathbf{B}}_{0}$ in (1) and (2) by penalizing a convex negative log-likelihood. This parametrization is made possible by the additional assumption that $(\bm{X},\bm{U})$ is jointly Gaussian. In Zhu [2020], the matrix $\boldsymbol{\mathbf{\Gamma}}$ in (2) is written as $\boldsymbol{\mathbf{\Gamma}}=\tilde{\boldsymbol{\mathbf{\Gamma}}}\boldsymbol{\mathbf{\Omega}}$ leading to a negative log-likelihood of $\boldsymbol{\mathbf{X}}$ that is jointly convex in $\tilde{\boldsymbol{\mathbf{\Gamma}}}$ and $\boldsymbol{\mathbf{\Omega}}$. However, their method does not immediately allow for covariate-adjusted network structures.
Our proposed method differs from these works as it allows both the mean and the precision matrix to depend on external covariates while also offering the advantage of solving a convex optimization problem.
1.2 Outline
We recall the Gaussian graphical regression framework of Zhang and Li [2022] and motivate the natural parametrization in Section 2. In Section 3 we provide the model specification and algorithm to induce a sparse-group structure in the graph. Our theoretical results are discussed in Section 4, namely that the natural parametrization allows for better theoretical scaling of $p$ and $q$ relative to $n$ under the same assumptions as in Zhang and Li [2022]. This is demonstrated through extensive simulations in Section 5. Finally we apply our method to perform a reanalysis of data from glioblastoma microforme (GBM) tissue samples in Section 6.
2 Parametrizations for Covariate-adjusted Graphical Models
Let $\bm{X}\in\mathbb{R}^{p}$ be a random vector of responses and $\bm{U}\in\mathbb{R}^{q}$ the corresponding covariates. In Zhang and Li [2022], the dependence of $\bm{X}$ on $\bm{U}$ is given by (1) and (3) so that both the mean and precision matrix are covariate-dependent. Write $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}$ and let $\Omega_{jk}(\boldsymbol{\mathbf{u}})$ denote the $(j,k)$-th element of $\boldsymbol{\mathbf{\Omega}}(\boldsymbol{\mathbf{u}})$, keeping in mind its dependence on $\boldsymbol{\mathbf{u}}$. Then (3) may be estimated via a neighborhood regression method [Meinshausen and Bühlmann, 2006], which amounts to estimating $p$ separate linear regression models of the form
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\beta_{jk0}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\beta_{jkh}u_{h}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\varepsilon_{j}$ (4)
where $\beta_{jkh}=-[\boldsymbol{\mathbf{B}}_{h}]_{jk}/\Omega_{jj}$ and $\varepsilon_{j}\sim N(0,1/\Omega_{jj})$ for all $j$, $k$, and $h$. This is termed Gaussian graphical regression in Zhang and Li [2022].
Note that a least squares criterion based on (4) is not jointly convex owing to the cross term $\beta_{jkh}\times\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k}$. Hence, Zhang and Li [2022] uses a two-stage estimation method. First, $\boldsymbol{\mathbf{\gamma}}_{j}$ is estimated for all $j$ via the model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\xi_{j},\quad\operatorname{\operatorname{\mathbb{E}}}\xi_{j}=0,$ (5)
and $\ell_{1}$ penalization is added so that $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ is sparse. Second, the observed response vectors $\boldsymbol{\mathbf{x}}_{j}$ are centered using $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ and the coefficients $\beta_{jkh}$ in (4) are estimated with $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ in place of $\boldsymbol{\mathbf{\gamma}}_{j}$ for all $j\in[p]$. As with other two-stage methods, the above procedure incurs an estimation error in the first stage because (5) ignores the dependence on the $\beta_{jkh}$ terms in (4). Assumptions on the scaling of the ambient dimensions $p$ and $q$ and the sparsity of the coefficient vectors are needed to suppress the model misspecification errors. Next, we will present a formulation of (1) so that the corresponding least squares criterion is convex, obviating the need for a two-stage procedure.
2.1 Convex formulation
Recall the form of the $p$-dimensional multivariate Gaussian likelihood expressed in terms of the natural parameters $(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}})$ where $\tilde{\boldsymbol{\mathbf{\theta}}}=\boldsymbol{\mathbf{\Omega}}\boldsymbol{\mathbf{\mu}}$ and $\tilde{\boldsymbol{\mathbf{\Theta}}}=-\boldsymbol{\mathbf{\Omega}}$:
$L(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}}\>|\>\boldsymbol{\mathbf{x}})=\exp\left\{\tilde{\boldsymbol{\mathbf{\theta}}}^{\top}\boldsymbol{\mathbf{x}}-\frac{1}{2}\boldsymbol{\mathbf{x}}^{\top}\tilde{\boldsymbol{\mathbf{\Theta}}}\boldsymbol{\mathbf{x}}-A(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}})\right\}.$
Our formulation is motivated by the fact that the cumulant function $A$ is jointly convex in $(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}})$ and therefore so is $-\log L(\tilde{\boldsymbol{\mathbf{\theta}}},\tilde{\boldsymbol{\mathbf{\Theta}}}\>|\>\boldsymbol{\mathbf{x}})$. Define
$\boldsymbol{\mathbf{\theta}}=\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})^{-1}\boldsymbol{\mathbf{\Omega}}\boldsymbol{\mathbf{\mu}}\quad\text{and}\quad\boldsymbol{\mathbf{\Theta}}=-\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})^{-1}\boldsymbol{\mathbf{\Omega}},$ (6)
where $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})$ is the $p\times p$ diagonal matrix of $\boldsymbol{\mathbf{\Omega}}$. Our model is to incorporate covariates as in (3) to $(\boldsymbol{\mathbf{\theta}},\boldsymbol{\mathbf{\Theta}})$ instead of $(\boldsymbol{\mathbf{\mu}},\boldsymbol{\mathbf{\Omega}})$, namely
$\boldsymbol{\mathbf{\theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}.$ (7)
To see that this leads to jointly convex nodewise regression problems, fix $j\in[p]$ and consider the partial regression of component $X_{j}$ against the covariates $\bm{U}$ and the remaining components $\bm{X}_{-j}$ in the general setting (1). For a matrix $\boldsymbol{\mathbf{M}}$ and sets of indices $\mathcal{I}$ and $\mathcal{J}$, denote by $\boldsymbol{\mathbf{M}}_{\mathcal{I},\mathcal{J}}$ the sub-matrix of $\boldsymbol{\mathbf{M}}$ consisting of rows indexed by $\mathcal{I}$ and columns indexed by $\mathcal{J}$. Letting $-j$ indicate all indices excluding $j$, we have the conditional distribution
$X_{j}-\mu_{j}\>|\>\bm{X}_{-j},\bm{U}\sim N\left(\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j}),\,\boldsymbol{\mathbf{\Sigma}}_{jj}-\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}\boldsymbol{\mathbf{\Sigma}}_{-j,j}\right),$ (8)
suppressing the dependence of $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\mu}}$ on the covariates $\boldsymbol{\mathbf{u}}$ for notational convenience. Defining the error terms $\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}})$ where $\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj}$, by the matrix block inversion formula we may write (8) as the linear model
$X_{j}=\mu_{j}-\Omega^{-1}_{jj}\boldsymbol{\mathbf{\Omega}}_{j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j})+\varepsilon_{j}=\theta_{j}+\boldsymbol{\mathbf{\Theta}}_{j,-j}\bm{X}_{-j}+\varepsilon_{j},$
which is readily seen to be jointly convex in the parameters $\theta_{j}$ and $\boldsymbol{\mathbf{\Theta}}_{j,-j}$ as defined in (6). In light of the specification of covariate dependence (7) and with $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}$ and $\beta_{jkh}=[\boldsymbol{\mathbf{B}}_{h}]_{jk}$, the above leads to the nodewise regression model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\beta_{jk0}X_{k}+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\beta_{jkh}u_{h}X_{k}+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}})$ (9)
which allows for simultaneous estimation of the parameters $\boldsymbol{\mathbf{\gamma}}_{j}$ and $\beta_{jkh}$ in a jointly convex setting.
Implicit in (6) is that $\Theta_{jj}=-1$ for all $j\in[p]$. This is satisfied in (7) by letting $[\boldsymbol{\mathbf{B}}_{0}]_{jj}=-1$ and $[\boldsymbol{\mathbf{B}}_{h}]_{jj}=0$ for all $h\in[q]$ and $j\in[p]$. By doing so, we assume as in Zhang and Li [2022] that the residual variance $\sigma^{2}_{\varepsilon_{j}}$ is not covariate-dependent, allowing us to write $\Omega_{jj}=\Omega_{jj}(\boldsymbol{\mathbf{u}})$ for $j\in[p]$.
3 Estimation
Suppose we collect $n$ independent observations $\{(\boldsymbol{\mathbf{x}}^{(i)},\boldsymbol{\mathbf{u}}^{(i)})\}_{i=1}^{n}$ of responses and covariates that follow the joint distribution in (1) with (6) and (7). Let $\boldsymbol{\mathbf{U}}=\left[\boldsymbol{\mathbf{u}}^{(1)},\dotsc,\boldsymbol{\mathbf{u}}^{(n)}\right]^{\top}\in\mathbb{R}^{n\times q}$ be the matrix of covariates and $\boldsymbol{\mathbf{X}}=\left[\boldsymbol{\mathbf{x}}^{(1)},\dotsc,\boldsymbol{\mathbf{x}}^{(n)}\right]^{\top}$ the matrix of responses with $\boldsymbol{\mathbf{x}}_{j}\in\mathbb{R}^{n}$ denoting the $j$-th column of $\boldsymbol{\mathbf{X}}$ and $\boldsymbol{\mathbf{u}}_{j}\in\mathbb{R}^{n}$ the $j$-th column of $\boldsymbol{\mathbf{U}}$. For $j\in[p]$, let $\boldsymbol{\mathbf{W}}_{-j}$ be the $n\times(p-1)(q+1)$ matrix of interactions of the remaining responses with the covariates. Concretely,
$\begin{split}\boldsymbol{\mathbf{W}}_{-j,0}&=\left[\boldsymbol{\mathbf{x}}_{1},\dotsc,\boldsymbol{\mathbf{x}}_{j-1},\boldsymbol{\mathbf{x}}_{j+1},\dotsc,\boldsymbol{\mathbf{x}}_{p}\right]\in\mathbb{R}^{n\times(p-1)},\\ \boldsymbol{\mathbf{W}}_{-j,h}&=\left[\boldsymbol{\mathbf{x}}_{1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{j-1}\odot\boldsymbol{\mathbf{u}}_{h},\ \boldsymbol{\mathbf{x}}_{j+1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{p}\odot\boldsymbol{\mathbf{u}}_{h}\right]\;\text{for $h\in[q]$},\\ \boldsymbol{\mathbf{W}}_{-j}&=\left[\boldsymbol{\mathbf{W}}_{-j,0},\ \boldsymbol{\mathbf{W}}_{-j,1},\ \dotsc,\ \boldsymbol{\mathbf{W}}_{-j,q}\right]\in\mathbb{R}^{n\times(p-1)(q+1)},\end{split}$ (10)
where $\odot$ denotes the elementwise product of two vectors. By writing
$\boldsymbol{\mathbf{\beta}}_{j,h}=(\beta_{j1h},\dotsc,\beta_{j,j-1,h},\beta_{j,j+1,h},\dotsc,\beta_{jph})^{\top}\in\mathbb{R}^{p-1}\;\text{and}\;\boldsymbol{\mathbf{\beta}}_{j}=(\boldsymbol{\mathbf{\beta}}_{j,0},\boldsymbol{\mathbf{\beta}}_{j,1},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q})^{\top}\in\mathbb{R}^{(p-1)(q+1)},$
we can view $\boldsymbol{\mathbf{\beta}}_{j}$ as $q+1$ blocks of $(p-1)$-element vectors $\boldsymbol{\mathbf{\beta}}_{j,h}$ for $h=\{0,1,\dotsc q\}$ so that (9) may be written in the block form
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\bm{X}_{-j}^{\top}\boldsymbol{\mathbf{\beta}}_{j,0}+\sum_{h=1}^{q}(u_{h}\bm{X}_{-j})^{\top}\boldsymbol{\mathbf{\beta}}_{j,h}+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}}).$ (11)
Figure 1 relates the blocks $\boldsymbol{\mathbf{\beta}}_{j,h}$ to $\boldsymbol{\mathbf{\Theta}}$ in (7).
The joint convexity of (11) in $(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j})$ allows for simultaneous estimation of the penalized problem
$\operatorname*{minimize}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+g_{j}(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j})$ (12)
where $g_{j}$ is an arbitrary convex penalty function. Moreover, by the relationship between $\boldsymbol{\mathbf{\beta}}_{j}$ and rows of $\boldsymbol{\mathbf{\Theta}}$, sparsity-inducing penalties $g_{j}$ in (12) will give us sparse estimates of $\boldsymbol{\mathbf{\Theta}}$. From (6), we see that $\boldsymbol{\mathbf{\Theta}}$ has the same sparsity pattern as $\boldsymbol{\mathbf{\Omega}}$. Hence we keep the conditional independence interpretation of the sparsity in $\boldsymbol{\mathbf{\Theta}}$ as in $\boldsymbol{\mathbf{\Omega}}$.
Figure 1: Decomposition of $\boldsymbol{\mathbf{\Theta}}$ into components $\boldsymbol{\mathbf{B}}_{h}$ according to (7). The block $\boldsymbol{\mathbf{\beta}}_{j,h}$ corresponds to the effects of covariate $u_{h}$ on the partial correlations of response $X_{j}$ while $\boldsymbol{\mathbf{\beta}}_{j,0}$ describes the population effect.
3.1 Estimating covariate-adjusted graphs
In the nodewise regression approach to GGMs, a symmetrization step is needed to ensure that the estimate of $\boldsymbol{\mathbf{\Omega}}$ is symmetric because each regression is fit independently [Meinshausen and Bühlmann, 2006]. In our approach, the nodewise regressions target $\boldsymbol{\mathbf{\Theta}}$, which is not necessarily symmetric due to the different scaling factor for each row in (6). To ensure symmetry in the estimate of $\boldsymbol{\mathbf{\Omega}}$, we perform a similar symmetrization step after performing all nodewise regressions. First, specify an estimate $\hat{\sigma}^{2}_{\varepsilon_{j}}$ of the error variance in (11). Then for an estimate $\hat{\boldsymbol{\mathbf{\beta}}}_{j}$, set $\tilde{\boldsymbol{\mathbf{\beta}}}_{j}=-\hat{\boldsymbol{\mathbf{\beta}}}_{j}/\hat{\sigma}^{2}_{\varepsilon_{j}}$ and define for all $h=0,1,\dotsc,q$ the symmetric matrix
$[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\left[\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\right]+\tilde{\beta}_{kjh}\left[\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\right],$ (13)
where the expression $\left[P\right]$ is equal to $1$ if $P$ is true and 00 otherwise. The mean vector and precision matrix may then be estimated by (6) via
$\hat{\boldsymbol{\mathbf{\Omega}}}(\boldsymbol{\mathbf{u}}^{(i)})$ $=\tilde{\boldsymbol{\mathbf{B}}}_{0}+\sum_{h=1}^{q}\tilde{\boldsymbol{\mathbf{B}}}_{h}u^{(i)}_{h}\quad\text{and}\quad\hat{\boldsymbol{\mathbf{\mu}}}(\boldsymbol{\mathbf{u}}^{(i)})=(\hat{\boldsymbol{\mathbf{\Omega}}}(\boldsymbol{\mathbf{u}}^{(i)}))^{-1}\operatorname{diag}(\hat{\boldsymbol{\mathbf{\Omega}}})\hat{\boldsymbol{\mathbf{\Gamma}}}\boldsymbol{\mathbf{u}}^{(i)}.$ (14)
The preceding steps are summarized in Algorithm 1 in Section S1.
Equation (13) is the “and-rule” to symmetrize the matrices $\tilde{\boldsymbol{\mathbf{B}}}_{h}$ for $h=0,1,\dotsc,q$; $[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}$ is nonzero if both $\tilde{\beta}_{jkh}$ and $\tilde{\beta}_{kjh}$ are nonzero. A less conservative estimate would be given by the “or-rule”, namely
$[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}{\left[\lvert\tilde{\beta}_{jkh}\rvert\geq\lvert\tilde{\beta}_{kjh}\rvert\right]}+\tilde{\beta}_{kjh}{\left[\lvert\tilde{\beta}_{jkh}\rvert\leq\lvert\tilde{\beta}_{kjh}\rvert\right]},$
so that $[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}$ is nonzero if either $\tilde{\beta}_{jkh}$ or $\tilde{\beta}_{kjh}$ is nonzero. We elect to use the more conservative rule and note that both approaches are asymptotically equivalent [Meinshausen and Bühlmann, 2006] and that the and-rule has been considered before in covariate-adjusted graphical models [Cai et al., 2013, Zhang and Li, 2022].
3.2 Sparsity structure
With the application to eQTL studies in mind, we will focus on a particular sparsity-inducing penalty $g_{j}$ in (15). We wish to identify elementwise sparsity within a group $h$, amounting to sparsity in the coefficient vector $\boldsymbol{\mathbf{\beta}}_{j,h}$. We may interpret sparse entries in $\boldsymbol{\mathbf{\beta}}_{j,h}$ to mean that the covariate $u_{h}$ affects the conditional independence between $X_{j}$ and some, but not all other responses $X_{k}$. Meanwhile, we wish to identify groupwise sparsity, where $\boldsymbol{\mathbf{\beta}}_{j,h}=\boldsymbol{\mathbf{0}}$ for a group $h$. This means that $u_{h}$ has no effect on the conditional independence between $X_{j}$ and the other responses.
We consider the following convex problem:
$\operatorname*{minimize}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}_{j}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}$ (15)
where $\lambda\geq 0$ and $\lambda_{g}\geq 0$ are tuning parameters and $\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}=\sum_{h=1}^{q}\lVert\boldsymbol{\mathbf{\beta}}_{j,h}\rVert_{2}$ is the group lasso penalty. Together, the penalty on $\boldsymbol{\mathbf{\beta}}_{j}$ is the sparse-group lasso penalty developed in Simon et al. [2013]. The group lasso penalty is not applied $\boldsymbol{\mathbf{\beta}}_{j,0}$ nor to $\boldsymbol{\mathbf{\gamma}}_{j}$.
We considered the following two estimates of $\sigma^{2}_{\varepsilon_{j}}$,
$\hat{\sigma}^{2}_{1}=\frac{\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\hat{\boldsymbol{\mathbf{\beta}}}_{j}\rVert_{2}^{2}}{n-\hat{s}_{\beta_{j}}-\hat{s}_{\gamma_{j}}}\quad\text{and}\quad\hat{\sigma}^{2}_{2}=\frac{\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\hat{\boldsymbol{\mathbf{\beta}}}_{j}\rVert_{2}^{2}}{n-\hat{s}_{\beta_{j}}-1},$
where $\hat{s}_{\gamma_{j}}$ and $\hat{s}_{\beta_{j}}$ are the number of nonzero elements of $\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$ and $\hat{\boldsymbol{\mathbf{\beta}}}_{j}$. Following the lasso literature [Reid et al., 2016, Yu and Bien, 2019], a straightforward estimate is given by $\hat{\sigma}^{2}_{1}$ as the value $\hat{s}_{\beta_{j}}+\hat{s}_{\gamma_{j}}$ roughly approximates the degrees of freedom of (15). In our experimentation, we have found $\hat{\sigma}^{2}_{2}$ to be slightly more consistent.
3.3 Implementation
We use block coordinate descent to solve (15) where the blocks are given by $\{\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j,0},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q}\}$. Since (15) is convex and separable in these blocks, this approach is guaranteed to converge to the optimal solution [Tseng, 2001]. Our R implementation uses the sparsegl package to solve (15) [Liang et al., 2024]. We use $k$-fold cross-validation to select the tuning parameter pair $(\lambda,\lambda_{g})$. To do so, we set a parameter $\lambda_{0}>0$ and mixture parameter $\alpha_{s}\in[0,1]$ so that the penalty in (15) may be written as
$\alpha_{s}\lambda_{0}(\lVert\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{1}+\lVert\boldsymbol{\mathbf{\beta}}_{j}\rVert_{1})+(1-\alpha_{s})\lambda_{0}\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}.$ (16)
Then for each $\alpha_{s}$, we run the method on a path of $\lambda_{0}$.
4 Theoretical Properties
In this section we will analyze the estimation error and support recovery of (15). For two sequences of real numbers $a_{n}$ and $b_{n}$, we write $a_{n}\precsim b_{n}$ if $a_{n}=O(b_{n})$, i.e. there exists constants $C>0$ and $N>0$ so that $a_{n}<Cb_{n}$ for all $n\geq N$. If $a_{n}\precsim b_{n}$ and $b_{n}\precsim a_{n}$, we write $a_{n}\asymp b_{n}$. We write $a_{n}=o(b_{n})$ if $\lim_{n\rightarrow\infty}a_{n}/b_{n}=0$. With the understanding that $j$ is fixed, we will suppress the subscript $j$ when referring to $\boldsymbol{\mathbf{\gamma}}_{j}$, $\boldsymbol{\mathbf{\beta}}_{j}$ and $\boldsymbol{\mathbf{\varepsilon}}_{j}$ for notational convenience.
Let $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ be the solution to (15) and let $(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})$ be the true parameters. Denote by $S_{\gamma}$ and $S_{\beta}$ the support sets of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ respectively. Further let $S_{\beta,g}$ index the active groups of $\boldsymbol{\mathbf{\beta}}$, that is $S_{\beta,g}=\left\{h:\boldsymbol{\mathbf{\beta}}_{j,h}\neq\boldsymbol{\mathbf{0}},\,h\in[q]\right\}$. Denote by $s_{\gamma},\,s_{\beta}$ and $s_{\beta,g}$ the cardinalities of these sets. For a square matrix $\boldsymbol{\mathbf{M}}$, denote by $\lambda_{\mathrm{min}}(\boldsymbol{\mathbf{M}})$ and $\lambda_{\mathrm{max}}(\boldsymbol{\mathbf{M}})$ its minimum and maximum eigenvalue, respectively.
We require the following assumptions on the model (1) with specification (7).
Assumption 1.
The covariates $\{\boldsymbol{\mathbf{u}}^{(i)}\}_{i=1}^{n}$ are i.i.d. mean zero random vectors with covariance matrix satisfying
$\phi_{0}\leq\lambda_{\mathrm{min}}(\operatorname{Cov}(\boldsymbol{\mathbf{u}}^{(i)}))\leq\lambda_{\mathrm{max}}(\operatorname{Cov}(\boldsymbol{\mathbf{u}}^{(i)}))\leq\phi_{1}$
for some constants $0<\phi_{0}\leq\phi_{1}<\infty$. Furthermore, there exists a constant $M>0$ such that $\lvert\boldsymbol{\mathbf{u}}^{(i)}_{j}\rvert\leq M$ for all $j\in[q]$ and $i\in[n]$.
Assumption 1 is the same as Assumptions 1 and 5 in Zhang and Li [2022] without requiring a bound on $\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}$. The boundedness of the covariates is not restrictive in the eQTL setting since SNPs are often binary coded.
Denote by $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ the $n\times(p(q+1)-1)$ matrix that is the concatenation of the covariate matrix $\boldsymbol{\mathbf{U}}$ with the interaction matrix $\boldsymbol{\mathbf{W}}_{-j}$. Define the following Gram matrix of covariates, responses, and interactions:
$\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}\left(\frac{[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]}{n}\right).$
Our next assumption bounds the eigenvalues of $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$, which is needed to characterize the joint distribution of $\boldsymbol{\mathbf{u}}^{(i)}$, $\boldsymbol{\mathbf{x}}^{(i)}$, and their interactions.
Assumption 2.
We assume there exist positive constants $m_{0},M_{0}$ such that
$m_{0}\leq\lambda_{\mathrm{min}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq\lambda_{\mathrm{max}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq M_{0}.$
Elementwise boundedness in Assumption 1 implies that $\boldsymbol{\mathbf{u}}^{(i)}$ is elementwise sub-Gaussian. Our theoretical analysis further requires that our design matrix $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ is elementwise sub-Gaussian. Assumption 3 ensures this since each entry of $\boldsymbol{\mathbf{W}}_{-j}$ is the product of a sub-Gaussian and a bounded random variable.
Assumption 3.
The marginal distribution of $\bm{X}$ is elementwise sub-Gaussian with bounded sub-Gaussian norm.
The following assumption controls how the sparsity of the true parameters may scale with the sample size $n$.
Assumption 4.
Let $p$ and $q$ be the number of responses and covariates, respectively. The true sparsities $s_{\gamma}$ and $s_{\beta}$ satisfy
  1. 1.
    $(s_{\gamma}+s_{\beta})\sqrt{\log(pq)}=o(\sqrt{n})$,
  2. 2.
    $(s_{\gamma}+s_{\beta})\log(pq)=o(n/\log n)$.
In practice we consider estimates $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ of varying sparsities and select from this pool of candidate models via cross-validation. Define $\hat{s}_{\gamma}^{\mathrm{max}}$ and $\hat{s}_{\beta}^{\mathrm{max}}$ to be the maximum sparsity of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ out of all candidate models, chosen so that $s_{\gamma}<s_{\gamma}^{\mathrm{max}}$ and $s_{\beta}<s_{\beta}^{\mathrm{max}}$. Our first theorem describes the $\ell_{2}$ estimation error of the nodewise solution.
Theorem 1.
Under Assumptions 1-4, with $(\hat{s}_{\gamma}^{\mathrm{max}}+\hat{s}_{\beta}^{\mathrm{max}})\log(pq)=O(\sqrt{n})$, and with
$\lambda=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\gamma}+s_{\beta}}+\frac{s_{\beta,g}}{s_{\gamma}+s_{\beta}}\log(eq/s_{\beta,g})\right)^{1/2}\quad\text{and}\quad\lambda_{g}=\lambda\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}$ (17)
we have
$\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{2}^{2}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}\precsim\frac{\sigma_{\varepsilon}^{2}}{n}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right).$
with probability at least $1-C_{1}\exp\left\{-C_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$ where $C$, $C_{1}$, and $C_{2}$ are positive constants.
The bound on the sparsity of candidate models helps to bound the errors in Theorem 1 and is also assumed in Zhang and Li [2022].
Our second result proves that our method achieves support recovery with high probability under some additional conditions. Denote the $(k,\ell)$-th entry of $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$ by $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$.
Assumption 5.
Define $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$. We assume that there exists a constant $c_{0}>2/m_{0}$ such that
$\max_{k\neq\ell}\lvert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\rvert\leq\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}.$
Theorem 2.
Suppose Assumptions 1-5 hold. Additionally, suppose $\log p\asymp\log q$ and that
$n\geq A_{1}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$
for some constant $A_{1}>0$. Then with $\lambda$ and $\lambda_{g}$ defined as in Theorem 1, we have
$\max\left\{\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{\infty},\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{\infty}\right\}\leq\frac{3}{m_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}m_{0}-2)}\right)$
with probability at least $1-C_{3}\exp(-C_{4}\log p)$ for some positive constants $C_{3},C_{4}$.
Theorem 2 implies that our method achieves support recovery with high probability if $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ satisfy a minimal signal strength condition.
Corollary 1.
Define
$\kappa_{0}$ $=\frac{3}{m_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}m_{0}-2)}\right),$
$\tilde{S}_{\gamma}$ =$=\left\{k:\lvert\hat{\gamma}_{k}\rvert>\kappa_{0}\right\},\quad\text{and}\quad\tilde{S}_{\beta}=\left\{k:\lvert\hat{\beta}_{k}\rvert>\kappa_{0}\right\}.$
If it holds that
$\min\left\{\min_{\ell\in S_{\gamma}}\lvert\gamma_{\ell}\rvert,\min_{k\in S_{\beta}}\lvert\beta_{k}\rvert\right\}\geq 2\kappa_{0},$
we have that $\operatorname{\operatorname{\mathbb{P}}}(S_{\gamma}=\tilde{S}_{\gamma}\;\text{and}\;S_{\beta}=\tilde{S}_{\beta})\geq 1-C_{3}\exp(-C_{4}\log p)$.
Zhang and Li [2022] provides two theorems related to the estimation error of $\hat{\boldsymbol{\mathbf{\beta}}}$; an oracle rate where the true $\boldsymbol{\mathbf{\Gamma}}$ in (3) is known and a rate where $\hat{\boldsymbol{\mathbf{\Gamma}}}$ is estimated via independent regressions in stage 1 of the procedure. It is shown that the oracle rate, given by
$\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}\precsim\frac{\sigma^{2}_{\varepsilon_{j}}}{n}(s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})),$
can be achieved by the two-stage rate under the assumptions $\log(pq)=O(n^{1/6})$, $s_{\beta}=o(n^{1/6})$, and $s_{\gamma}=o(n^{1/3})$. By comparison, Theorem 1 shows that our method achieves the oracle rate under milder assumptions; more explicitly, Assumption 4 is satisfied when $\log(pq)=O(n^{1/3})$ and $s_{\gamma}+s_{\beta}=o(n^{1/3})$. The milder scaling assumptions in $\log(pq)$ and $s_{\beta}$ in our formulation stems from the fact that (15) is jointly convex in $(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})$, allowing for their simultaneous estimation. By comparison, a more stringent scaling assumption is needed to control the estimation error of $\boldsymbol{\mathbf{\gamma}}_{j}$ incurred in stage 1 of Zhang and Li [2022]. These comments apply as well to the support recovery results in Theorem 2.
5 Simulation Study
In this section, we compare our method Covariate-adjusted Sparse Precision with Natural Estimation (cspine) with the Gaussian graphical model regression method developed in Zhang and Li [2022] (RegGMM) under extensive simulations. We are interested in the behavior of the two methods under data generated according to both models.
For each simulation setting, we generate a sparse $\boldsymbol{\mathbf{\Gamma}}\in\mathbb{R}^{p\times q}$ by independently determining if $\Gamma_{jk}$ is non-zero with probability $0.3$. The non-zero entries are sampled from $N(0,1)$. Then $\boldsymbol{\mathbf{\Gamma}}$ is scaled by a constant to ensure that the signal-to-noise ratio is approximately $1$ for each observation.
We generate the population graph $\boldsymbol{\mathbf{B}}_{0}$ using a preferential attachment algorithm [Barabási and Albert, 1999] with power $1$. We select $q_{e}=5$ out of $q$ covariates to have nonzero effects and generate the covariate-adjusted components $\boldsymbol{\mathbf{B}}_{h}$ for $h\in[q_{e}]$ as Erdős-Renyi graphs [Erdős and Rényi, 1959] with edge probability $v_{e}=0.01$. This choice of graph structure ensures that the covariate-dependent graphs are more sparse than the population graph; see Clauset et al. [2009] in particular for a discussion of power law graph models in applications. After determining the graph structure, we generate entries in $\{\boldsymbol{\mathbf{B}}_{h}\}$ by independently and uniformly sampling from $[-0.5,-0.35]\cup[0.35,0.5]$ and then dividing each row $j$ by $\sum_{h}\sum_{k\neq j}\lvert\beta_{jkh}\rvert\times 1.5$ before symmetrizing each $\{\boldsymbol{\mathbf{B}}_{h}\}$ by taking the average of $\beta_{jkh}$ and $\beta_{kjh}$. These steps help ensure a diagonally dominant precision matrix, and mirror the data model of Zhang and Li [2022].
We select half of the $q$ covariates to be binary and half to be continuous from $\operatorname{Unif}(0,1)$ and then standardize the continuous covariates across the $n$ observations to have zero mean and unit variance. With $\boldsymbol{\mathbf{\Gamma}}$ and $\{\boldsymbol{\mathbf{B}}_{h}\}$ in hand, we generate a precision matrix $\boldsymbol{\mathbf{\Omega}}^{(i)}=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}^{(i)}$. Then we generate a dataset under the original model (3) by setting $\boldsymbol{\mathbf{\mu}}^{(i)}=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$ and and a second dataset under (7), termed the “natural model”, via $\boldsymbol{\mathbf{\mu}}^{(i)}=\boldsymbol{\mathbf{\Sigma}}^{(i)}\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$. Finally, we sample a vector of responses $\boldsymbol{\mathbf{x}}^{(i)}\sim N_{p}(\boldsymbol{\mathbf{\mu}}^{(i)},\boldsymbol{\mathbf{\Sigma}}^{(i)})$ for $i\in[n]$. In both models, we set $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}}^{(i)})=1$ so that $\sigma^{2}_{\varepsilon_{j}}=1$.
For each of the settings $p=25,q=50$ and $p=25,q=100$, we generate $100$ independent data sets of $n=200$ and $n=400$ samples and run both RegGMM and cspine on each data set. For a fair comparison, both methods have two tuning parameters selected via $5$-fold cross-validation on a path of $100$ $\lambda$ parameters and $10$ mixture parameters. With the notation in (16), this means cross-validating over $100$ values of $\lambda_{0}$ and $\alpha_{s}=0.1,0.2,\dotsc,1$.
We report the mean and standard error of the following metrics after applying (13): TPR, the true positive rate of detected edges across all $\{\boldsymbol{\mathbf{B}}_{h}\}$; $\text{TPR}_{\text{pop}}$, the true positive rate of detected edges in the population graph $\boldsymbol{\mathbf{B}}_{0}$; $\text{FPR}_{\text{pop}}$, the false positive rate of detected edges in $\boldsymbol{\mathbf{B}}_{0}$; $\text{TPR}_{\text{cov}}$, the true positive rate of detected edges in covariate graphs $\boldsymbol{\mathbf{B}}_{h}$, $h\in[q]$; the nodewise estimation error given by $\boldsymbol{\mathbf{\beta}}_{\text{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{j}-\boldsymbol{\mathbf{\beta}}_{j}\rVert_{2}$; and the average error of the estimated precision matrix given by $\boldsymbol{\mathbf{\Omega}}_{\text{err}}=\sum_{i=1}^{n}\lVert{\hat{\boldsymbol{\mathbf{\Omega}}}}^{(i)}-\boldsymbol{\mathbf{\Omega}}^{(i)}\rVert^{2}_{F,\text{off}}/n$, where the norm is taken over off-diagonal entries. The overall false positive rate of detected edges is less than $0.005$ in all cases.
In Table 1, we see that cspine outperforms RegGMM under the natural model setting across all metrics. Table 2 compares the two methods under the setting (3) where RegGMM is correctly specified and cspine is misspecified. We see that despite the misspecification, cspine is competitive with RegGMM in edge detection and estimation error due to the convex formulation of the nodewise regressions. Both methods are conservative in their detection of edges in the covariate networks.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.73}$ $(0.08)$ $\mathbf{0.89}$ $(0.07)$ $\mathbf{0.07}$ $(0.02)$ $\mathbf{0.45}$ $(0.21)$ $\mathbf{5.66}$ $(0.50)$ $\mathbf{1.09}$ $(0.15)$
RegGMM $0.62$ $(0.09)$ $0.80$ $(0.10)$ $0.16$ $(0.06)$ $0.30$ $(0.18)$ $7.43$ $(0.61)$ $1.82$ $(0.26)$
(25, 100) cspine $\mathbf{0.66}$ $(0.10)$ $\mathbf{0.82}$ $(0.09)$ $\mathbf{0.05}$ $(0.02)$ $\mathbf{0.40}$ $(0.24)$ $\mathbf{5.98}$ $(0.55)$ $\mathbf{1.22}$ $(0.17)$
RegGMM $0.55$ $(0.08)$ $0.71$ $(0.08)$ $0.14$ $(0.05)$ $0.28$ $(0.19)$ $7.73$ $(0.56)$ $1.95$ $(0.28)$
400 (25, 50) cspine $\mathbf{0.77}$ $(0.10)$ $\mathbf{0.96}$ $(0.04)$ $\mathbf{0.09}$ $(0.02)$ $\mathbf{0.52}$ $(0.24)$ $\mathbf{4.50}$ $(0.59)$ $\mathbf{0.66}$ $(0.08)$
RegGMM $0.67$ $(0.08)$ $0.86$ $(0.07)$ $0.17$ $(0.07)$ $0.41$ $(0.20)$ $5.99$ $(0.55)$ $1.22$ $(0.17)$
(25, 100) cspine $\mathbf{0.72}$ $(0.09)$ $\mathbf{0.94}$ $(0.05)$ $\mathbf{0.08}$ $(0.02)$ $\mathbf{0.44}$ $(0.21)$ $\mathbf{4.96}$ $(0.60)$ $\mathbf{0.76}$ $(0.09)$
RegGMM $0.62$ $(0.08)$ $0.85$ $(0.08)$ $0.22$ $(0.07)$ $0.32$ $(0.17)$ $6.80$ $(0.58)$ $1.40$ $(0.20)$
Table 1: Mean and standard error of performance metrics over $100$ data sets.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.53}$ $(0.06)$ $\mathbf{0.77}$ $(0.08)$ $\mathbf{0.12}$ $(0.02)$ $0.09$ $(0.10)$ $\mathbf{7.70}$ $(0.29)$ $\mathbf{2.18}$ $(0.24)$
RegGMM $0.50$ $(0.07)$ $0.70$ $(0.10)$ $0.14$ $(0.05)$ $\mathbf{0.12}$ $(0.11)$ $8.02$ $(0.41)$ $2.20$ $(0.25)$
(25, 100) cspine $\mathbf{0.44}$ $(0.07)$ $\mathbf{0.67}$ $(0.09)$ $\mathbf{0.09}$ $(0.02)$ $0.04$ $(0.06)$ $\mathbf{7.97}$ $(0.32)$ $\mathbf{2.43}$ $(0.25)$
RegGMM $0.39$ $(0.07)$ $0.59$ $(0.12)$ $0.19$ $(0.07)$ $\mathbf{0.05}$ $(0.06)$ $8.80$ $(0.45)$ $2.69$ $(0.32)$
400 (25, 50) cspine $\mathbf{0.60}$ $(0.05)$ $\mathbf{0.92}$ $(0.05)$ $0.18$ $(0.03)$ $0.12$ $(0.10)$ $6.71$ $(0.28)$ $1.54$ $(0.18)$
RegGMM $0.59$ $(0.07)$ $0.80$ $(0.10)$ $\mathbf{0.14}$ $(0.05)$ $\mathbf{0.28}$ $(0.15)$ $\mathbf{6.64}$ $(0.42)$ $\mathbf{1.49}$ $(0.19)$
(25, 100) cspine $\mathbf{0.55}$ $(0.06)$ $\mathbf{0.88}$ $(0.07)$ $\mathbf{0.15}$ $(0.02)$ $0.12$ $(0.10)$ $\mathbf{6.83}$ $(0.26)$ $\mathbf{1.69}$ $(0.22)$
RegGMM $\mathbf{0.55}$ $(0.07)$ $0.81$ $(0.09)$ $0.22$ $(0.07)$ $\mathbf{0.19}$ $(0.14)$ $7.15$ $(0.42)$ $\mathbf{1.69}$ $(0.23)$
Table 2: Performance metrics over $100$ data sets under specification (3).
Next, we explore the error bound in Theorem 1 through simulations. We fixed $p=25$, $q=10$, and $n=300$. In the first set of simulations, we fixed a collection $\{\boldsymbol{\mathbf{B}}_{h}\}$ and generated $300$ data sets from $\boldsymbol{\mathbf{\Gamma}}$ matrices of varying sparsities. In the second set of simulations, we fixed a sparse $\boldsymbol{\mathbf{\Gamma}}$ and generated data sets from $300$ different collections of $\{\boldsymbol{\mathbf{B}}_{h}\}$ with varying sparsity levels. In Figure 2, we plot the error $\boldsymbol{\mathbf{\gamma}}_{\text{err}}+\boldsymbol{\mathbf{\beta}}_{\text{err}}$ where $\boldsymbol{\mathbf{\gamma}}_{\text{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{2}$ against the number of non-zero entries in $\boldsymbol{\mathbf{\Gamma}}$ and $\{\boldsymbol{\mathbf{B}}_{h}\}$ The shape of the plot verifies the theoretical results of Section 4, that the $\ell_{2}$ error scales with the square-root of the sparsity levels.
Figure 2: Estimation error of cspine over $300$ data sets plotted against the number of non-zero entries in (a) $\boldsymbol{\mathbf{\Gamma}}$ and (b) $\{\boldsymbol{\mathbf{B}}_{h}\}$. In both cases, the $\ell_{2}$ estimation error bound increases at a rate that is roughly the square-root of the sparsity level.
6 GBM eQTL Reanalysis
Glioblastoma multiforme (GBM) is the most malignant type of brain cancer and patient prognosis is typically very poor. Although there has been research on the genetic signaling pathways involved in the proliferation of GBM, it remains largely incurable; see Hanif et al. [2017] for a survey. It is important to understand the conditional independence structure of genes involved in GBM in order to discover new drug therapies [Kwiatkowska et al., 2013]. Our estimated graphs describe the conditional independence of co-expressions in a gene network; hence, we refer to estimated networks and effects of SNPs on this network.
We reanalyze a GBM eQTL data set that was reported in Zhang and Li [2022]. The data set contains microarray and SNP profiling data of $n=401$ GBM patients from the REMBRANDT trial (GSE108476) We use the expression levels of $p=73$ genes known to belong to the human glioma pathway according to the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [Kanehisa and Goto, 2000]; the genes and pathways are detailed in Table S3. We also consider $q=118$ SNPs that are local to these $73$ genes. The SNPs are binary-coded, with 00 indicating homozygous major alleles at that locus and $1$ otherwise. Our data set is slightly different from the one used in Zhang and Li [2022]; we have a larger cohort size (compared to $n=178$ in Zhang and Li 2022) and do not include age and sex as covariates.
We ran cspine on the data set using $5$-fold cross-validation over $\alpha_{s}=0.1,0.2,\dotsc,1$, and $100$ values of $\lambda_{0}$. Our method identified $56$ SNPs that potentially modify expressions in the network. However, many of the identified edges in these networks have small weights. Since cross-validation tends to select dense models, we focus on interpreting those entries of $\{\tilde{\boldsymbol{\mathbf{B}}}_{h}\}$ above $0.005$ in magnitude. With this threshold, $10$ SNPs are estimated to have effects on the network; see Table S4.
We also ran RegGMM on the data set for comparison, using $5$-fold cross-validation for a path of $100$ $\lambda$ parameters and $10$ mixture parameters. RegGMM identified $16$ SNPs with nonzero effects on the network. However, similarly to cspine, many of the estimated edges have very small weights. We choose to threshold these edges as well.
Figure 3 shows the estimated population (covariate-independent) network from cspine and RegGMM. It can be seen that many estimated edges overlap and that cspine estimates a somewhat denser network. For example, SHC4 and CALML4 are highly connected nodes in both population networks.
Figure 3: Population network from GBM eQTL data estimated by cspine (left) and RegGMM (right). The graph structure is determined by the estimates of $\boldsymbol{\mathbf{B}}_{0}$ in (3) and (7). Solid blue lines and dashed red lines indicate positive and negative edge weights, respectively.
Four SNPs are identified by RegGMM to affect the network of co-expressions. Out of the four, rs1267622 is also identified by cspine to have a nonzero effect. It is interesting to look at the estimated effect of rs1267622, a variant of the BRAF gene, by the two methods, shown in Figure 4.
Figure 4: Estimated covariate networks for the SNP rs1267622, a variant local to the BRAF gene. The graph structure corresponds to the sparsity pattern of the matrix $\boldsymbol{\mathbf{B}}_{h}$ corresponding to this SNP in (3) and (7). Solid blue lines and dashed red lines indicate positive and negative effects of this SNP on the partial correlation of co-expressions. An effect on the co-expression of GADD45A and GADD45B is detected by cspine while RegGMM detects an effect between PDGFRA and SOS2. The remaining two edges, between PDGRFA and PDGFB and PDGRFA and SOS2, are detected by both methods.
The results from both methods suggest that this SNP may modify the co-expressions of PDGFRA and PDGFB as well as PDGFRA and SOS2. This is plausible since all three of these genes lie in the Ras-Raf-MEK-ERK pathway along with BRAF (Table S3). There is evidence that variants on this pathway are associated with the proliferation of certain cancers [Gonzalez-Hormazabal et al., 2019]. Beyond this, the two methods differ; cspine estimates this variant to modify the co-expressions of GADD45A and GADD45B while RegGMM suggests that the co-expressions of PDGFRA and CAMK1 are modified.
[…]
The estimated connections among these four genes in the population network are shown in Figure 5. While RegGMM and cspine detect mostly the same edges in the population, only cspine finds that these co-expressions are modified by the presence of SNPs. These effects are summarized in Figure 5 (right). For instance, it is estimated that a variant local to the CALML4 gene may mediate the induction of the PI3K/Akt/mTOR pathway by expressions in E2F1. PIK3CA and E2F1 are part of interconnected signaling pathways implicated in cancer progression; E2F1 is involved in regulating the expression of PIK3CA and downstream signaling components [Ladu et al., 2008]. Hence, this discovery may give a clue for the role of CALML4 on regulating this pathway.
Figure 5: Estimated population network from cspine (left) and RegGMM (middle) highlighting four select genes. An edge indicates the partial correlation between the expressions of two genes conditional on the rest of the network (not shown). Right: the estimated effect of SNPs on co-expressions among the selected genes according to cspine. The gene to which the SNP is local is given in parentheses. Solid blue lines and dashed red lines indicate positive and negative edges, respectively. RegGMM did not detected any effects of SNPs on these co-expressions.
7 Discussion
In this work, we contributed to the covariate-adjusted Gaussian graphical model literature by developing a framework allowing for jointly convex optimization of the mean and precision matrix. Our theoretical work implies that the convex formulation allows for more relaxed scaling assumptions in the sparsities of $\boldsymbol{\mathbf{\beta}}_{j}$ in relation to the sample size $n$ and this is confirmed by our simulation results. Our method relies on tuning the pair $(\alpha_{s},\lambda_{0})$ over a grid as described in 3.3, which may be too computationally intensive for large datasets. If the practitioner has preconceptions about how many covariates are relevant in explaining the network structure of the responses, he may fix $\alpha_{s}$ to specify a desired group sparsity level. Alternatively, adapting tuning-free methods such as the square-root lasso [Belloni et al., 2011] to our method would be an important contribution for large datasets. A direction for theoretical work would be to establish the minimax rate of (15). Although our estimator achieves the same rate as in Zhang and Li [2022], the optimality of this rate has yet to be established. Our R package cspine is available on GitHub [^1] https://github.com/roobnloo/cspine.
Acknowledgements
We are grateful to Emma Zhang for generously making the RegGMM software and GBM eQTL data set available to us.
Use was made of computational facilities purchased with funds from the National Science Foundation (CNS-1725797) and administered by the Center for Scientific Computing (CSC). The CSC is supported by the California NanoSystems Institute and the Materials Research Science and Engineering Center (MRSEC; NSF DMR 2308708) at UC Santa Barbara.
Supplemental Materials
A convex formulation of covariate-adjusted Gaussian graphical models via natural parametrization
S1 Algorithm outline
Input :  $\boldsymbol{\mathbf{U}}\in\mathbb{R}^{n\times q}$ matrix of covariates.
$\boldsymbol{\mathbf{X}}\in\mathbb{R}^{n\times p}$ matrix of responses.
$g_{j}\colon\mathbb{R}^{q}\times\mathbb{R}^{(p-1)(q+1)}\rightarrow\mathbb{R}$ convex penalty functions for $j\in[p]$.
$H(\hat{\boldsymbol{\mathbf{\gamma}}}_{j},\hat{\boldsymbol{\mathbf{\beta}}}_{j})$ a function to estimate the error variance $\sigma^{2}_{\varepsilon_{j}}$.
Output :  $\hat{\boldsymbol{\mathbf{\Gamma}}}\in\mathbb{R}^{p\times q}$, $\tilde{\boldsymbol{\mathbf{B}}}_{0},\tilde{\boldsymbol{\mathbf{B}}}_{1},\dotsc\tilde{\boldsymbol{\mathbf{B}}}_{q}\in\mathbb{R}^{p\times p}$, and $\hat{\sigma}^{2}_{\varepsilon_{j}}$ for $j\in[p]$.
[…]
       Define $\boldsymbol{\mathbf{W}}_{-j}$ as in (10) and solve
$(\hat{\boldsymbol{\mathbf{\gamma}}}_{j},\hat{\boldsymbol{\mathbf{\beta}}}_{j})=\operatorname*{arg\,min}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+g_{j}(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j}).$
      Set $\hat{\sigma}^{2}_{\varepsilon_{j}}=H(\hat{\boldsymbol{\mathbf{\gamma}}}_{j},\hat{\boldsymbol{\mathbf{\beta}}}_{j})$ and $\tilde{\beta}_{jkh}=-\hat{\beta}_{jkh}/\hat{\sigma}^{2}_{\varepsilon_{j}}$ for $k\neq j$, $h=0,1,\dotsc,q$.
end for
Set $\hat{\boldsymbol{\mathbf{\Gamma}}}=[\hat{\boldsymbol{\mathbf{\gamma}}}_{1},\dotsc,\hat{\boldsymbol{\mathbf{\gamma}}}_{p}]^{\top}$.
Initialize $\tilde{\boldsymbol{\mathbf{B}}}_{0}=\tilde{\boldsymbol{\mathbf{B}}}_{1}=\dotsb=\tilde{\boldsymbol{\mathbf{B}}}_{q}=\mathbf{0}$.
[…]
       Apply the and-rule:
$[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\}}+\tilde{\beta}_{kjh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\}},\quad j\neq k$
end for
return $\hat{\boldsymbol{\mathbf{\Gamma}}},\tilde{\boldsymbol{\mathbf{B}}}_{0},\tilde{\boldsymbol{\mathbf{B}}}_{1},\dotsc,\tilde{\boldsymbol{\mathbf{B}}}_{q},\hat{\sigma}^{2}_{\varepsilon_{1}},\dotsc,\hat{\sigma}^{2}_{\varepsilon_{p}}$.
Algorithm 1 Covariate-adjusted Sparse Precision with Natural Estimation (cspine)
S2 Additional Simulation Results
In addition to the metrics reported in Table 1, we report here the support recovery metrics for the estimated precision matrix and the estimation error of the mean given by $\boldsymbol{\mathbf{\mu}}_{\text{err}}=\sum_{i=1}^{n}\lVert\hat{\boldsymbol{\mathbf{\mu}}}_{i}-\boldsymbol{\mathbf{\mu}}_{i}\rVert_{2}^{2}/n$ where $\hat{\boldsymbol{\mathbf{\mu}}}_{i}$ is computed from (14). The results are shown in Table S1 for data generated under the natural model (7) and in Table S2 for data generated under the original model (3). We see that cspine underperforms in mean prediction error in both data models. This is a reasonable finding since our method does not focus on estimating the mean. Instead, our method targets the precision matrix and its components $\{\boldsymbol{\mathbf{B}}_{h}\}$ while incorporating the covariate effects on the mean.
$n$ $(p,q)$ Method $\boldsymbol{\mathbf{\Omega}}_{\text{TPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{FPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$ $\boldsymbol{\mathbf{\mu}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.84}$ $(0.06)$ $\mathbf{0.10}$ $(0.03)$ $\mathbf{1.09}$ $(0.15)$ $10.14$ $(1.49)$
RegGMM $0.74$ $(0.08)$ $0.18$ $(0.06)$ $1.82$ $(0.26)$ $\mathbf{8.28}$ $(1.13)$
(25, 100) cspine $\mathbf{0.79}$ $(0.07)$ $\mathbf{0.08}$ $(0.02)$ $\mathbf{1.22}$ $(0.17)$ $11.94$ $(1.12)$
RegGMM $0.68$ $(0.07)$ $0.17$ $(0.04)$ $1.95$ $(0.28)$ $\mathbf{10.11}$ $(1.01)$
400 (25, 50) cspine $\mathbf{0.93}$ $(0.04)$ $\mathbf{0.12}$ $(0.03)$ $\mathbf{0.66}$ $(0.08)$ $5.80$ $(1.05)$
RegGMM $0.83$ $(0.06)$ $0.20$ $(0.07)$ $1.22$ $(0.17)$ $\mathbf{4.96}$ $(0.72)$
(25, 100) cspine $\mathbf{0.89}$ $(0.04)$ $\mathbf{0.11}$ $(0.03)$ $\mathbf{0.76}$ $(0.09)$ $7.98$ $(1.05)$
RegGMM $0.79$ $(0.06)$ $0.24$ $(0.07)$ $1.40$ $(0.20)$ $\mathbf{6.76}$ $(0.91)$
Table S1: Additional metrics under the natural model setting (7).
$n$ $(p,q)$ Method $\boldsymbol{\mathbf{\Omega}}_{\text{TPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{FPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$ $\boldsymbol{\mathbf{\mu}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.69}$ $(0.07)$ $\mathbf{0.14}$ $(0.03)$ $\mathbf{2.18}$ $(0.24)$ $6.20$ $(0.67)$
RegGMM $0.65$ $(0.09)$ $0.16$ $(0.05)$ $2.20$ $(0.25)$ $\mathbf{5.89}$ $(0.66)$
(25, 100) cspine $\mathbf{0.54}$ $(0.07)$ $\mathbf{0.12}$ $(0.02)$ $\mathbf{2.43}$ $(0.25)$ $9.95$ $(1.16)$
RegGMM $0.51$ $(0.09)$ $0.21$ $(0.06)$ $2.69$ $(0.32)$ $\mathbf{9.58}$ $(1.12)$
400 (25, 50) cspine $\mathbf{0.79}$ $(0.06)$ $0.20$ $(0.03)$ $1.54$ $(0.18)$ $3.34$ $(0.36)$
RegGMM $0.75$ $(0.07)$ $\mathbf{0.16}$ $(0.05)$ $\mathbf{1.49}$ $(0.19)$ $\mathbf{3.07}$ $(0.31)$
(25, 100) cspine $0.73$ $(0.06)$ $\mathbf{0.16}$ $(0.03)$ $\mathbf{1.69}$ $(0.22)$ $5.83$ $(0.63)$
RegGMM $\mathbf{0.74}$ $(0.07)$ $0.24$ $(0.07)$ $\mathbf{1.69}$ $(0.23)$ $\mathbf{5.58}$ $(0.54)$
Table S2: Additional metrics under specification (3).
S3 Additional Results from GBM eQTL Analysis
Table S3 displays the genes belonging to known signaling pathways associated with GBM. The table is copied from Zhang and Li [2022]. Table S4 lists the SNPs that are detected by cspine to affect network co-expression.
name genes references
PI3K/Akt/mTOR PIK3CA, PIK3CB, PIK3CD, Network [2008]
signaling pathway PIK3R3, PTEN, AKT1, AKT2, AKT3
MTOR, IGF1, PRKCA
EGF, EGFR, GRB2, Brennan et al. [2013]
SOS1, SOS2, IGF1
Ras-Raf-MEK-ERK SHC1, SHC2, SHC3, SHC4
signaling pathway MAPK1, MAPK3, MAP2K1, MAP2K2
HRAS, KRAS, NRAS,
RAF1, ARAF, BRAF, PRKCA
CALM1,CALML3, CALML4, CALML5, Maklad et al. [2019]
calcium (Ca+2) CALML6, CAMK1,CAMK4, CAMK1D,
signaling pathway CAMK1G,CAMK2A, CAMK2B,
CAMK2D,CAMK2G, PRKCA
p53 TP53, MDM2, DDB2, PTEN, IGF1 Network [2008]
signaling pathway CDK4, CDK6, CDKN1A, CDKN2A
Table S3: Gene signaling pathways related to GBM.
SNP co-expressed genes
rs1267622 (PDGFRA, PDGFB), (PDGRFA, SOS2), (GADD45A, GADD45B)
rs10519201 (E2F3, PIK3CA), (E2F3, BAX)
rs10488141 (SOS2, E2F1), (SOS2, GADD45B)
rs2076655 (E2F1, CALML5)
rs10509346 (E2F3, CALML5)
rs10518759 (E2F1, PIK3CA)
rs10512510 (E2F3, BRAF)
rs9303504 (E2F1, CALML5)
rs4834352 (E2F1, E2F3)
rs2075109 (PIK3CA, CALML5)
Table S4: Estimated SNP effects on gene co-expressions according to cspine.
S4 Proofs
Our proof strategies for Theorems 1 and 2 follow that of Zhang and Li [2022], with modifications made to accommodate our concatenated design matrix $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$. First we state a few lemmas.
Lemma 1 (Bellec et al. [2018] Lemma 1).
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be any convex function and let
$\hat{\boldsymbol{\mathbf{\beta}}}\in\operatorname*{arg\,min}_{\boldsymbol{\mathbf{\beta}}\in\mathbb{R}^{d}}\left\{\lVert\boldsymbol{\mathbf{y}}-\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}+g(\boldsymbol{\mathbf{\beta}})\right\}$
where $\boldsymbol{\mathbf{H}}\in\mathbb{R}^{n\times d}$ and $\boldsymbol{\mathbf{y}}\in\mathbb{R}^{n}$. Then for all $\boldsymbol{\mathbf{\beta}}\in\mathbb{R}^{d}$,
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{y}}-\boldsymbol{\mathbf{H}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}+g(\hat{\boldsymbol{\mathbf{\beta}}})+\frac{1}{2n}\lVert\boldsymbol{\mathbf{H}}(\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}})\rVert_{2}^{2}\leq\frac{1}{2n}\lVert\boldsymbol{\mathbf{y}}-\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}+g(\boldsymbol{\mathbf{\beta}}).$
Lemma 2 (Graybill and Marsaglia [1957] Theorem F).
Let $\boldsymbol{\mathbf{\varepsilon}}\sim N_{p}(\boldsymbol{\mathbf{0}},\sigma^{2}\boldsymbol{\mathbf{I}}_{p})$ and let $A$ be a $p\times p$ idempotent matrix with rank $r\leq p$. Then $\boldsymbol{\mathbf{\varepsilon}}^{\top}A\boldsymbol{\mathbf{\varepsilon}}/\sigma^{2}\sim\chi^{2}_{r}$.
Lemma 3 (Laurent and Massart [2000] Lemma 1).
Suppose that $U\sim\chi^{2}_{r}$. For any $x>0$ it holds that
$\operatorname{\operatorname{\mathbb{P}}}(U-r\geq 2\sqrt{rx}+2x)\leq e^{-x}.$
Lemma 4 (Vershynin [2011] Proposition 5.16).
Let $X_{1},\dotsc,X_{n}$ be independent, mean zero sub-exponential random variables. Let $v_{1}=\max_{i}\lVert X_{i}\rVert_{\psi_{1}}$ where $\lVert\cdot\rVert_{\psi_{1}}$ is the sub-exponential norm. Then there exists a constant $c$ such that for any $t>0$ we have
$\operatorname{\operatorname{\mathbb{P}}}\left(\left\lvert\sum_{i=1}^{n}X_{i}\right\rvert\geq t\right)\leq 2\exp\left\{-c\min\left(\frac{t^{2}}{v_{1}^{2}n},\frac{t}{v_{1}}\right)\right\}.$
Lemma 5 comes from Theorem 4.1 in Kuchibhotla and Chakrabortty [2022] applied to marginally sub-Gaussian random vectors.
Lemma 5 (Kuchibhotla and Chakrabortty [2022] Theorem 4.1).
Let $\boldsymbol{\mathbf{Z}}_{1},\dotsc,\boldsymbol{\mathbf{Z}}_{n}$ be independent random vectors in $\mathbb{R}^{p}$. Assume each element of $\boldsymbol{\mathbf{Z}}_{i}$ is sub-Gaussian with bounded sub-Gaussian norm for all $i\in[n]$. Let $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}=\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n$ and $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}}=\operatorname{\operatorname{\mathbb{E}}}[\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n]$. Define
$\Upsilon_{n}=\max_{j,k}\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}(Z_{ij}Z_{ik}).$
Then we have
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\lvert\boldsymbol{\mathbf{v}}^{\top}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}})\boldsymbol{\mathbf{v}}\right\rvert\precsim k\sqrt{\frac{\Upsilon_{n}\log p}{n}}+\frac{k\log n\log p}{n}$
with probability at least $1-O(1/p)$.
We may apply Lemma 5 because by Assumption 3, $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ is elementwise sub-Gaussian, each entry being the product of a sub-Gaussian and a bounded random variable. Furthermore, we have by the Cauchy-Schwarz inequality that
$\max_{\ell,k}\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{i\ell}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{ik}\right)\leq\max_{\ell_{1},\ell_{2},\ell_{3},\ell_{4}}\operatorname{\operatorname{\mathbb{E}}}\left(X_{\ell_{1}}^{(1)^{2}}X_{\ell_{2}}^{(1)^{2}}U_{\ell_{3}}^{(1)^{2}}U_{\ell_{4}}^{(1)^{2}}\right)=O(1)$
since the entries of $\boldsymbol{\mathbf{X}}^{(1)}$ and $\boldsymbol{\mathbf{U}}^{(1)}$ have bounded moments. Thus $\Upsilon_{n}=O(1)$ in our setting.
Lemma 6 (Loh and Wainwright [2012] Lemma 12).
Let $\boldsymbol{\mathbf{\Sigma}}\in\mathbb{R}^{p\times p}$ be a symmetric matrix such that $\lvert\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}\boldsymbol{\mathbf{v}}\rvert\leq\delta_{1}$ for all $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p}$ with $\lVert\boldsymbol{\mathbf{v}}\rVert_{2}=1$ and $\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq 2s$. It holds for all $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p}$ that
$\lvert\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}\boldsymbol{\mathbf{v}}\rvert\leq 27\delta_{1}\left(\lVert\boldsymbol{\mathbf{v}}\rVert_{2}^{2}+\frac{1}{s}\lVert\boldsymbol{\mathbf{v}}\rVert_{1}^{2}\right).$
Define the matrices:
$\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}],\quad\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}}).$
Lemma 7.
For a set of indices $S\subset[p(q+1)-1]$, denote by $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{S}$ the submatrix of $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]$ with columns indexed by $S$. Let $\lVert\cdot\rVert_{\mathrm{op}}$ denote the matrix operator norm. Under Assumptions 1-4, there exist constants $M_{uw}$ and $C_{0}$ such that with probability at least $1-C_{0}\exp(-\log(pq))$ we have
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}_{-j}]_{S}\right\rVert_{\mathrm{op}}^{2}\leq M_{uw}$
for all $S$ satisfying $\lvert S\rvert\leq\hat{s}^{\mathrm{max}}_{\gamma}+\hat{s}^{\mathrm{max}}_{\beta}$, provided that $(\hat{s}_{\gamma}^{\mathrm{max}}+\hat{s}_{\beta}^{\mathrm{max}})\log(pq)=O(\sqrt{n})$, as assumed in Theorem 1.
Proof.
Letting $k=\lvert S\rvert$, it suffices to show that $\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\boldsymbol{\mathbf{v}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ is bounded. We may write
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\boldsymbol{\mathbf{v}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ $=\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\{\boldsymbol{\mathbf{v}}^{\top}\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}\right\}.$
By Assumption 2, the second term is bounded. For the first term, by Lemma 5 we have
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\{\boldsymbol{\mathbf{v}}^{\top}\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}\right\}\precsim\left(\frac{k^{2}\log(pq)}{n}\right)^{1/2}+\frac{k\log(pq)}{n/\log n}$
with probability at least $1-C_{0}\exp(-\log(pq))$ and we see that the right-hand side is bounded by Assumption 4. ∎
S4.1 Proof of Theorem 1
For ease of notation, we will drop the dependence of $\boldsymbol{\mathbf{\gamma}}_{j}$, $\boldsymbol{\mathbf{\beta}}_{j}$, $\boldsymbol{\mathbf{\varepsilon}}_{j}$ and $\boldsymbol{\mathbf{W}}_{-j}$ on $j$. Let $S_{\beta},S_{\gamma},\hat{S}_{\beta},\hat{S}_{\gamma}$ be the support sets of $\boldsymbol{\mathbf{\beta}},\boldsymbol{\mathbf{\gamma}},\hat{\boldsymbol{\mathbf{\beta}}},\hat{\boldsymbol{\mathbf{\gamma}}}$, respectively. Let $S_{\beta,g}=\{h\in[q]:\boldsymbol{\mathbf{\beta}}_{h}\neq\boldsymbol{\mathbf{0}}\}$ index the blocks $\boldsymbol{\mathbf{\beta}}_{h}$ of $\boldsymbol{\mathbf{\beta}}$ that are not identically zero and let $\hat{S}_{\beta,g}$ be the corresponding block indices for $\hat{\boldsymbol{\mathbf{\beta}}}$. For any vector $\boldsymbol{\mathbf{v}}$ and set of block indices $S$, let $\boldsymbol{\mathbf{v}}_{(S)}$ denote the sub-vector containing blocks in $S$. Let $s_{\beta},s_{\gamma},s_{\beta,g},\hat{s}_{\beta},\hat{s}_{\gamma},\hat{s}_{\beta,g}$ be the number of elements in $S_{\beta},S_{\gamma},S_{\beta,g},\hat{S}_{\beta},\hat{S}_{\gamma},\hat{S}_{\beta,g}$, respectively. Our proof occurs in three steps.
Step 1
In this step we bound the error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}/n$ by the stochastic term $\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle/n$, which is then bounded by a projection of $\boldsymbol{\mathbf{\varepsilon}}$ onto the columns of $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$.
Since our penalty function
$g(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})=\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}$
is convex, by Lemma 1 we have
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}+g(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})+\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\leq\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}+g({\boldsymbol{\mathbf{\gamma}}},{\boldsymbol{\mathbf{\beta}}})$
where $\boldsymbol{\mathbf{\nu}}=\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\Delta}}=\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}$. Since $\boldsymbol{\mathbf{\varepsilon}}=\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\beta}}$ we may write
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}$ $=\frac{1}{2n}\lVert\boldsymbol{\mathbf{\varepsilon}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$
$=\frac{1}{2n}\lVert\boldsymbol{\mathbf{\varepsilon}}\rVert_{2}^{2}-\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}.$
Plugging this into the previous expression and substituting the penalty expression then yields
$\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}.$
Notice that $\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S^{c}_{\beta}}\rVert_{1}$, $\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}}\rVert_{1}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}^{c}}\rVert_{1}$, and $\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}=\lVert\boldsymbol{\mathbf{\beta}}_{S_{\beta}}\rVert_{1}$. Hence we can express the above as
$\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\underbrace{\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}}_{\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\underbrace{\lambda\lVert\boldsymbol{\mathbf{\beta}}_{S_{\beta}}\rVert_{1}}_{\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}.$
Thus we have
$\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda(\lVert\boldsymbol{\mathbf{\beta}}_{S_{\beta}}\rVert_{1}-\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}}\rVert_{1})+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}$
using the triangle inequality for the $\ell_{1}$ norm. The same development holds for $\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}$. Finally, notice that
$\lVert\boldsymbol{\mathbf{\Delta}}_{(S^{c}_{\beta,g})}\rVert_{1,2}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{(S^{c}_{\beta,g})}\rVert_{1,2},\;\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{(S_{\beta,g})}\rVert_{1,2}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{(S^{c}_{\beta,g})}\rVert_{1,2},\;\text{and}\;\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}=\lVert\boldsymbol{\mathbf{\beta}}_{(S_{\beta,g})}\rVert_{1,2}.$
Hence the previous development holds for the $\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$ term by the triangle inequality of $\lVert\cdot\rVert_{1,2}$. All in all we have
$\begin{split}&\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g}^{c})}\rVert_{1,2}\\ &\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}.\end{split}$ (S1)
Now let $\mathcal{I}$ and $\mathcal{J}$ be arbitrary index sets of the columns of $\boldsymbol{\mathbf{U}}$ and $\boldsymbol{\mathbf{W}}$ respectively. Denote by $\mathcal{P}_{\mathcal{I},\mathcal{J}}$ the orthogonal projection onto the columns of $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ indexed by $(\mathcal{I},\mathcal{J})$. Let $\mathcal{I}_{0}=S_{\gamma}\cup\hat{S}_{\gamma}$ and $\mathcal{J}_{0}=S_{\beta}\cup\hat{S}_{\beta}$ denote the unions of the true and estimated support sets of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$. We seek to bound the stochastic term
$\begin{split}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle&=\langle\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}}),\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle\\ &\leq\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}\leq\frac{1}{2a_{1}}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{a_{1}}{2}\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}.\end{split}$ (S2)
The last inequality follows from the fact that $2xy\leq ax^{2}+y^{2}/a$ holds for any constant $a>0$ and real numbers $x$ and $y$.
Following Zhang and Li [2022], we first bound the term $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ with a counting argument. For fixed $s^{\prime}_{\gamma}$, $s^{\prime}_{\beta}$, and $s^{\prime}_{\beta,g}$, we will bound the cardinality of the set
$\mathcal{H}(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})=\{(\mathcal{I},\mathcal{J})\subset[q]\times[(p-1)(q+1)]:\lvert\mathcal{I}\rvert=s^{\prime}_{\gamma},\,\lvert\mathcal{J}\rvert=s^{\prime}_{\beta},\,\lvert g(\mathcal{J})\rvert=s^{\prime}_{\beta,g}\}$
where $g(\mathcal{J})$ is the number of nonzero groups of $\boldsymbol{\mathbf{\beta}}_{\mathcal{J}}$. For ease of notation, we will write $\mathcal{H}$ while keeping in mind its dependence on $s^{\prime}_{\gamma},s^{\prime}_{\beta},$ and $s^{\prime}_{\beta,g}$. We will show that
$\log\lvert\mathcal{H}\rvert\leq s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)$
by considering the two cases $s^{\prime}_{\beta}=s^{\prime}_{\beta,g}$ and $s^{\prime}_{\beta}>s^{\prime}_{\beta,g}$ separately. These are the only cases since we cannot have more nonzero groups than nonzero elements.
  1. 1.
    $s^{\prime}_{\beta}=s^{\prime}_{\beta,g}$: In this case, we have $\lvert\mathcal{H}\rvert\leq\binom{q}{s^{\prime}_{\gamma}}\binom{q}{s^{\prime}_{\beta,g}}(p-1)^{s^{\prime}_{\beta}}$. Hence
    $\log\lvert\mathcal{H}\rvert$ $\leq\log\binom{q}{s^{\prime}_{\gamma}}+\log\binom{q}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(p-1)$
    $\leq s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)$
    where we use $\log\binom{n}{k}\leq k\log(en/k)$ which follows from Stirling’s approximation.
  2. 2.
    $s^{\prime}_{\beta}>s^{\prime}_{\beta,g}$: In this case, the cardinality is bounded by
    $\lvert\mathcal{H}\rvert\leq\binom{q}{s^{\prime}_{\gamma}}\binom{q}{s^{\prime}_{\beta,g}}\binom{(p-1)(s^{\prime}_{\beta,g}+1)}{s^{\prime}_{\beta}}.$
    Since by Stirling’s approximation
    $\log\binom{(p-1)(s^{\prime}_{\beta,g}+1)}{s^{\prime}_{\beta}}\leq s^{\prime}_{\beta}\log\frac{e(p-1)(s^{\prime}_{\beta,g}+1)}{s^{\prime}_{\beta}}\leq s^{\prime}_{\beta}\log(ep),$
    we have
    $\log\lvert\mathcal{H}\rvert\leq s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)$
    as desired.
Define $k_{0}$ to be the exponential of the right hand side of the above inequality, so that $\lvert\mathcal{H}\rvert\leq k_{0}$. For any $(\mathcal{I},\mathcal{J})\in\mathcal{H}$, since $\mathcal{P}_{\mathcal{I},\mathcal{J}}$ is idempotent, Lemma 2 implies
$\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}/\sigma^{2}_{\varepsilon}\sim\chi_{d}^{2}$
where $d\leq\lvert\mathcal{I}\rvert+\lvert\mathcal{J}\rvert=s^{\prime}_{\gamma}+s^{\prime}_{\beta}$ is the rank of $\mathcal{P}_{\mathcal{I},\mathcal{J}}$. By Lemma 3 we have for arbitrary $t^{\prime}>0$ that
$\operatorname{\operatorname{\mathbb{P}}}\left(\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq\sigma^{2}_{\varepsilon}(2\sqrt{dt^{\prime}}+d+2t^{\prime})\right)\leq e^{-t^{\prime}}.$
Since $2\sqrt{dt^{\prime}}\leq d+t^{\prime}$ and $d\leq s^{\prime}_{\gamma}+s^{\prime}_{\beta}\leq\log k_{0}$, we have
$\operatorname{\operatorname{\mathbb{P}}}\left(\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq\sigma^{2}_{\varepsilon}(2\log k_{0}+3t^{\prime})\right)\leq e^{-t^{\prime}}.$
Taking the supremum over $\mathcal{H}$ and applying the union bound yields
$\operatorname{\operatorname{\mathbb{P}}}\left(\sup_{(\mathcal{I},\mathcal{J})\in\mathcal{H}}\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq\sigma^{2}_{\varepsilon}(2\log k_{0}+3t^{\prime})\right)\leq\lvert\mathcal{H}\rvert e^{-t^{\prime}}.$
Now set $t^{\prime}=t/3+\log k_{0}$ for $t>0$. Then $\lvert\mathcal{H}\rvert e^{-t^{\prime}}=\lvert\mathcal{H}\rvert/{k_{0}}\cdot e^{-t/3}\leq e^{-t/3}$ since $\lvert\mathcal{H}\rvert\leq k_{0}$. Substituting these expressions into the previous bound (and recalling that we have fixed $s^{\prime}_{\gamma}$, $s^{\prime}_{\beta}$, and $s^{\prime}_{\beta,g}$) yields
$\operatorname{\operatorname{\mathbb{P}}}\left(\sup_{(\mathcal{I},\mathcal{J})\in\mathcal{H}(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})}\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq 5\sigma^{2}_{\varepsilon}{\left[s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)\right]}+\sigma^{2}_{\varepsilon}t\right)\leq e^{-t/3},$ (S3)
recalling that
$\log k_{0}=s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep).$
This gives a concentration bound of $\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ over all possible subsets $\mathcal{I}$ of columns of $\boldsymbol{\mathbf{U}}$ and $\mathcal{J}$ of columns of $\boldsymbol{\mathbf{W}}$ satisfying $\lvert\mathcal{I}\rvert=s^{\prime}_{\gamma}$, $\lvert\mathcal{J}\rvert=s^{\prime}_{\beta}$, and $\lvert g(\mathcal{J})\rvert=s^{\prime}_{\beta,g}$. Recalling that $\mathcal{I}_{0}$ and $\mathcal{J}_{0}$ are the support sets of $\boldsymbol{\mathbf{\nu}}$ and $\boldsymbol{\mathbf{\Delta}}$, we can now bound $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$. Define
$r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})=\left(\sup_{(\mathcal{I},\mathcal{J})\in\mathcal{H}(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})}\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}-5\sigma^{2}_{\varepsilon}\left\{s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)\right\}\right)_{+}$
and
$r=\sup_{s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}}r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}).$
It is clear that
$\lvert\mathcal{I}_{0}\rvert\leq s_{\gamma}+\hat{s}_{\gamma},\;\lvert\mathcal{J}_{0}\rvert\leq s_{\beta}+\hat{s}_{\beta},\;\text{and}\;\lvert g(\mathcal{J}_{0})\rvert\leq s_{\beta,g}+\hat{s}_{\beta,g}.$
Thus we have
$\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\leq 5\sigma^{2}_{\varepsilon}\left\{(s_{\gamma}+\hat{s}_{\gamma})\log\frac{eq}{s_{\gamma}}+(s_{\beta}+\hat{s}_{\beta})\log(ep)+(s_{\beta,g}+\hat{s}_{\beta,g})\log\frac{eq}{s_{\beta,g}}\right\}+r.$ (S4)
We also have the following concentration inequality on $r$ for $t>0$, which comes from the definition of $r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})$ along with (S3):
$\operatorname{\operatorname{\mathbb{P}}}(r\geq t\sigma^{2}_{\varepsilon})$ $\leq\sum_{s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}}\operatorname{\operatorname{\mathbb{P}}}(r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})\geq t\sigma^{2}_{\varepsilon})\leq\sum_{s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}}e^{-t/3}\leq q^{3}p\,c_{1}e^{-t/3},$
where $c_{1}>0$ is a constant and the sum is taken over $s^{\prime}_{\gamma}\in[q],\,s^{\prime}_{\beta}\in[(p-1)(q+1)],\,s^{\prime}_{\beta,g}\in[q]$. Then for a sufficiently large constant $\tilde{M}$ and some $c_{2}>0$, by letting
$t=\tilde{M}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))$
we have
$\begin{split}&\operatorname{\operatorname{\mathbb{P}}}\left\{r\geq\tilde{M}\sigma^{2}_{\varepsilon}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}\\ &\qquad\leq c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}.\end{split}$ (S5)
Step 2
We now use the KKT optimality conditions to bound $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ in terms of $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$. This takes care of the stochastic term after plugging back into (S1) and (S2). To ease the notation when describing the conditions, define the vectors
$\boldsymbol{\mathbf{r}}_{\gamma}=\frac{1}{n}\boldsymbol{\mathbf{U}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}),\quad\boldsymbol{\mathbf{r}}_{\beta}=\frac{1}{n}\boldsymbol{\mathbf{W}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})$
Let $\hat{\boldsymbol{\mathbf{\beta}}}_{h}$ be the $h$-th block of $\hat{\boldsymbol{\mathbf{\beta}}}$. By the KKT conditions, we know that an optimizer $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ of (15) must satisfy
$\left(\boldsymbol{\mathbf{r}}_{\gamma}\right)_{\ell}$ $=\lambda\,\operatorname{sign}(\hat{\gamma}_{\ell})$ for $\hat{\gamma}_{\ell}\neq 0$ (S6)
$(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}$ $=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell})$ for $(\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell}\neq 0$ (S7)
$(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}$ $=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell})+\lambda_{g}\,\frac{(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}}{\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{h}\rVert_{2}}$ for ($\text{for $(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}\neq 0,\,h\in[q]$}.$ (S8)
Squaring both sides of (S6) and summing over $\ell$ gives
$\lambda^{2}\hat{s}_{\gamma}=\frac{1}{n^{2}}\lVert\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\rVert_{2}^{2}$
and doing the same for (S7) and (S8) gives
$\lambda^{2}\hat{s}_{\beta}+\lambda_{g}^{2}\hat{s}_{\beta,g}\leq\frac{1}{n^{2}}\lVert\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\rVert_{2}^{2}$
since in (S8) the cross term $\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{j,h})_{\ell})\times(\hat{\boldsymbol{\mathbf{\beta}}}_{j,h})_{\ell}$ is nonnegative. We have shown that
$\begin{split}\lambda^{2}\hat{s}_{\gamma}+\lambda^{2}\hat{s}_{\beta}+\lambda_{g}^{2}\hat{s}_{\beta,g}&\leq\frac{1}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\right\rVert_{2}^{2}\\ &=\frac{1}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}(\boldsymbol{\mathbf{\varepsilon}}-\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}-\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}})\right\rVert_{2}^{2}.\end{split}$ (S9)
Using that $\lVert\boldsymbol{\mathbf{a}}-\boldsymbol{\mathbf{b}}\rVert_{2}^{2}\leq 2\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}+2\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}$ for any vectors $\boldsymbol{\mathbf{a}}$ and $\boldsymbol{\mathbf{b}}$ along with Lemma 7, we have with probability at least $1-C_{0}\exp(-\log(pq))$ that
$\begin{split}\lambda^{2}\hat{s}_{\gamma}+\lambda^{2}\hat{s}_{\beta}+\lambda_{g}^{2}\hat{s}_{\beta,g}&\leq\frac{2}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{2}^{2}+\frac{2}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}\left(\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}+\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}}\right)\right\rVert_{2}^{2}\\ &\leq\frac{2}{n}M_{uw}\left\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\right\rVert_{2}^{2}+\frac{2}{n}M_{uw}\left\lVert\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}+\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}}\right\rVert_{2}^{2},\end{split}$ (S10)
where the last inequality uses that $\hat{S}_{\gamma}\subset\mathcal{I}_{0}$ and $\hat{S}_{\beta}\subset\mathcal{J}_{0}$.
Now let
$\lambda$ $=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\gamma}+s_{\beta}}+\frac{s_{\beta,g}}{s_{\gamma}+s_{\beta}}\log(eq/s_{\beta,g})\right)^{1/2},$
$\lambda_{g}$ $=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\beta,g}}+\log(eq/s_{\beta,g})\right)^{1/2},$
where $C=\sqrt{5M_{uw}a_{2}}$ for some $a_{2}>2$. Combining (S10) and (S4) gives
$\begin{split}&\left(1-\frac{2}{a_{2}}\right)\left\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\right\rVert_{2}^{2}\\ &\qquad\leq 5\sigma^{2}_{\varepsilon}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)+\frac{2}{a_{2}}\left\lVert\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}+\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}}\right\rVert_{2}^{2}+r\end{split}$ (S11)
It is then straightforward to multiply both sides by $a_{1}a_{2}/(2(a_{2}-2))$ and plug into (S1) and (S2) to get
$\begin{split}&\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S^{c}_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S^{c}_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S^{c}_{\beta,g}}\rVert_{1,2}\\ &\qquad\leq\frac{1}{2a_{1}}\cdot\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{5a_{1}a_{2}}{2(a_{2}-2)}\cdot E_{j}+\frac{a_{1}}{a_{2}-2}\cdot\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\\ &\qquad\quad+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}\end{split}$ (S12)
where $E_{j}=(\sigma^{2}_{\varepsilon}/n)\cdot\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$.
We also have that
$\frac{\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}+\frac{\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}}{\sqrt{s_{\beta,g}}}$ $\leq\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{2}$
$\leq 2\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}\leq 2m_{0}^{-1/2}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
where the last inequality follows from Assumption 2. With our chosen values of $\lambda$ and $\lambda_{g}$, we have after multiplying through by $C\sqrt{E_{j}}$ that
$\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}$ $\leq 2Cm_{0}^{-1/2}\sqrt{E_{j}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
$\leq a_{3}\frac{C^{2}}{m_{0}}E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$
for any $a_{3}>0$. Plugging this into (LABEL:eqn:s11) gives
$\begin{split}&\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\\ &\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{m_{0}}\right)E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}.\end{split}$ (S13)
Step 3
We now bound the difference between $\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}/n$ and $\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$. Let $L>0$ be an arbitrarily large constant. By Assumption 4 and Lemma 5, we have with probability at least $1-C^{\prime}\exp(-\log(pq))$ that
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq 2(s_{\gamma}+s_{\beta}),\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}=1}\left\lvert\boldsymbol{\mathbf{v}}^{\top}\left(\frac{[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]}{n}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}\right\rvert\leq\frac{1}{27L}$
for sufficiently large $n$. By Lemma 6, it holds with probability at least $1-C^{\prime}\exp(-\log(pq))$ that
$\begin{split}&\left\lvert(\boldsymbol{\mathbf{\nu}}^{\top},\,\boldsymbol{\mathbf{\Delta}}^{\top})\left(\frac{[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]}{n}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rvert\\ &\qquad\leq\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right)\end{split}$ (S14)
for sufficiently large $n$. Plugging this into (S13) gives
$\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}-\frac{1}{a_{3}}\right)\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{m_{0}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
$\qquad+\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\left(\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}-\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\right)$
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{m_{0}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
$\qquad+\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\cdot\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right).$
Recalling that the above holds for any $a_{1}>0$, $a_{2}>2$, and $a_{3}>0$, choose $a_{1}=2$, $a_{2}=6$, and $a_{3}=6$ to get
$\frac{1}{2}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right)+\frac{r}{n}.$
By the concentration bound on $r$ in (S5), we in fact have
$\frac{1}{2}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right)$ (S15)
with probability at least $1-c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$. Then by Assumption 2 we have with the same high probability
$\frac{m_{0}}{2}(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2})\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right).$ (S16)
Next, taking $a_{1}=2-\sqrt{2}$, $a_{2}=6$ in (LABEL:eqn:s11) cancels out the $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}/n$ terms. Appealing again to (S5) yields
$\frac{\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}+\frac{\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g}^{c})}\rVert_{1,2}}{\sqrt{s_{\beta,g}}}\leq\sqrt{E_{j}}+\frac{\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}+\frac{\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}}{\sqrt{s_{\beta,g}}}.$
with probability at least $1-c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$.
Adding $(\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1})/\sqrt{s_{\gamma}+s_{\beta}}$ to both sides and using
$\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}\leq\sqrt{s_{\gamma}+s_{\beta}}(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2})\quad\text{and}\quad\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}\leq\sqrt{s_{\beta,g}}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}$
yields
$\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}\leq\sqrt{E_{j}}+2\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}+3\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}$
and after squaring both sides we have
$\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\leq k_{0}E_{j}+k_{1}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+k_{2}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$ (S17)
for some absolute constants $k_{0},k_{1},k_{2}>0$. Plugging (S17) into the right-hand side of (S16) yields
$\frac{m_{0}}{2}(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2})\precsim\left(1+\frac{k_{0}}{L}\right)E_{j}+\frac{k_{1}+1}{L}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\frac{k_{2}+1}{L}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$
Finally, plugging in the expression for $E_{j}$ and recalling that $L$ is arbitrarily large yields
$\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\precsim\frac{\sigma^{2}_{\varepsilon}}{n}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$
with probability at least $1-C_{1}\exp\left\{-C_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$ for some positive constants $C_{1},C_{2}$ as desired. $\square$
S4.2 Proof of Theorem 2
The proof of Theorem 2 occurs in three steps.
Step 1
Recall that $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]/n$. We wish to show that with high probability
$\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\leq\frac{3}{2}(\lambda+\lambda_{g}).$ (S18)
To ease the notation for this step, define the vectors
$\boldsymbol{\mathbf{r}}_{\gamma}=\frac{1}{n}\boldsymbol{\mathbf{U}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}),\quad\boldsymbol{\mathbf{r}}_{\beta}=\frac{1}{n}\boldsymbol{\mathbf{W}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})$
Let $\hat{\boldsymbol{\mathbf{\beta}}}_{h}$ indicate the $h$-th block of $\hat{\boldsymbol{\mathbf{\beta}}}$. By the KKT conditions, we know that an optimizer $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ of (15) must satisfy
$\begin{cases}\left(\boldsymbol{\mathbf{r}}_{\gamma}\right)_{\ell}=\lambda\,\operatorname{sign}(\hat{\gamma}_{\ell})&\hat{\gamma}_{\ell}\neq 0\\ \left\lvert(\boldsymbol{\mathbf{r}}_{\gamma})_{\ell}\right\rvert\leq\lambda&\hat{\gamma}_{\ell}=0\\ (\boldsymbol{\mathbf{r}}_{\beta})_{\ell}=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell})&(\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell}\neq 0\\ (\boldsymbol{\mathbf{r}}_{\beta})_{\ell}=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell})+\lambda_{g}\,\frac{(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}}{\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{h}\rVert_{2}}&(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}\neq 0,\,h\in[q]\\ \left\lvert(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}\right\rvert\leq\lambda+\lambda_{g}&(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}=0,\,h\in[q]\\ \end{cases}$
so we have for all $\ell\in[p(q+1)-1]$
$\left\lvert\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})/n\right)_{\ell}\right\rvert\leq\lambda+\lambda_{g}.$
Recalling that
$\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}=\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\beta}}=\boldsymbol{\mathbf{\varepsilon}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},$
it follows by the triangle inequality that
$\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}-\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}$ $\leq\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}-\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}$
$=\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\right\rVert_{\infty}$
$\leq\lambda+\lambda_{g}.$
[…]
$\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\leq\frac{\lambda+\lambda_{g}}{2}.$ (S19)
To see this, let $M_{\Sigma}$ denote the bound on the sub-Gaussian norm of elements of $\bm{X}$ as assumed in Assumption 3. Recall that each element of $\bm{U}$ is bounded by a constant $M$ by Assumption 1. Then the products of random variables $U_{\ell}$ and $X_{\ell}$ with the random noise $\varepsilon$ are sub-exponential with norm satisfying
$\max_{\ell}\lVert U_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq M\sigma_{\varepsilon}$
$\max_{\ell}\lVert X_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\max_{\ell}\lVert X_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq MM_{\Sigma}\sigma_{\varepsilon}.$
Define the vector $\boldsymbol{\mathbf{v}}=[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}/n$, where each element $v_{i}$ is a sum of sub-exponential functions with bounded norm. By Lemma 4 we have
$\operatorname{\operatorname{\mathbb{P}}}\left\{\lvert v_{i}\rvert>\frac{1}{2}(\lambda+\lambda_{g})\right\}\leq 2\exp\left(-c\min\left(\frac{n(\lambda+\lambda_{g})^{2}}{4M^{2}M_{\Sigma}^{2}\sigma_{\varepsilon}^{2}},\frac{n(\lambda+\lambda_{g})}{2MM_{\Sigma}\sigma_{\varepsilon}}\right)\right).$
[…]
$\lambda+\lambda_{g}\geq C\sigma_{\varepsilon}\sqrt{\frac{\log p}{n}}\iff\frac{(\lambda+\lambda_{g})^{2}n}{C^{2}\sigma_{\varepsilon}^{2}}\geq\log p,$
we have
$\frac{(\lambda+\lambda_{g})^{2}n}{C^{2}\sigma_{\varepsilon}^{2}}\cdot C_{0}^{\prime\prime}\geq C_{0}^{\prime\prime}\log p$
where $C^{\prime\prime}_{0}=C^{2}/(4M^{2}M_{\Sigma}^{2})$. Similarly, the inequality
$\frac{(\lambda+\lambda_{g})n}{C\sigma_{\varepsilon}}\geq\sqrt{n\log p}$
implies
$\frac{(\lambda+\lambda_{g})n}{C\sigma_{\varepsilon}}\cdot\sqrt{C_{0}^{\prime\prime}}\geq\sqrt{C_{0}^{\prime\prime}}\cdot\sqrt{n\log p}\geq\tilde{A}^{-1/2}\sqrt{C_{0}^{\prime\prime}}\log p$
provided that $\log p\leq\tilde{A}n$ for some constant $\tilde{A}>0$, which we assumed in the hypothesis of Theorem 2. The above shows that we can pick either argument of the minimum in Lemma 4 to develop and hence
$\operatorname{\operatorname{\mathbb{P}}}\left\{\lvert v_{i}\rvert>\frac{1}{2}(\lambda+\lambda_{g})\right\}\leq 2\exp\left(-C_{0}^{\prime}\log p\right).$
for some constant $C^{\prime}_{0}$ that increases with $C$, recalling that $C$ comes from (17). Then by the union bound we have
$\operatorname{\operatorname{\mathbb{P}}}\left(\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\geq\frac{\lambda+\lambda_{g}}{2}\right)$ $=\operatorname{\operatorname{\mathbb{P}}}\left(\max_{i}\lvert v_{i}\rvert\geq\frac{\lambda+\lambda_{g}}{2}\right)$
$\leq(p(q+1))\cdot 2\exp\left(-C^{\prime}_{0}\log p\right)$
$\leq 2\exp\left(-C^{\prime}_{0}\log p+\log p+\log(q+1)\right)$
$\leq 2\exp(-c_{3}\log p)$
for some $c_{3}>0$. In the last step we used $\log p\asymp\log q$. This shows (S19) holds with high probability.
Next we will show the cone condition
$\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}\leq 4\tau_{j}(\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1})$ (S20)
holds with high probability, where $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$ as defined in Assumption 5. Since $(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})$ is optimal, we have that
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq\frac{1}{2n}\lVert\boldsymbol{\mathbf{\varepsilon}}_{j}\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}.$
Rearranging the above leads to
$\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}+\frac{1}{n}\langle\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},\boldsymbol{\mathbf{\varepsilon}}\rangle.$ (S21)
By the same argument leading to (S19), the events
$\mathcal{A}_{1}=\left\{\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\leq\frac{\lambda}{2}\right\}\quad\text{and}\quad\mathcal{A}_{2}=\left\{\frac{1}{n}\left\lVert\boldsymbol{\mathbf{W}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\leq\frac{\lambda}{2}\right\}$
hold with probability at least $1-2\exp(-c_{4}\log p)$ and $1-2\exp(-c_{5}\log p)$ respectively for some constants $c_{4}>0$ and $c_{5}>0$.
Conditional on $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, we have by Hölder’s inequality
$\frac{1}{n}\langle\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},\boldsymbol{\mathbf{\varepsilon}}\rangle$ $=\frac{1}{n}\langle\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}},\boldsymbol{\mathbf{\varepsilon}}\rangle+\frac{1}{n}\langle\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},\boldsymbol{\mathbf{\varepsilon}}\rangle$
$\leq\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\rVert_{\infty}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\frac{1}{n}\lVert\boldsymbol{\mathbf{W}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\rVert_{\infty}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}$
$\leq\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}$
so that (S21) becomes
$\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}+\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}.$
Since $\lambda_{g}/\lambda=\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$, we have
$2\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+2\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+2\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq 2\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+2\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+2\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}+\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}.$
Now subtract by the left-hand side and add $\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}$, $\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}$, and $\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}\lVert\boldsymbol{\mathbf{\Delta}}_{-0}\rVert_{1,2}$ to find
$\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}+\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\boldsymbol{\mathbf{\Delta}}_{-0}\rVert_{1,2}$ $\leq 2\left(\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}-\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}\right)$
$\quad+2\left(\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}-\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}\right)$
$\quad+2\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\left(\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}-\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}+\lVert\boldsymbol{\mathbf{\Delta}}_{-0}\rVert_{1,2}\right),$
Then using the triangle inequality and the fact that the terms in parentheses vanish outside of the respective support sets $S_{\gamma},S_{\beta},$ and $S_{\beta,g}$, we have
$\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}\leq 4\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+4\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+4\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}.$
Since $\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}\leq\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}$, the above yields
$\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}\leq 4\tau_{j}(\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1})$
as desired, recalling again that $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$ as defined in Assumption 5
Step 2
In this step, we bound the diagonal difference $\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert$. Notice that $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ is a sum of sub-exponential random variables with sub-exponential norm bounded by $M^{2}M_{\Sigma}^{2}$, where $M$ is from Assumption 1 and $M_{\Sigma}$ from Assumption 3. We have by Lemma 4
$\operatorname{\operatorname{\mathbb{P}}}\left(\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert>\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\right)$ $\leq 2\exp\left(-cn\min\left(\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)^{2}}{4M^{2}M_{\Sigma}^{2}},\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)}{2MM_{\Sigma}}\right)\right)$
$\leq 2\exp\left(-cn\min\left(\frac{m_{0}^{2}}{4M^{2}M_{\Sigma}^{2}},\frac{m_{0}}{2MM_{\Sigma}}\right)\right)$
$\leq 2\exp(-c_{6}n)$
where we used $m_{0}\leq\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ by Assumption 2. Since
$\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert\leq\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$
implies
$\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$
which by Assumption 2 implies
$\frac{m_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2M_{0},$
it follows from the above concentration bound that
$\operatorname{\operatorname{\mathbb{P}}}\left(\frac{m_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2M_{0}\right)\geq 1-2\exp(-c_{6}n).$ (S22)
Next we bound the off-diagonal terms. Again by Lemma 4
$\operatorname{\operatorname{\mathbb{P}}}\left(\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)\rvert\geq 2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)\right)$
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)^{2}}{M^{2}M_{\Sigma}^{2}},\frac{2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)}{MM_{\Sigma}}\right)\right)$
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4M_{0}^{2}}{M^{2}M_{\Sigma}^{2}},\frac{4M_{0}}{MM_{\Sigma}}\right)\right)$
$\qquad\leq 2\exp(-c_{7}n).$
By Assumption 5, we have
$-2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq 2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$
$\implies-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq 3\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$
$\implies-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}$
Combining this with the above concentration bound yields
$\operatorname{\operatorname{\mathbb{P}}}\left(-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\leq\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\right)\geq 1-2\exp(-c_{7}n).$ (S23)
Step 3
Define the index set $\tilde{S}=S_{\gamma}\cup\{q+i\mid i\in S_{\beta}\}$. The set $\{q+i\mid i\in S_{\beta}\}$ indexes the columns of the submatrix $\boldsymbol{\mathbf{W}}$ of the matrix $[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ corresponding to the support set of $\boldsymbol{\mathbf{\beta}}$. With slight abuse of notation, for a vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}$ let $\boldsymbol{\mathbf{v}}_{\tilde{S}}\in\mathbb{R}^{p(q+1)-1}$ be the vector that equals $\boldsymbol{\mathbf{v}}$ on the set $\tilde{S}$ and is zero on $\tilde{S}^{c}$, so that it has $s_{\gamma}+s_{\beta}$ nonzero elements. Define $\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}$ analogously.
We will show that conditional on $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and the bounds
$\begin{split}\frac{m_{0}}{2}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2M_{0},\\ -\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})},\end{split}$ (S24)
we have
$\inf_{\boldsymbol{\mathbf{v}}\in\mathcal{V}}\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}\rVert_{2}}{\sqrt{n}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}}\geq\sqrt{\frac{m_{0}}{2}-\frac{1}{c_{0}}}>0.$
where $\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}\mid\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}\leq 4\tau_{j}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\}$
First we have that
$\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$ $=\frac{\boldsymbol{\mathbf{v}}_{\tilde{S}}^{\top}\operatorname{diag}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\boldsymbol{\mathbf{v}}_{\tilde{S}}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}+\frac{\boldsymbol{\mathbf{v}}_{\tilde{S}}^{\top}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\operatorname{diag}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}))\boldsymbol{\mathbf{v}}_{\tilde{S}}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}.$
Then by $\lVert\boldsymbol{\mathbf{a}}+\boldsymbol{\mathbf{b}}\rVert_{2}^{2}=\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}+2\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{b}}+\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}$ we have
$\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}\rVert_{2}^{2}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$ $\geq\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}+2\frac{\boldsymbol{\mathbf{v}}_{\tilde{S}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{2}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{1+8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{m_{0}}{2}-\frac{(1+8\tau_{j})(s_{\beta}+s_{\gamma})}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}=\frac{m_{0}}{2}-\frac{1}{c_{0}}>0.$
In the last step we used that $c_{0}>2/m_{0}$ in Assumption 5. We have shown that
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\geq\left(\frac{m_{0}}{2}-\frac{1}{c_{0}}\right)\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}$ (S25)
Final step
It is true that for $\ell\in[p(q+1)-1]$ we have
$\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}=\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}+\sum_{k\neq\ell}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{k}.$
Then by (S24) and the triangle inequality we have
$\left\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}\right\rvert-\left\lvert\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert$ $\leq\left\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}-\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert$
$\leq\left\lvert\sum_{k\neq\ell}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{k}\right\rvert$
$\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\sum_{k\neq\ell}\left\lvert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{k}\right\rvert.$
Rearranging terms and applying (S24) yields
$\frac{m_{0}}{2}\left\lvert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}\right\rvert\leq\left\lvert\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert+\frac{3}{c_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$
Since this holds for all $\ell\in[p(q+1)-1]$, we have shown that
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\leq\frac{2}{m_{0}}\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}+\frac{6}{c_{0}m_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$ (S26)
Combining (S18) and (S20) we have
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$ $\leq\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}$
$\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{1}$
$\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\sqrt{s_{\gamma}+s_{\beta}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
and combining this with (S25) gives us
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\left(\frac{2c_{0}}{c_{0}m_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
and therefore
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})\left(\frac{c_{0}}{c_{0}m_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}.$
On the other hand, by (S20) we have
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}\leq(1+4\tau_{j})\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{1}\leq(1+4\tau_{j})\sqrt{s_{\beta}+s_{\gamma}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
so combining this with the above yields
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})^{2}\left(\frac{c_{0}}{c_{0}m_{0}-2}\right)(s_{\gamma}+s_{\beta})$
Finally, plugging this into (S26) and with (S18) yields
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}$ $\leq\frac{3}{m_{0}}(\lambda+\lambda_{g})+\frac{18(\lambda+\lambda_{g})(1+4\tau_{j})^{2}}{c_{0}m_{0}(1+8\tau_{j})}\left(\frac{c_{0}}{c_{0}m_{0}-2}\right)$
$=\frac{3}{m_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}m_{0}-2)}\right)$
as desired. $\square$