### v2

# Convex estimation of Gaussian graphical regression models with covariates

Abstract
Gaussian graphical models (GGMs) are widely used to recover the conditional independence structure among random variables. Recent work has sought to incorporate auxiliary covariates to improve estimation, particularly in applications such as co-expression quantitative trait locus (eQTL) studies, where both gene expression levels and their conditional dependence structure may be influenced by genetic variants. Existing approaches to covariate-adjusted GGMs either restrict covariate effects to the mean structure or lead to nonconvex formulations when jointly estimating the mean and precision matrix. In this paper, we propose a convex framework that simultaneously estimates the covariate-adjusted mean and precision matrix via a natural parametrization of the multivariate Gaussian likelihood. The resulting formulation enables joint convex optimization and yields improved theoretical guarantees under high-dimensional scaling, where the sparsity and dimension of covariates grow with the sample size. We support our theoretical findings with numerical simulations and demonstrate the practical utility of the proposed method through a reanalysis of an eQTL study of glioblastoma multiforme (GBM), an aggressive form of brain cancer.
1 Introduction
Graphical models use graphs to represent conditional independence relationships among components of a random vector $\bm{X}=(X_{1},\dots,X_{p})$. In the Gaussian case, this leads to Gaussian graphical models (GGMs), where conditional independences correspond to zeros in the precision matrix $\boldsymbol{\mathbf{\Omega}}=\boldsymbol{\mathbf{\Sigma}}^{-1}$. Specifically, for a Gaussian random vector $\bm{X}$ and $i\neq j$, the variables $X_{i}$ and $X_{j}$ are conditionally independent given $\bm{X}_{\{1,\dots,p\}\setminus\{i,j\}}$ if and only if $\Omega_{ij}=0$ [lauritzen]. There has been extensive work on the estimation of GGMs in high-dimensional settings [yuanlin, friedman-glasso, meinshausen], where sparsity is typically induced on the precision matrix to improve interpretability and estimation accuracy.
A major application of GGMs arises in genomics, where the goal is to infer gene interaction networks from gene expression data. Such problems are inherently high-dimensional, with the number of genes often exceeding the number of samples [schafer2005]. A prominent example is provided by co-expression quantitative trait locus (eQTL) studies, in which gene expression measurements are analyzed jointly with external genetic markers. These markers may confound gene expression levels, either through local genetic effects or through associations with unmeasured factors influencing gene expression. This setting motivates the development of Gaussian graphical models that incorporate covariate information, enabling adjustment for such confounding in network estimation.
Specifically, let $\bm{X}\in\mathbb{R}^{p}$ be the vector of responses and $\bm{U}\in\mathbb{R}^{q}$ the associated vector of covariates. We consider the general model
$\bm{X}\>|\>\bm{U}=\boldsymbol{\mathbf{u}}\sim N(\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}}),\boldsymbol{\mathbf{\Omega}}^{-1}(\boldsymbol{\mathbf{u}})),$ (1)
which allows both the mean vector and the precision matrix of $\bm{X}$ to depend on a realization of covariates $\bm{U}=\boldsymbol{\mathbf{u}}$. A commonly used specification assumes that only the mean depends on $\bm{U}$, while the precision matrix remains constant. This specification is also known as the conditional Gaussian graphical model [yin-li]:
$\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Omega}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}.$ (2)
Estimation of the unknown parameters in (2) is challenging, as the negative log-likelihood associated with (1) is not jointly convex in $(\boldsymbol{\mathbf{\Gamma}},\boldsymbol{\mathbf{B}}_{0})$. Existing approaches typically rely on alternating optimization [rothman, yin-li, chen18] or two-stage procedures [cai13, chen2016]. Despite this nonconvexity, minimax-optimal rates for joint estimation under (1) with (2) have been established [chen18, lv]. Several works have sought to convexify the problem. For instance, wang_joint_2015 propose a convex formulation based on an initial estimate of $\boldsymbol{\mathbf{\Gamma}}$, while zhu_convex_2020 introduce a reparameterization that yields a jointly convex negative log-likelihood.
Model (1) with specification (2) can also be viewed as a multivariate response regression model, where the primary goal is to improve estimation of $\boldsymbol{\mathbf{\Gamma}}$ by leveraging structural assumptions on $\boldsymbol{\mathbf{B}}_{0}$. In contrast, the covariate-adjusted GGM framework focuses on recovering the sparsity pattern of the precision matrix while accounting for covariate effects through $\bm{U}$. Several methods in this setting incorporate structure-inducing penalties into the likelihood of (1) associated with the specification (2); see liu_estimation_2025 for a recent review highlighting connections and distinctions between these closely related lines of work.
While substantial progress has been made in modeling heterogeneous means, relatively little work allows the graph structure itself to vary with covariates; most existing approaches instead assume a common precision matrix $\boldsymbol{\mathbf{B}}_{0}$ across subjects. However, subject-specific graph structures are plausible in applications such as eQTL studies. For example, gene C may mediate the co-expression of genes A and B only in the presence of a single-nucleotide polymorphism (SNP) local to gene C. In such cases, genes A and B may be conditionally independent given the rest of the network unless a genetic variant is present near gene C [fehrmann, kolberg, rockman].
A small number of works have considered models in which both the mean and precision depend on covariates. Bayesian approaches, for example, allow subject-specific precision matrices by imposing regression or clustering structures on $\boldsymbol{\mathbf{\Omega}}$ [ni_bayesian_2022, niu_covariate-assisted_2024, wang_bayesian_2022, zeng_bayesian_2025]. However, these methods do not directly address joint estimation of both the mean and precision as functions of covariates within a unified framework.
This work is closely related to zhangli, which extends (2) to allow both the mean and the network structure of the responses to depend on covariates. Specifically, their model specifies (1) by taking
$\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Omega}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}$ (3)
where $\boldsymbol{\mathbf{B}}_{h}\in\mathbb{R}^{p\times p},h=0,1\dotsc,q$ are sparse, symmetric matrices. As in conditional GGMs, joint estimation of the parameters $(\boldsymbol{\mathbf{\Gamma}},\boldsymbol{\mathbf{B}}_{0},\boldsymbol{\mathbf{B}}_{1},\dotsc,\boldsymbol{\mathbf{B}}_{q})$ under (3) leads to a nonconvex objective. To address this, zhangli adopt a two-stage estimation procedure. Building on this framework, zhang-multi-task-2025 propose a multi-task estimator with a cross-task group penalty, while meng_statistical_2025 develop debiased nodewise estimators for inference on edge–covariate effects.
In this work, we develop Covariate-adjusted Sparse Precision with Natural Estimation (cspine), a method for estimating covariate-adjusted GGMs with heterogeneous network effects based on a jointly convex formulation of (1). Like zhangli, our method uses nodewise regression [meinshausen] to estimate the covariate-adjusted graph structure. However, by adopting a natural parametrization of the Gaussian likelihood, we obtain a formulation in which each nodewise regression reduces to a convex optimization problem, thereby avoiding the nonconvexity inherent in the existing approaches.
Our main contributions and the organization of the paper are as follows. In Section 2, we review the framework of zhangli and introduce the proposed natural parametrization, under which the likelihood in (1) is jointly convex in all unknown parameters. Building on this formulation, Section 3 presents the proposed model and estimation procedure with sparsity-inducing penalties. Section 4 establishes theoretical guarantees for the proposed estimator, showing that the convex formulation leads to improved scaling of the ambient dimensions $p$ and $q$ relative to the sample size $n$ under assumptions comparable to those in zhangli. These theoretical advantages are further supported by extensive simulations in Section 5. Finally, in Section 6, we illustrate the practical utility of the proposed approach through an analysis of glioblastoma multiforme (GBM) gene expression data. Proofs of all theoretical results, along with additional numerical results, are provided in the Appendix.
Throughout this paper, we adopt the following notational conventions unless otherwise specified. Scalars are denoted by regular (non-bold) letters, vectors by bold lowercase letters, and matrices by bold uppercase letters. For any positive integer $m$, let $m]=\{1,\ldots,m\}$, and for $j\in[m]$, let j-j denote the index set $m]\setminus{j}$. For a matrix $\boldsymbol{\mathbf{M}}$ and index sets $\mathcal{I}$ and $\mathcal{J}$, we write $\boldsymbol{\mathbf{M}}_{\mathcal{I},}$, $\boldsymbol{\mathbf{M}}_{\mathcal{J}}$, and $\boldsymbol{\mathbf{M}}_{\mathcal{I},\mathcal{J}}$ for the submatrices of $\boldsymbol{\mathbf{M}}$ consisting of rows indexed by $\mathcal{I}$, columns indexed by $\mathcal{J}$, and both rows indexed by $\mathcal{I}$ and columns indexed by $\mathcal{J}$, respectively.
2 Model formulation and natural parametrization
In this section, we first review the model formulation of zhangli for estimating covariate-adjusted GGMs in (1) via a nodewise regression approach, and discuss the challenges of their two-stage estimation procedure arising from the nonconvexity of the formulation. We then introduce a convex formulation of (1) based on a natural parametrization.
2.1 Nodewise regression in covariate-adjusted GGMs
Suppressing the dependence of $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\mu}}$ on the covariates $\bm{U}$ for notational simplicity, the Gaussian model in (1) implies that, for each $j\in[p]$,
$X_{j}-\mu_{j}\>|\>\bm{X}_{-j},\bm{U}\sim N\left(\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j}),\,\Sigma_{jj}-\boldsymbol{\mathbf{\Sigma}}_{j,-j}\boldsymbol{\mathbf{\Sigma}}^{-1}_{-j,-j}\boldsymbol{\mathbf{\Sigma}}_{-j,j}\right).$ (4)
By the block matrix inversion formula, (4) is equivalent to the linear model with Gaussian errors
$X_{j}=\mu_{j}-\Omega^{-1}_{jj}\boldsymbol{\mathbf{\Omega}}_{j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j})+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}}),\quad\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj},$ (5)
where the expression is understood conditional on $(\bm{X}_{-j},\bm{U})$. Based on this representation, nodewise regression methods for estimating GGMs [meinshausen] can be naturally extended to incorporate covariate adjustments, where the regression coefficients and error variance are used to recover the $j$-th column of the precision matrix $\boldsymbol{\mathbf{\Omega}}$.
2.2 Non-convexity of the nodewise regression under (3)
Write $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}\in\mathbb{R}^{p\times q}$. Under the specification (3) of zhangli, the nodewise regression model (5) becomes
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\eta_{jk0}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\eta_{jkh}u_{h}(X_{k}-\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{k})+\varepsilon_{j},$ (6)
[…]
To address this difficulty, zhangli separate the estimation of $\boldsymbol{\mathbf{\gamma}}_{j}$ from that of $\eta_{jkh}$. Specifically, if $\boldsymbol{\mathbf{\Gamma}}$ were known, the responses may first be centered via $Z_{j}=X_{j}-\boldsymbol{\mathbf{u}}^{T}\boldsymbol{\mathbf{\gamma}}_{j}$, reducing the model to the oracle form
$Z_{j}=\sum_{k\neq j}^{p}\eta_{jk0}Z_{k}+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\eta_{jkh}u_{h}Z_{k}+\varepsilon_{j},$ (7)
for which the least squares criterion is convex in $\eta$’s.
In practice, however, $\boldsymbol{\mathbf{\Gamma}}$ is not known, and a two-stage procedure is employed. In the first stage, each $\boldsymbol{\mathbf{\gamma}}_{j}$ is estimated under a simplified working model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\varepsilon_{j}.$
Comparing with the full model (6), this working model imposes a strong assumption on the specification (3), effectively treating the precision matrix $\boldsymbol{\mathbf{B}}_{0}$ as diagonal and all other matrices $\boldsymbol{\mathbf{B}}_{h}$ as zero matrices, thereby assuming conditional independence among all pairs of responses, which is furthermore not dependent on covariates. As a result, the first-stage estimator does not account for the network structure and its potential dependence on covariates, and thus may incur non-negligible bias when such dependencies are present.
Given the resulting estimate of $\boldsymbol{\mathbf{\Gamma}}$, the second stage proceeds by forming centered responses $Z_{j}$ and estimating the coefficients $\eta_{jkh}$ using the oracle model (7). This stage treats the estimate of $\boldsymbol{\mathbf{\Gamma}}$ as fixed, thereby inheriting any estimation error from the first stage.
The two-stage procedure introduces additional estimation error, as inaccuracies in estimating $\boldsymbol{\mathbf{\gamma}}_{j}$ propagate to the estimation of $\eta_{jkh}$. As shown in the theoretical analysis of zhangli, the concentration guarantees are weakened relative to the oracle setting in which $\boldsymbol{\mathbf{\Gamma}}$ is known, since the overall error probability depends on the tuning parameters from the first stage. We revisit these limitations in Section 4.1.
2.3 A convex formulation via natural parametrization
[…]
$\boldsymbol{\mathbf{\theta}}=\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})^{-1}\boldsymbol{\mathbf{\Omega}}\boldsymbol{\mathbf{\mu}}\quad\text{and}\quad\boldsymbol{\mathbf{\Theta}}=-\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})^{-1}\boldsymbol{\mathbf{\Omega}},$ (8)
where $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})$ is the $p\times p$ diagonal matrix formed from $\boldsymbol{\mathbf{\Omega}}$. Under this parametrization, the nodewise regression model in (5) can be written as
$X_{j}=\mu_{j}-\Omega^{-1}_{jj}\boldsymbol{\mathbf{\Omega}}_{j,-j}(\bm{X}_{-j}-\boldsymbol{\mathbf{\mu}}_{-j})+\varepsilon_{j}=\theta_{j}+\boldsymbol{\mathbf{\Theta}}_{j,-j}\bm{X}_{-j}+\varepsilon_{j},$
which yields a least squares criterion that is jointly convex in $\theta_{j}$ and $\boldsymbol{\mathbf{\Theta}}_{j,-j}$.
In contrast to (3), we model covariate dependence through $(\boldsymbol{\mathbf{\theta}},\boldsymbol{\mathbf{\Theta}})$ rather than $(\boldsymbol{\mathbf{\mu}},\boldsymbol{\mathbf{\Omega}})$ by specifying
$\boldsymbol{\mathbf{\theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}},\quad\boldsymbol{\mathbf{\Theta}}(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h},$ (9)
where $\boldsymbol{\mathbf{B}}_{0}\in\mathbb{R}^{p\times p}$ represents the baseline graph structure and $\boldsymbol{\mathbf{B}}_{h}\in\mathbb{R}^{p\times p}$ captures the effect of covariate $u_{h}$ on the graph structure. Writing $\boldsymbol{\mathbf{\Gamma}}=(\boldsymbol{\mathbf{\gamma}}_{1},\dotsc,\boldsymbol{\mathbf{\gamma}}_{p})^{\top}$ and letting $\beta_{jkh}=[\boldsymbol{\mathbf{B}}_{h}]_{jk}$ for $k\in[p]\setminus\{j\}$ and $h\in\{0\}\cup[q]$, this specification leads to the following nodewise regression model
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\sum_{k\neq j}^{p}\beta_{jk0}X_{k}+\sum_{k\neq j}^{p}\sum_{h=1}^{q}\beta_{jkh}u_{h}X_{k}+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}}),\quad\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj}$ (10)
for $j\in[p]$. That is, $X_{j}$ is regressed on the covariates $\boldsymbol{\mathbf{u}}$, all other responses $X_{k}$ for $k\neq j$, and the pairwise interactions between the covariates and the other responses. Remarkably, the least squares criterion associated with (10) is jointly convex in $\boldsymbol{\mathbf{\gamma}}_{j}$ and $\beta_{jkh}$, enabling simultaneous estimation of all parameters and avoiding the nonconvexity inherent in the formulation of zhangli.
The parameterization (8) implicitly imposes that $\Theta_{jj}=-1$ for all $j\in[p]$. Under the specification (9), this is enforced by setting $\boldsymbol{\mathbf{B}}_{0}]_{jj}=-1$ and $\boldsymbol{\mathbf{B}}_{h}]_{jj}=0$ for all $h\in[q]$ and $j\in[p]$. Consequently, the residual variance $\sigma^{2}_{\varepsilon_{j}}=1/\Omega_{jj}$, i.e., the conditional variance, does not depend on covariates, an assumption also adopted by zhangli. Moreover, $\boldsymbol{\mathbf{\Theta}}$ and $\boldsymbol{\mathbf{\Omega}}$ share the same sparsity structure, so that estimating the sparsity pattern of $\boldsymbol{\mathbf{\Theta}}$ is equivalent to that of $\boldsymbol{\mathbf{\Omega}}$. Accordingly, we focus on recovering the sparsity pattern of $\boldsymbol{\mathbf{\Theta}}$ in what follows.
Finally, we note that the proposed parametrization is closely related to, yet distinct from, existing convex formulations. The transformation in (8) is conceptually related to the natural parametrization in yu2019 for univariate-response regression and its multivariate extension in zhu_convex_2020. However, the scaling by $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}})^{-1}$ in (8) is crucial for ensuring that the nodewise regression model in (10) remains jointly convex in the presence of covariate adjustments. This feature distinguishes our approach from the reparametrizations in yu2019 and zhu_convex_2020, which do not accommodate covariate-dependent variance or covariance structures.
3 Estimation
For each $j\in[p]$, define the concatenated parameter vector $\boldsymbol{\mathbf{\beta}}_{j}=(\boldsymbol{\mathbf{\beta}}^{\top}_{j,0},\boldsymbol{\mathbf{\beta}}^{\top}_{j,1},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q}^{\top})^{\top}\in\mathbb{R}^{(p-1)(q+1)}$, where the $h$-th block collects the parameters $\boldsymbol{\mathbf{\beta}}_{j,h}=(\beta_{j1h},\dotsc,\beta_{j,j-1,h},\beta_{j,j+1,h},\dotsc,\beta_{jph})^{\top}\in\mathbb{R}^{p-1}$. Under this notation, the nodewise regression model (10) can be written compactly as
$X_{j}=\boldsymbol{\mathbf{u}}^{\top}\boldsymbol{\mathbf{\gamma}}_{j}+\bm{X}_{-j}^{\top}\boldsymbol{\mathbf{\beta}}_{j,0}+\sum_{h=1}^{q}(u_{h}\bm{X}_{-j})^{\top}\boldsymbol{\mathbf{\beta}}_{j,h}+\varepsilon_{j},\quad\varepsilon_{j}\sim N(0,\sigma^{2}_{\varepsilon_{j}}).$ (11)
Figure 1 illustrates how each block $\boldsymbol{\mathbf{\beta}}_{j,h}^{\top}$ corresponds to the $j$-th row (and, by symmetry, the $j$-th column) of $\boldsymbol{\mathbf{\Theta}}$ in (9).
Suppose we collect $n$ independent observations $\{(\boldsymbol{\mathbf{x}}^{(i)},\boldsymbol{\mathbf{u}}^{(i)})\}_{i=1}^{n}$ of responses and covariates from the joint distribution in (1). Let $\boldsymbol{\mathbf{U}}=\left[\boldsymbol{\mathbf{u}}^{(1)},\dotsc,\boldsymbol{\mathbf{u}}^{(n)}\right]^{\top}\in\mathbb{R}^{n\times q}$ and $\boldsymbol{\mathbf{X}}=\left[\boldsymbol{\mathbf{x}}^{(1)},\dotsc,\boldsymbol{\mathbf{x}}^{(n)}\right]^{\top}\in\mathbb{R}^{n\times p}$ denote the covariate and response matrices, respectively. Let $\boldsymbol{\mathbf{u}}_{h}\in\mathbb{R}^{n}$ and $\boldsymbol{\mathbf{x}}_{k}\in\mathbb{R}^{n}$ denote the $h$-th and $k$-th columns of $\boldsymbol{\mathbf{U}}$ and $\boldsymbol{\mathbf{X}}$. For $j\in[p]$, define the design matrix
$\boldsymbol{\mathbf{W}}_{-j}=\left[\boldsymbol{\mathbf{W}}_{-j,0},\ \boldsymbol{\mathbf{W}}_{-j,1},\ \dotsc,\ \boldsymbol{\mathbf{W}}_{-j,q}\right]\in\mathbb{R}^{n\times(p-1)(q+1)},$
where
$\begin{split}\boldsymbol{\mathbf{W}}_{-j,0}&=\left[\boldsymbol{\mathbf{x}}_{1},\dotsc,\boldsymbol{\mathbf{x}}_{j-1},\boldsymbol{\mathbf{x}}_{j+1},\dotsc,\boldsymbol{\mathbf{x}}_{p}\right]\in\mathbb{R}^{n\times(p-1)},\\ \boldsymbol{\mathbf{W}}_{-j,h}&=\left[\boldsymbol{\mathbf{x}}_{1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{j-1}\odot\boldsymbol{\mathbf{u}}_{h},\ \boldsymbol{\mathbf{x}}_{j+1}\odot\boldsymbol{\mathbf{u}}_{h},\ \dotsc,\ \boldsymbol{\mathbf{x}}_{p}\odot\boldsymbol{\mathbf{u}}_{h}\right]\in\mathbb{R}^{n\times(p-1)}\;\text{for $h\in[q]$},\end{split}$ (12)
and $\odot$ denotes the elementwise multiplication.
We consider estimating $\boldsymbol{\mathbf{\gamma}}_{j}$ and $\boldsymbol{\mathbf{\beta}}_{j}$ in (11) by solving the following penalized least squares problem
$\operatorname*{minimize}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+g_{j}(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j}),$ (13)
where $g_{j}$ is a convex penalty function. In this paper, we focus on penalties that induce sparsity in $\boldsymbol{\mathbf{\beta}}_{j}$, promoting interpretable graph estimates. By construction, sparsity in $\boldsymbol{\mathbf{\beta}}_{j}$ corresponds to sparsity in the $j$-th row (and, by symmetry, the $j$-th column) of $\boldsymbol{\mathbf{\Theta}}$, and hence in $\boldsymbol{\mathbf{\Omega}}$ via (8).
Figure 1: Decomposition of $\boldsymbol{\mathbf{\Theta}}$ into components $\boldsymbol{\mathbf{B}}_{h}$ according to (9). The block $\boldsymbol{\mathbf{\beta}}_{j,h}$ corresponds to the effects of covariate $u_{h}$ on the partial correlations of response $X_{j}$ while $\boldsymbol{\mathbf{\beta}}_{j,0}$ describes the population effect.
3.1 Sparse-group penalized estimation
Motivated by applications to eQTL studies, we consider a structured sparsity-inducing penalty for $g_{j}$ in (14). Our goal is to capture both elementwise and groupwise sparsity in the coefficients $\boldsymbol{\mathbf{\beta}}_{j}$. Elementwise sparsity within each group $\boldsymbol{\mathbf{\beta}}_{j,h}$ allows for the possibility that a covariate $u_{h}$ affects the conditional dependence between $X_{j}$ and only a subset of the remaining $\{X_{k}:k\neq j\}$. In contrast, groupwise sparsity, where $\boldsymbol{\mathbf{\beta}}_{j,h}=\boldsymbol{\mathbf{0}}$, indicates that the covariate $u_{h}$ has no effect on any conditional dependence structure involving $X_{j}$.
To this end, we consider the convex optimization problem
$\operatorname*{minimize}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}_{j}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}$ (14)
where $\lambda\geq 0$ and $\lambda_{g}\geq 0$ are tuning parameters, and $\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}=\sum_{h=1}^{q}\lVert\boldsymbol{\mathbf{\beta}}_{j,h}\rVert_{2}$ denotes the group lasso penalty [yuan2006model]. Together, the $\ell_{1}$ and group penalties on $\boldsymbol{\mathbf{\beta}}_{j}$ form a sparse-group lasso penalty [simon]. In contrast, only an $\ell_{1}$ penalty is applied to $\boldsymbol{\mathbf{\beta}}_{j,0}$ and $\boldsymbol{\mathbf{\gamma}}_{j}$, allowing for unrestricted elementwise sparsity in the baseline precision matrix $\boldsymbol{\mathbf{B}}_{0}$ and in the covariate effects on the mean vector.
To estimate the diagonal entries of $\boldsymbol{\mathbf{\Omega}}$, an estimate of the noise variance $\sigma^{2}_{\varepsilon_{j}}$ is required. While several estimators are available in the literature on error variance estimation in high-dimensional linear regression [reid, yu2019], we adopt
$\hat{\sigma}^{2}_{\varepsilon_{j}}=\frac{\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\hat{\boldsymbol{\mathbf{\beta}}}_{j}\rVert_{2}^{2}}{n-\widehat{\mathrm{df}}}$, (15)
where $\widehat{\mathrm{df}}$ denotes an estimate of the degrees of freedom of (14) as in li_highly_2018. Alternative approaches for estimating the degrees of freedom, or more generally the noise variance in (11), may also be employed.
3.2 Implementation details
We solve (14) using a blockwise coordinate descent algorithm, iteratively minimizing the objective with respect to each block in $\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j,0},\dotsc,\boldsymbol{\mathbf{\beta}}_{j,q}$} while holding the others fixed. Each block update can be efficiently computed via a majorized proximal gradient method [simon, liang_sparsegl_2024]. Since (14) is convex and separable across blocks, the algorithm is guaranteed to converge to the optimal solution [Tseng].
The penalized nodewise regression problem (14) involves two tuning parameters: $\lambda$ controlling elementwise sparsity and $\lambda_{g}$ controlling groupwise sparsity. We select $(\lambda,\lambda_{g})$ via $k$-fold cross-validation. Equivalently, we reparametrize the penalty using a global sparsity penalty $\lambda_{0}>0$ and a mixing parameter $\alpha_{s}\in[0,1]$, so that the penalty in (14) becomes
$\alpha_{s}\lambda_{0}(\lVert\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{1}+\lVert\boldsymbol{\mathbf{\beta}}_{j}\rVert_{1})+(1-\alpha_{s})\lambda_{0}\lVert\boldsymbol{\mathbf{\beta}}_{j,-0}\rVert_{1,2}.$ (16)
For each fixed $\alpha_{s}$, we compute solution along a regularization path in $\lambda_{0}$.
As in standard nodewise regression approaches to GGMs, each row of the precision matrix is estimated separately, and the resulting estimate is not necessarily symmetric [meinshausen]. In our setting, each nodewise regression in (14) estimates one row of $\boldsymbol{\mathbf{\Theta}}$, and asymmetry arises due to both independent row estimation and the different row scaling factors in (8). To enforce symmetry, we apply a post hoc symmetrization step after performing all nodewise regressions. Specifically, with a solution $\hat{\boldsymbol{\mathbf{\beta}}}_{j}$ to (14) and the corresponding error variance estimate in (15), set $\tilde{\boldsymbol{\mathbf{\beta}}}_{j}=-\hat{\boldsymbol{\mathbf{\beta}}}_{j}/\hat{\sigma}^{2}_{\varepsilon_{j}}$ and construct symmetric matrices $\tilde{\boldsymbol{\mathbf{B}}}_{h}\in\mathbb{R}^{p\times p}$ for all $h=0,1,\dotsc,q$ with entries
$\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathds{1}\left[\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\right]+\tilde{\beta}_{kjh}\mathds{1}\left[\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\right$],
where the expression $\mathds{1}\left[P\right]$ is the indicator function of the argument $P$: it equals to $1$ if $P$ is true and 0 otherwise. This corresponds to the “and-rule” symmetrization, under which $\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}$ is nonzero only if both $\tilde{\beta}_{jkh}$ and $\tilde{\beta}_{kjh}$ are nonzero. An alternative is the “or-rule,” defined by
$\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathds{1}{\left[\lvert\tilde{\beta}_{jkh}\rvert\geq\lvert\tilde{\beta}_{kjh}\rvert\right]}+\tilde{\beta}_{kjh}\mathds{1}{\left[\lvert\tilde{\beta}_{jkh}\rvert\leq\lvert\tilde{\beta}_{kjh}\rvert\right]}$,
which retains an edge if either coefficient is nonzero. We adopt the more conservative “and-rule”, noting that both approaches are asymptotically equivalent [meinshausen] and have been widely used in the GGM literature [cai2011constrained, liu2012tiger, Li_2021, tran2022completely, cai13, zhangli].
Finally, the mean vector and precision matrix are recovered via (8) as
$\hat{\boldsymbol{\mathbf{\Omega}}}(\boldsymbol{\mathbf{u}}^{(i)})$ $=\tilde{\boldsymbol{\mathbf{B}}}_{0}+\sum_{h=1}^{q}\tilde{\boldsymbol{\mathbf{B}}}_{h}u^{(i)}_{h}\quad\text{and}\quad\hat{\boldsymbol{\mathbf{\mu}}}(\boldsymbol{\mathbf{u}}^{(i)})=(\hat{\boldsymbol{\mathbf{\Omega}}}(\boldsymbol{\mathbf{u}}^{(i)}))^{-1}\operatorname{diag}(\hat{\boldsymbol{\mathbf{\Omega}}})\hat{\boldsymbol{\mathbf{\Gamma}}}\boldsymbol{\mathbf{u}}^{(i)}.$ (17)
The full procedure is summarized in Algorithm 1 in Section S1.
4 Theoretical properties
In this section, we present theoretical guarantees for the estimator defined in (14), which characterize the accuracy of the recovered graph structure and its dependence on covariates. We begin by introducing notation and stating assumptions on the data-generating process. We then present our main results, including $\ell_{2}$ estimation error bounds and support recovery guarantees, followed by a discussion comparing our results with those of zhangli.
For two sequences of real numbers $a_{n}$ and $b_{n}$, we write $a_{n}\precsim b_{n}$ if $a_{n}=O(b_{n})$, i.e. there exists constants $C>0$ and $N>0$ so that $a_{n}<Cb_{n}$ for all $n\geq N$. If $a_{n}\precsim b_{n}$ and $b_{n}\precsim a_{n}$, we write $a_{n}\asymp b_{n}$. We write $a_{n}=o(b_{n})$ if $\lim_{n\rightarrow\infty}a_{n}/b_{n}=0$. Our analysis holds for any $j\in[p]$. For notational simplicity, we suppress the subscript $j$ when referring to $\boldsymbol{\mathbf{\gamma}}_{j}$, $\boldsymbol{\mathbf{\beta}}_{j}$, $\boldsymbol{\mathbf{\varepsilon}}_{j}$ and $\boldsymbol{\mathbf{W}}_{-j}$ in this section.
[…]
Assumption 1.
Suppose $\{\boldsymbol{\mathbf{u}}^{(i)}\}_{i=1}^{n}$ are i.i.d. mean zero random vectors with uniformly bounded entries: there exists a constant $M_{u}>0$ such that $\lvert u^{(i)}_{h}\rvert\leq M_{u}$ for all $h\in[q]$ and $i\in[n]$.
Assumption 2.
The responses $\boldsymbol{\mathbf{x}}^{(i)}$ are elementwise sub-Gaussian with bounded sub-Gaussian norm: there exists a constant $K_{X}>0$ such that $\lVert x^{(i)}_{j}\rVert_{\psi_{2}}\leq K_{X}$ for all $j\in[p]$ and $i\in[n]$.
Let $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\in\mathbb{R}^{n\times(p(q+1)-1)}$ denote the augmented design matrix in (14). Define the population Gram matrix $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]/n\right)$.
Assumption 3.
There exist constants $0<\phi_{0}^{\ast}\leq\phi_{1}^{\ast}<\infty$ such that
$\phi_{0}^{\ast}\leq\lambda_{\mathrm{min}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq\lambda_{\mathrm{max}}(\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\leq\phi_{1}^{\ast}.$
Assumption 4.
The sparsity levels satisfy
$(s_{\gamma}+s_{\beta})\log(pq)=o(\sqrt{n}/\log n).$
Assumption 12 controls the tail behavior of the covariates, responses, and their pairwise interactions, thus facilitating their concentration behavior. Similar assumptions are also imposed in zhangli. Assumption 3 ensures the augmented design matrix is well-conditioned, preventing excessive collinearity among predictors in the nodewise regression (14). Assumption 3 is closely related to the bounded eigenvalue assumptions on $\operatorname{Cov}(\bm{U})$ and $\operatorname{Cov}(\bm{X}\>|\>\bm{U})$ in zhangli. In Section S5, we provide interpretable sufficient conditions under which Assumption 3 holds. Finally, Assumption 4 governs the scaling of sparsity levels with the sample size. Combined with Theorem 4.1 of kuchi, it allows us to control sparse deviations between sample and population Gram matrices (see Lemmas 5 and 7 in Appendix S4). A sufficient condition for Assumption 4 is $\log(pq)=O(n^{1/6})$ and $s_{\gamma}+s_{\beta}=o(n^{1/6})$, as assumed in zhangli.
[…]
Theorem 1.
Under Assumptions 1-4, with $(\hat{s}_{\gamma}^{\mathrm{max}}+\hat{s}_{\beta}^{\mathrm{max}})\log(pq)=O(\sqrt{n})$, and with
$\lambda=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\gamma}+s_{\beta}}+\frac{s_{\beta,g}}{s_{\gamma}+s_{\beta}}\log(eq/s_{\beta,g})\right)^{1/2},\quad\lambda_{g}=\lambda\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}},$ (18)
we have that
$\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{2}^{2}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}\precsim\frac{\sigma_{\varepsilon}^{2}}{n}\left\{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right\}$
holds with probability at least $1-C_{1}\exp\left[-C_{2}\{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\}\right]$ where $C$, $C_{1}$, and $C_{2}$ are positive constants.
We next establish support recovery guarantees, under an additional mutual incoherence condition. Denote the $(k,\ell)$-th entry of $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$ by $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$.
Assumption 5.
Define $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$. Suppose
$\max_{k\neq\ell}\lvert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\rvert\leq\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}$
for a constant $c_{0}>2/\phi^{\ast}_{0}$.
Assumption 5 imposes a mutual incoherence condition [Van2009On] on the augmented design matrix. It controls correlations between active and inactive coordinates and is essential for deriving $\ell_{\infty}$ error bounds necessary for support recovery.
Theorem 2.
Suppose the assumptions of Theorem 1, together with Assumption 5, hold, and that $\log p\asymp\log q$. Further assume that
$n\geq A_{1}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$
for some constant $A_{1}>0$. With the same tuning parameters $\lambda$ and $\lambda_{g}$ as in (18), we have that
$\max\left\{\lVert\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}\rVert_{\infty},\lVert\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}\rVert_{\infty}\right\}\leq\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}\phi^{\ast}_{0}-2)}\right)$
holds with probability at least $1-C_{3}\exp(-C_{4}\log p)$ for some positive constants $C_{3},C_{4}$.
As in the standard lasso regression, with an additional minimal signal strength condition, the $\ell_{\infty}$ bound in Theorem 2 implies that the support of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$ can be recovered with high probability by thresholding the estimated coefficients at an appropriate level.
Corollary 1.
Define
$\kappa_{0}=\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}\phi^{\ast}_{0}-2)}\right).$
Let $\tilde{S}_{\gamma}=\left\{k:\lvert\hat{\gamma}_{k}\rvert>\kappa_{0}\right\}$ and $\tilde{S}_{\beta}=\left\{k:\lvert\hat{\beta}_{k}\rvert>\kappa_{0}\right\}$. If
$\min\left\{\min_{\ell\in S_{\gamma}}\lvert\gamma_{\ell}\rvert,\min_{k\in S_{\beta}}\lvert\beta_{k}\rvert\right\}\geq 2\kappa_{0},$
then $\operatorname{\operatorname{\mathbb{P}}}(S_{\gamma}=\tilde{S}_{\gamma}\;\text{and}\;S_{\beta}=\tilde{S}_{\beta})\geq 1-C_{3}\exp(-C_{4}\log p)$.
4.1 Understanding the theoretical guarantees
We conclude this section with a discussion of the $\ell_{2}$ estimation error and a comparison with the results of zhangli. The same insights apply to the support recovery guarantees in Theorem 2.
In zhangli, two rates of estimation error are established: an oracle rate and a two-stage rate. The oracle rate is achieved when the true $\boldsymbol{\mathbf{\Gamma}}$ in (3) is known, so that the centered responses can be constructed exactly and the estimation in (7) can be carried out directly. The resulting rate is
$\lVert\hat{\boldsymbol{\mathbf{\eta}}}_{j}-\boldsymbol{\mathbf{\eta}}_{j}\rVert_{2}^{2}\precsim\frac{\sigma^{2}_{\varepsilon_{j}}}{n}(s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})),$ (19)
which holds with probability at least $1-C_{1}\exp[-C_{2}\{s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\}]$ for some positive constants $C_{1},C_{2}$.
In practice, however, $\boldsymbol{\mathbf{\Gamma}}$ is unknown and must be estimated. As reviewed in Section 2.2, zhangli employ a two-stage procedure in which $\boldsymbol{\mathbf{\Gamma}}$ is estimated via the misspecified lasso regression
$\hat{\boldsymbol{\mathbf{\gamma}}}_{j}=\operatorname*{arg\,min}_{\boldsymbol{\mathbf{\gamma}}_{j}}\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}\right\rVert_{2}^{2}+\lambda_{1}\lVert\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{1},$ (20)
where $\lambda_{1}\geq 0$, and the resulting estimate is used to construct centered responses $\hat{\boldsymbol{\mathbf{z}}}_{j}=\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}_{j}$, which are then used in the second-stage regression (7).
Although the resulting two-stage estimator can attain the oracle rate (19), the associated probability bound depends critically on the tuning parameter $\lambda_{1}$ in (20). Specifically, supposing that $\lambda_{1}\propto\sqrt{\kappa_{1}\log q/n}$ for some $\kappa_{1}>0$. Standard lasso results imply that the first-stage estimation error satisfies $\lVert\hat{\boldsymbol{\mathbf{z}}}_{j}-\boldsymbol{\mathbf{z}}_{j}\rVert_{2}^{2}/n\precsim s_{\gamma}\cdot\kappa_{1}\log q/n$ with probability at least $1-3\exp(-\kappa_{1}\log q)$. To ensure that this error does not dominate the second-stage estimation, a union bound must be applied over all $p\times q$ columns of the design matrix. This leads to the overall success probability of at least $1-C_{3}\exp[C_{4}\{\log p-(\kappa_{1}-1)\log q\}]$ for positive constants $C_{3},C_{4}$.
This dependence introduces a nontrivial tradeoff: if $\kappa_{1}$ is too small, the concentration probability deteriorates. If $\kappa_{1}$ is too large, the first-stage penalty becomes overly aggressive, increasing the bias in the first-stage estimator and potentially inflating the overall error. In addition, when $\log p$ dominates $\log q$, the success probability may no longer decay to zero, further weakening the theoretical guarantees.
By contrast, our estimator matches the oracle rate without requiring a separate centering stage, thereby avoiding error propagation from the first stage and eliminating the need to tune an additional penalty parameter. Furthermore, the oracle rate in zhangli is established under the scaling conditions $\log(pq)=O(n^{1/6})$ and $s_{\gamma}+s_{\beta}=o(n^{1/6})$. In comparison, Assumption 4 is strictly weaker and allows for a wider range of scaling regimes. It permits $\log(pq)$ and $s_{\gamma}+s_{\beta}$ to grow at a faster rates relative to $n$, e.g., $\log(pq)=O(n^{1/4}/\log(n))$ and $s_{\gamma}+s_{\beta}=o(n^{1/4})$, while still ensuring consistency. Note however that the scaling in zhangli may be tightened beyond what is reported.
These theoretical advantages in our framework stem from the joint convexity of (14), which enables simultaneous estimation of $(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})$ without intermediate approximation steps. In contrast, the two-stage procedure in zhangli requires additional control over the first-stage estimation error, leading to more restrictive conditions. We provide additional empirical evidence of these advantage in Section 5, particularly in settings where accurate first stage estimation in zhangli is challenging.
5 Simulation studies
We compare our method, termed cspine, with zhangli (referred to as RegGMM) through extensive simulation studies. Comparisons with other methods considered in zhangli are omitted, as RegGMM is shown therein to outperform these alternatives. We use the R implementation of RegGMM, which is generously shared by the authors. Our R implementation uses the sparsegl package to solve (14) [liang_sparsegl_2024].
5.1 Data generation
For each simulation setting, we generate a sparse coefficient matrix $\boldsymbol{\mathbf{\Gamma}}\in\mathbb{R}^{p\times q}$ by independently setting each entry $\Gamma_{jk}$ to be nonzero with probability $0.3$, with nonzero entries drawn from $N(0,1)$. The matrix $\boldsymbol{\mathbf{\Gamma}}$ is then rescaled so that the signal-to-noise ratio is approximately $1$ for each observation.
We generate the baseline precision matrix $\boldsymbol{\mathbf{B}}_{0}$ using a preferential attachment algorithm [barabasi] with power parameter equal to $1$. To introduce covariate-dependent effects, we select $q_{e}=5$ covariates to have nonzero influence and construct the matrices $\boldsymbol{\mathbf{B}}_{h}$ for $h\in[q_{e}]$ as Erdős–R’enyi graphs [erdos] with edge probability $v_{e}=0.01$. This choice ensures that the covariate-dependent graphs are sparser than the population graph [clauset].
Given the graph structure, the nonzero entries of $\boldsymbol{\mathbf{B}}_{h}$} are generated independently from the distribution $\mathrm{Unif}([-0.5,-0.35]\cup[0.35,0.5])$. To ensure diagonal dominance of the resulting precision matrix, each row $j$ is scaled by $\sum_{h}\sum_{k\neq j}\lvert\beta_{jkh}\rvert\times 1.5$ followed by symmetrization of each $\boldsymbol{\mathbf{B}}_{h}$ via averaging $\beta_{jkh}$ and $\beta_{kjh}$. This construction mirrors the data-generating mechanism in zhangli.
Among the $q$ covariates, half are generated as binary variables and half are drawn independently from $\mathrm{Unif}(0,1)$. The continuous covariates are then standardized across the $n$ observations to have zero mean and unit variance. For each observation $i\in[n]$, we construct matrix $\boldsymbol{\mathbf{B}}_{0}+\sum_{h=1}^{q}\boldsymbol{\mathbf{B}}_{h}u_{h}^{(i)}$.
5.2 Model specification and experimental setup
We evaluate both methods under two data-generating specifications. Recall that zhangli adopt the model (3), while our method is based on the alternative parametrization (9). In both models, we enforce all entries of $\operatorname{diag}(\boldsymbol{\mathbf{\Omega}}^{(i)})$ to be $1$, so that $\sigma^{2}_{\varepsilon_{j}}=1$. Under (3), we set $\boldsymbol{\mathbf{\mu}}^{(i)}=\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$. Under our specification (8) and (9), referred to as “natural model”, we set $\boldsymbol{\mathbf{\mu}}^{(i)}=[\boldsymbol{\mathbf{\Omega}}^{(i)}]^{-1}\boldsymbol{\mathbf{\Gamma}}\boldsymbol{\mathbf{u}}^{(i)}$. For both models, we generate responses as $\boldsymbol{\mathbf{x}}^{(i)}\sim N_{p}(\boldsymbol{\mathbf{\mu}}^{(i)},[\boldsymbol{\mathbf{\Omega}}^{(i)}]^{-1})$ for $i\in[n]$.
We consider two regimes: $(p,q)=(25,50)$ and $(25,100)$. For each setting, we generate $100$ independent datasets with sample sizes $n=200$ and $n=400$. Both RegGMM and cspine are applied to each dataset. For a fair comparison, both methods employ two tuning parameters selected via $5$-fold cross-validation. Specifically, using (16), we evaluate a grid consisting of $100$ values of $\lambda_{0}$ and $10$ values of the mixing parameter $\alpha_{s}\in{0.1,0.2,\dotsc,1}$. We report the mean and standard error (over $100$ independent datasets) of the following performance metrics:
  • TPR: true positive rate of detected edges across all $\boldsymbol{\mathbf{B}}_{h}$};
  • $\text{TPR}_{\mathrm{pop}}$: true positive rate for the population graph $\boldsymbol{\mathbf{B}}_{0}$;
  • $\text{FPR}_{\mathrm{pop}}$: false positive rate for $\boldsymbol{\mathbf{B}}_{0}$;
  • $\text{TPR}_{\mathrm{cov}}$: true positive rate for the covariate-dependent graphs $\boldsymbol{\mathbf{B}}_{h}$, $h\in[q]$;
  • $\boldsymbol{\mathbf{\beta}}_{\mathrm{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{j}-\boldsymbol{\mathbf{\beta}}_{j}\rVert_{2}$: nodewise estimation error;
Additional metrics are reported in Section S2. In all simulation settings considered, the overall false positive rate of detected edges is below $0.005$ and is therefore not reported.
5.3 Results
As expected, Table 1 shows that cspine outperforms RegGMM under the natural model across all metrics, where cspine is correctly specified while RegGMM is misspecified. Table 2 presents results under the setting (3), in which RegGMM is correctly specified and cspine is misspecified. Despite this misspecification, cspine remains competitive with RegGMM in both edge detection and estimation error, due to the convex formulation of the nodewise regressions.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.73}$ $(0.08)$ $\mathbf{0.89}$ $(0.07)$ $\mathbf{0.07}$ $(0.02)$ $\mathbf{0.45}$ $(0.21)$ $\mathbf{5.66}$ $(0.50)$
RegGMM $0.62$ $(0.09)$ $0.80$ $(0.10)$ $0.16$ $(0.06)$ $0.30$ $(0.18)$ $7.43$ $(0.61)$
(25, 100) cspine $\mathbf{0.66}$ $(0.10)$ $\mathbf{0.82}$ $(0.09)$ $\mathbf{0.05}$ $(0.02)$ $\mathbf{0.40}$ $(0.24)$ $\mathbf{5.98}$ $(0.55)$
RegGMM $0.55$ $(0.08)$ $0.71$ $(0.08)$ $0.14$ $(0.05)$ $0.28$ $(0.19)$ $7.73$ $(0.56)$
400 (25, 50) cspine $\mathbf{0.77}$ $(0.10)$ $\mathbf{0.96}$ $(0.04)$ $\mathbf{0.09}$ $(0.02)$ $\mathbf{0.52}$ $(0.24)$ $\mathbf{4.50}$ $(0.59)$
RegGMM $0.67$ $(0.08)$ $0.86$ $(0.07)$ $0.17$ $(0.07)$ $0.41$ $(0.20)$ $5.99$ $(0.55)$
(25, 100) cspine $\mathbf{0.72}$ $(0.09)$ $\mathbf{0.94}$ $(0.05)$ $\mathbf{0.08}$ $(0.02)$ $\mathbf{0.44}$ $(0.21)$ $\mathbf{4.96}$ $(0.60)$
RegGMM $0.62$ $(0.08)$ $0.85$ $(0.08)$ $0.22$ $(0.07)$ $0.32$ $(0.17)$ $6.80$ $(0.58)$
Table 1: Mean and standard error of performance metrics over $100$ data sets generated under specification (9), where cspine is correctly specified and RegGMM is misspecified.
$n$ $(p,q)$ Method TPR $\text{TPR}_{\text{pop}}$ $\text{FPR}_{\text{pop}}$ $\text{TPR}_{\text{cov}}$ $\boldsymbol{\mathbf{\beta}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.53}$ $(0.06)$ $\mathbf{0.77}$ $(0.08)$ $\mathbf{0.12}$ $(0.02)$ $0.09$ $(0.10)$ $\mathbf{7.70}$ $(0.29)$
RegGMM $0.50$ $(0.07)$ $0.70$ $(0.10)$ $0.14$ $(0.05)$ $\mathbf{0.12}$ $(0.11)$ $8.02$ $(0.41)$
(25, 100) cspine $\mathbf{0.44}$ $(0.07)$ $\mathbf{0.67}$ $(0.09)$ $\mathbf{0.09}$ $(0.02)$ $0.04$ $(0.06)$ $\mathbf{7.97}$ $(0.32)$
RegGMM $0.39$ $(0.07)$ $0.59$ $(0.12)$ $0.19$ $(0.07)$ $\mathbf{0.05}$ $(0.06)$ $8.80$ $(0.45)$
400 (25, 50) cspine $\mathbf{0.60}$ $(0.05)$ $\mathbf{0.92}$ $(0.05)$ $0.18$ $(0.03)$ $0.12$ $(0.10)$ $6.71$ $(0.28)$
RegGMM $0.59$ $(0.07)$ $0.80$ $(0.10)$ $\mathbf{0.14}$ $(0.05)$ $\mathbf{0.28}$ $(0.15)$ $\mathbf{6.64}$ $(0.42)$
(25, 100) cspine $\mathbf{0.55}$ $(0.06)$ $\mathbf{0.88}$ $(0.07)$ $\mathbf{0.15}$ $(0.02)$ $0.12$ $(0.10)$ $\mathbf{6.83}$ $(0.26)$
RegGMM $\mathbf{0.55}$ $(0.07)$ $0.81$ $(0.09)$ $0.22$ $(0.07)$ $\mathbf{0.19}$ $(0.14)$ $7.15$ $(0.42)$
Table 2: Performance metrics over $100$ data sets under specification (3), where RegGMM is correctly specified and cspine is misspecified.
We next provide numerical verification of the error bound in Theorem 1 through simulation. We fix $p=25$, $q=10$, and $n=300$. In the first set of experiments, we fix a collection $\{\boldsymbol{\mathbf{B}}_{h}\}$ and generate $300$ matrices $\boldsymbol{\mathbf{\Gamma}}$ of varying sparsity levels, using each to generate one data set. In the second set of experiments, we fix a sparse $\boldsymbol{\mathbf{\Gamma}}$ and generate data sets from $300$ different collections of $\{\boldsymbol{\mathbf{B}}_{h}\}$ with varying sparsity levels. In both settings, the data are generated according to specification (9).
In Figure 2, we plot the total estimation error $\boldsymbol{\mathbf{\gamma}}_{\text{err}}+\boldsymbol{\mathbf{\beta}}_{\text{err}}$, where $\boldsymbol{\mathbf{\gamma}}_{\text{err}}=\sum_{j=1}^{p}\lVert\hat{\boldsymbol{\mathbf{\gamma}}}_{j}-\boldsymbol{\mathbf{\gamma}}_{j}\rVert_{2}$, against the sparsity level of $s_{\boldsymbol{\mathbf{\Gamma}}}$ on the left panel, and against the sparsity level of $\{\boldsymbol{\mathbf{B}}_{h}\}$ on the right panel. The observed trend is consistent with the theoretical estimation error in Theorem 1, showing that the $\ell_{2}$ error scales with the square root of the sparsity level.
Finally, we illustrate the advantage of jointly optimizing over $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$, in contrast to the two-stage method of RegGMM. In Figure 3, we plot $\boldsymbol{\mathbf{\beta}}_{\text{err}}$ against $s_{\boldsymbol{\mathbf{\Gamma}}}$ for both methods under model specification (3) of zhangli. The data are generated with $p=25$, $q=50$, and $n=300$, with $300$ replications across varying levels of $s_{\boldsymbol{\mathbf{\Gamma}}}$.
We observe that the estimation error of RegGMM increases as $\boldsymbol{\mathbf{\Gamma}}$ becomes denser. This is consistent with the theoretical finding in zhangli, where accurate estimation of $\{\boldsymbol{\mathbf{B}}_{h}\}$ in the second stage depends on controlling the first-stage estimation error $\lVert\hat{\boldsymbol{\mathbf{\Gamma}}}-\boldsymbol{\mathbf{\Gamma}}\rVert_{F}$, which becomes more challenging as $\boldsymbol{\mathbf{\Gamma}}$ becomes less sparse. In constrast, cspine exhibits stable performance across different sparsity levels of $\boldsymbol{\mathbf{\Gamma}}$, highlighting the advantage of the joint convex formulation in (14), which avoids error propagation from a separate centering stage.
Figure 2: Estimation error of cspine over $300$ data sets plotted against the number of non-zero entries in (a) $\boldsymbol{\mathbf{\Gamma}}$ and (b) $\{\boldsymbol{\mathbf{B}}_{h}\}$. In both cases, the $\ell_{2}$ estimation error bound increases at a rate that is roughly the square-root of the sparsity level.
Figure 3: Estimation error of cspine and RegGMM over $300$ data sets plotted against the number of non-zero entries in $\boldsymbol{\mathbf{\Gamma}}$ with a small horizontal jitter applied. cspine is seen to have better scaling properties in the sparsity of $\boldsymbol{\mathbf{\Gamma}}$. Data were generated via the model specification (3) of zhangli.
6 GBM eQTL analysis
Glioblastoma multiforme (GBM) is the most malignant type of brain cancer and patient prognosis is typically very poor. Although there has been research on the genetic signaling pathways involved in the proliferation of GBM, it remains largely incurable; see glio-hanif for a survey. It is important to understand the conditional independence structure of genes involved in GBM in order to discover new drug therapies [glio-kwiakowska]. Our estimated graphs describe the conditional independence of co-expressions in a gene network; hence, we refer to estimated networks and effects of SNPs on this network.
We reanalyze a GBM eQTL data set that was reported in zhangli. The data set contains microarray and SNP profiling data of $n=401$ GBM patients from the REMBRANDT trial (GSE108476). We use the expression levels of $p=73$ genes known to belong to the human glioma pathway according to the Kyoto Encyclopedia of Genes and Genomes (KEGG) database [KEGG]; the genes and pathways are detailed in Table S3. We also consider $q=118$ SNPs that are local to these $73$ genes. The SNPs are binary-coded, with 0 indicating homozygous major alleles at that locus and $1$ otherwise. Our data set is slightly different from the one used in zhangli; we have a larger cohort size (compared to $n=178$ in zhangli) and do not include age and sex as covariates.
We ran cspine on the data set using $5$-fold cross-validation over $\alpha_{s}=0.1,0.2,\dotsc,1$, and $100$ values of $\lambda_{0}$ in (16). Since cross-validation tends to select relatively dense models, we focus on entries of $\tilde{\boldsymbol{\mathbf{B}}}_{h}$} with magnitude exceeding $0.01$. Under this threshold, our method identifies $28$ SNPs that potentially modify expression patterns in the network. For comparison, we also apply RegGMM to the dataset, using $5$-fold cross-validation over a grid of $100$ $\lambda_{0}$ values and $10$ values of $\alpha_{s}$. After applying the same thresholding procedure, RegGMM identifies $12$ SNPs with nonzero effects on the network. The results of both methods are summarized in Table S4.
Figure 4 displays the estimated population networks, corresponding to $\boldsymbol{\mathbf{B}}_{0}$, obtained from cspine (left) and RegGMM (right). The two networks share many common structures; for example, E2F3, an oncogenic transcription factor, and CALML4, a protein-coding gene, are highly connected nodes in both population networks. RegGMM produces a substantially denser population network, with $750$ total edges compared to $476$ edges identified by cspine.
Figure 4: Population network from GBM eQTL data estimated by cspine (left) and RegGMM (right). The graph structure is determined by the estimates of $\boldsymbol{\mathbf{B}}_{0}$ in (3) and (9). Solid blue lines and dashed red lines indicate positive and negative edge weights, respectively.
Among the SNPs identified as affecting the conditional dependence network, six are detected by both methods. We highlight two reprepresentative examples of these SNPs and their estimated effects on the network.
The first is rs10505980, a variant local to the KRAS gene, whose estimated effects are shown in Figure 5. Both methods detect that this variant influences the co-expression between PIK3CA and SHC2, genes involved in the Ras–Raf–MEK–ERK signaling pathway. In addition, cspine identifies an effect on the co-expression between PIK3CB and CAMK1G. Notably, PIK3CA and PIK3CB are key components of the PI3K/MTOR pathway, and their mutations are known to be oncogenic, with targeted inhibition being an active area of cancer research [samuels, whale_functional_2017]. In particular, inhibition of this pathway has been associated with improved survival in glioblastoma (GBM) [batsios_pi3kmtor_2019]. These findings suggest that rs10505980 may play a role in regulating the co-expression of oncogenic pathways involving PIK3CB and CAMK1G.
A second SNP identified by both methods is rs1267622, located near BRAF, with estimated effects shown in Figure 6. Both methods suggest that this variant modifies the co-expression between PDGFRA and SOS2. Additionally, RegGMM detects an effect on the co-expression between PDGFB and SOS2. Platelet-derived growth factors (PDGFA/B) and their receptors (PDGFRA/B) are known to play important roles in gliomas, where receptor activation triggers the Ras–Raf–MEK–ERK signaling cascade, in which SOS2 participates [soni_genetics_2005, westermark_platelet-derived_2014]. These results suggest that rs1267622 may influence the co-expression of genes in the PDGF receptor pathway and its downstream signaling processes.
Figure 5: Estimated covariate networks for the SNP rs10505980. The graph structure corresponds to the sparsity pattern of the matrix $\boldsymbol{\mathbf{B}}_{h}$ corresponding to this SNP in (9) and (3). Solid blue lines and dashed red lines indicate positive and negative effects of this SNP on the partial correlation of co-expressions. An effect on the co-expression of PIK3CA and SHC2 is detected by both cspine and RegGMM. Our method cspine additionally identifies an effect on the co-expression of PIK3CB and CAMK1G.
Figure 6: Estimated covariate networks for the SNP rs1267622. The graph structure corresponds to the sparsity pattern of the matrix $\boldsymbol{\mathbf{B}}_{h}$ corresponding to this SNP in (3) and (9). Solid blue lines and dashed red lines indicate positive and negative effects of this SNP on the partial correlation of co-expressions. An effect on the co-expression of GADD45A and GADD45B is detected by cspine while RegGMM detects an effect between PDGFRA and SOS2. The remaining two edges, between PDGRFA and PDGFB and PDGRFA and SOS2, are detected by both methods.
7 Discussion
In this work, we propose a new framework for covariate-adjusted Gaussian graphical models based on a natural parametrization that yields a jointly convex formulation. This approach enables unified estimation of both the covariated-dependent mean and covariate-dependent precision matrix, in contrast to the two-stage procedure of zhangli. Our theoretical analysis shows that the convex formulation permits more relaxed scaling conditions on the sparsity level of the true parameter relative to the sample size $n$. These advantages are supported by our simulation results, which demonstrate improved performance in regimes where stagewise methods are challenged by error propagation.
The proposed framework is flexible and does not impose restrictions on the form of the penalty functions $g_{j}$ in (13), allowing for a wide range of sparsity structures. Incorporating more structured penalties is a promising direction for future work. For instance, one may consider hierarchical penalties that reflect the relationship between covariate effects on the mean and precision matrix, where a covariate influencing the precision matrix is expected to also affect the mean, but not necessarily vice versa. Another direction is to develop tuning-free variants of the proposed method, for example by adapting ideas from the square-root lasso [sqrtlasso]. Additionally, methods for statistical inference could be developed by applying debiasing techniques to each nodewise regression [meng_statistical_2025]. From a theoretical perspective, it would be of interest to establish minimax optimality of the estimator in (14). While our results achieve the same rate as in zhangli, the optimality of this rate remains an open question.
Acknowledgements
We are grateful to Emma Zhang for generously making the RegGMM software and GBM eQTL data set available to us.
Use was made of computational facilities purchased with funds from the National Science Foundation (CNS-1725797) and administered by the Center for Scientific Computing (CSC). The CSC is supported by the California NanoSystems Institute and the Materials Research Science and Engineering Center (MRSEC; NSF DMR 2308708) at UC Santa Barbara.
Data availability statement
The code and data supporting Section 5 can be found on the corresponding author’s GitHub. [^1] https://github.com/roobnloo/cspine-sim The data that support the findings Section 6 are available from the corresponding author, RL, upon reasonable request.
Supplemental Materials
Convex estimation of Gaussian graphical regression models with covariates
S1 Algorithm outline
Input :  $\boldsymbol{\mathbf{U}}\in\mathbb{R}^{n\times q}$ matrix of covariates.
$\boldsymbol{\mathbf{X}}\in\mathbb{R}^{n\times p}$ matrix of responses.
$\boldsymbol{\mathbf{W}}_{-j}\in\mathbb{R}^{n\times(p-1)(q+1)}$ matrix of interactions.
$g_{j}\colon\mathbb{R}^{q}\times\mathbb{R}^{(p-1)(q+1)}\rightarrow\mathbb{R}$ convex penalty functions for $j\in[p]$.
$H(\hat{\boldsymbol{\mathbf{\gamma}}}_{j},\hat{\boldsymbol{\mathbf{\beta}}}_{j})$ a function to estimate the error variance $\sigma^{2}_{\varepsilon_{j}}$.
Output :  $\hat{\boldsymbol{\mathbf{\Gamma}}}\in\mathbb{R}^{p\times q}$, $\tilde{\boldsymbol{\mathbf{B}}}_{0},\tilde{\boldsymbol{\mathbf{B}}}_{1},\dotsc\tilde{\boldsymbol{\mathbf{B}}}_{q}\in\mathbb{R}^{p\times p}$, and $\hat{\sigma}^{2}_{\varepsilon_{j}}$ for $j\in[p]$.
for $j=1,2,\dotsc,p$ do
   solve
$(\hat{\boldsymbol{\mathbf{\gamma}}}_{j},\hat{\boldsymbol{\mathbf{\beta}}}_{j})=\operatorname*{arg\,min}_{\boldsymbol{\mathbf{\gamma}}_{j},\,\boldsymbol{\mathbf{\beta}}_{j}}\,\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}_{j}-\boldsymbol{\mathbf{W}}_{-j}\boldsymbol{\mathbf{\beta}}_{j}\right\rVert_{2}^{2}+g_{j}(\boldsymbol{\mathbf{\gamma}}_{j},\boldsymbol{\mathbf{\beta}}_{j}).$
  Set $\hat{\sigma}^{2}_{\varepsilon_{j}}=H(\hat{\boldsymbol{\mathbf{\gamma}}}_{j},\hat{\boldsymbol{\mathbf{\beta}}}_{j})$ and $\tilde{\beta}_{jkh}=-\hat{\beta}_{jkh}/\hat{\sigma}^{2}_{\varepsilon_{j}}$ for $k\neq j$, $h=0,1,\dotsc,q$.
end for
Set $\hat{\boldsymbol{\mathbf{\Gamma}}}=[\hat{\boldsymbol{\mathbf{\gamma}}}_{1},\dotsc,\hat{\boldsymbol{\mathbf{\gamma}}}_{p}]^{\top}$.
Initialize $\tilde{\boldsymbol{\mathbf{B}}}_{0}=\tilde{\boldsymbol{\mathbf{B}}}_{1}=\dotsb=\tilde{\boldsymbol{\mathbf{B}}}_{q}=\mathbf{0}$.
for $h=0,1,\dotsc,q$ do
   Apply the and-rule:
$\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{jk}=[\tilde{\boldsymbol{\mathbf{B}}}_{h}]_{kj}=\tilde{\beta}_{jkh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert<\lvert\tilde{\beta}_{kjh}\rvert\}}+\tilde{\beta}_{kjh}\mathbf{1}_{\{\lvert\tilde{\beta}_{jkh}\rvert>\lvert\tilde{\beta}_{kjh}\rvert\}},\quad j\neq k$
end for
return $\hat{\boldsymbol{\mathbf{\Gamma}}},\tilde{\boldsymbol{\mathbf{B}}}_{0},\tilde{\boldsymbol{\mathbf{B}}}_{1},\dotsc,\tilde{\boldsymbol{\mathbf{B}}}_{q},\hat{\sigma}^{2}_{\varepsilon_{1}},\dotsc,\hat{\sigma}^{2}_{\varepsilon_{p}}$.
Algorithm 1 Covariate-adjusted Sparse Precision with Natural Estimation (cspine)
S2 Additional Simulation Results
In addition to the metrics reported in Table 1, we report here the support recovery metrics for the estimated precision matrix and the estimation error of the mean given by $\boldsymbol{\mathbf{\mu}}_{\text{err}}=\sum_{i=1}^{n}\lVert\hat{\boldsymbol{\mathbf{\mu}}}_{i}-\boldsymbol{\mathbf{\mu}}_{i}\rVert_{2}^{2}/n$ where $\hat{\boldsymbol{\mathbf{\mu}}}_{i}$ is computed from (17). The results are shown in Table S1 for data generated under the natural model (9) and in Table S2 for data generated under the original model (3). We see that cspine underperforms in mean prediction error in both data models. This is a reasonable finding since our method does not focus on estimating the mean. Instead, our method targets the precision matrix and its components $\{\boldsymbol{\mathbf{B}}_{h}\}$ while incorporating the covariate effects on the mean.
$n$ $(p,q)$ Method $\boldsymbol{\mathbf{\Omega}}_{\text{TPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{FPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$ $\boldsymbol{\mathbf{\mu}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.84}$ $(0.06)$ $\mathbf{0.10}$ $(0.03)$ $\mathbf{1.09}$ $(0.15)$ $10.14$ $(1.49)$
RegGMM $0.74$ $(0.08)$ $0.18$ $(0.06)$ $1.82$ $(0.26)$ $\mathbf{8.28}$ $(1.13)$
(25, 100) cspine $\mathbf{0.79}$ $(0.07)$ $\mathbf{0.08}$ $(0.02)$ $\mathbf{1.22}$ $(0.17)$ $11.94$ $(1.12)$
RegGMM $0.68$ $(0.07)$ $0.17$ $(0.04)$ $1.95$ $(0.28)$ $\mathbf{10.11}$ $(1.01)$
400 (25, 50) cspine $\mathbf{0.93}$ $(0.04)$ $\mathbf{0.12}$ $(0.03)$ $\mathbf{0.66}$ $(0.08)$ $5.80$ $(1.05)$
RegGMM $0.83$ $(0.06)$ $0.20$ $(0.07)$ $1.22$ $(0.17)$ $\mathbf{4.96}$ $(0.72)$
(25, 100) cspine $\mathbf{0.89}$ $(0.04)$ $\mathbf{0.11}$ $(0.03)$ $\mathbf{0.76}$ $(0.09)$ $7.98$ $(1.05)$
RegGMM $0.79$ $(0.06)$ $0.24$ $(0.07)$ $1.40$ $(0.20)$ $\mathbf{6.76}$ $(0.91)$
Table S1: Additional metrics under the natural model setting (9).
$n$ $(p,q)$ Method $\boldsymbol{\mathbf{\Omega}}_{\text{TPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{FPR}}$ $\boldsymbol{\mathbf{\Omega}}_{\text{err}}$ $\boldsymbol{\mathbf{\mu}}_{\text{err}}$
200 (25, 50) cspine $\mathbf{0.69}$ $(0.07)$ $\mathbf{0.14}$ $(0.03)$ $\mathbf{2.18}$ $(0.24)$ $6.20$ $(0.67)$
RegGMM $0.65$ $(0.09)$ $0.16$ $(0.05)$ $2.20$ $(0.25)$ $\mathbf{5.89}$ $(0.66)$
(25, 100) cspine $\mathbf{0.54}$ $(0.07)$ $\mathbf{0.12}$ $(0.02)$ $\mathbf{2.43}$ $(0.25)$ $9.95$ $(1.16)$
RegGMM $0.51$ $(0.09)$ $0.21$ $(0.06)$ $2.69$ $(0.32)$ $\mathbf{9.58}$ $(1.12)$
400 (25, 50) cspine $\mathbf{0.79}$ $(0.06)$ $0.20$ $(0.03)$ $1.54$ $(0.18)$ $3.34$ $(0.36)$
RegGMM $0.75$ $(0.07)$ $\mathbf{0.16}$ $(0.05)$ $\mathbf{1.49}$ $(0.19)$ $\mathbf{3.07}$ $(0.31)$
(25, 100) cspine $0.73$ $(0.06)$ $\mathbf{0.16}$ $(0.03)$ $\mathbf{1.69}$ $(0.22)$ $5.83$ $(0.63)$
RegGMM $\mathbf{0.74}$ $(0.07)$ $0.24$ $(0.07)$ $\mathbf{1.69}$ $(0.23)$ $\mathbf{5.58}$ $(0.54)$
Table S2: Additional metrics under specification (3).
S3 Additional Results from GBM eQTL Analysis
Table S3 displays the genes belonging to known signaling pathways associated with GBM. The table is copied from zhangli. Table S4 lists the SNPs that are detected by cspine to affect network co-expression.
name genes references
PI3K/Akt/mTOR PIK3CA, PIK3CB, PIK3CD, network
signaling pathway PIK3R3, PTEN, AKT1, AKT2, AKT3
MTOR, IGF1, PRKCA
EGF, EGFR, GRB2, brennan
SOS1, SOS2, IGF1
Ras-Raf-MEK-ERK SHC1, SHC2, SHC3, SHC4
signaling pathway MAPK1, MAPK3, MAP2K1, MAP2K2
HRAS, KRAS, NRAS,
RAF1, ARAF, BRAF, PRKCA
CALM1,CALML3, CALML4, CALML5, maklad
calcium (Ca+2) CALML6, CAMK1,CAMK4, CAMK1D,
signaling pathway CAMK1G,CAMK2A, CAMK2B,
CAMK2D,CAMK2G, PRKCA
p53 TP53, MDM2, DDB2, PTEN, IGF1 network
signaling pathway CDK4, CDK6, CDKN1A, CDKN2A
Table S3: Gene signaling pathways related to GBM copied from zhangli and represented in https://www.genome.jp/pathway/hsa05214.
SNP Co-expressed genes
cspine RegGMM
rs1267622 (SOS2, PDGFRA) (PDGFRA, PDGFB) (SOS2, PDGFB) (PDGFB, PDGFRA)
rs6701524 (AKT1, PIK3CD) (AKT1, MAP2K2) (MAP2K2, AKT1)
rs10505980 (PIK3CB, CAMK1G) (PIK3CA, SHC2) (PIK3CA, SHC2)
rs10519201 (E2F3, PIK3CA) (CDK6, PIK3CA)
rs10492796 (CDKN2A, CAMK2A) (CDKN2A, CAMK2A)
rs10488137 (CDKN2A, CAMK2A) (CDKN2A, CAMK2A)
rs10512510 (PIK3CD, PDGFRB) (E2F1, SOS1)
rs723211 (GADD45B, IGF1R)
rs723210 (GADD45B, AKT3)
rs3730504 (GADD45B, IGF1R)
rs198563 (CDKN2A, PIK3R3)
rs5871 (PIK3CB, CALML5)
rs10518759 (PIK3CA, CAMK1G)
rs1567810 (PIK3CB, IGF1R)
rs9303511 (SOS1, SHC2)
rs2078153 (PIK3R3, EGFR)
rs958686 (GADD45B, SOS1)
rs10496180 (GADD45B, CAMK2A)
rs559160 (CDKN2A, PIK3R3)
rs10488881 (GADD45B, AKT3)
rs759947 (PIK3CD, NRAS)
rs759948 (MAP2K1, PDGFRA)
rs727923 (PIK3R3, CAMK2A)
rs42670 (MAP2K1, CALML5)
rs39605 (DDB2, CALML5)
rs10488141 (PIK3CB, SOS2)
rs7785163 (CALML5, IGF1R)
rs1267600 (CALML5, PDGFRB)
rs10509346 (CALML5, PDGFRA)
rs473698 (CDK6, CAMK2A)
rs503314 (CCND1, RAF1)
rs10489985 (HRAS, SOS1)
rs2051785 (CDK6, AKT3)
rs9326835 (CDK6, CAMK2A)
Table S4: Estimated effects of SNPs on gene co-expressions.
S4 Proofs
Our proof strategies for Theorems 1 and 2 follow that of zhangli, with modifications made to accommodate our concatenated design matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$. We will first recall several lemmas from the literature and then prove some supporting lemmas before moving onto the main proofs.
S4.1 Technical lemmas
Lemma 1 (bellec Lemma 1).
Let $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be any convex function and let
$\hat{\boldsymbol{\mathbf{\beta}}}\in\operatorname*{arg\,min}_{\boldsymbol{\mathbf{\beta}}\in\mathbb{R}^{d}}\left\{\lVert\boldsymbol{\mathbf{y}}-\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}+g(\boldsymbol{\mathbf{\beta}})\right\}$
where $\boldsymbol{\mathbf{H}}\in\mathbb{R}^{n\times d}$ and $\boldsymbol{\mathbf{y}}\in\mathbb{R}^{n}$. Then for all $\boldsymbol{\mathbf{\beta}}\in\mathbb{R}^{d}$,
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{y}}-\boldsymbol{\mathbf{H}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}+g(\hat{\boldsymbol{\mathbf{\beta}}})+\frac{1}{2n}\lVert\boldsymbol{\mathbf{H}}(\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}})\rVert_{2}^{2}\leq\frac{1}{2n}\lVert\boldsymbol{\mathbf{y}}-\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}+g(\boldsymbol{\mathbf{\beta}}).$
Lemma 2 (graybill Theorem F).
Let $\boldsymbol{\mathbf{\varepsilon}}\sim N_{p}(\boldsymbol{\mathbf{0}},\sigma^{2}\boldsymbol{\mathbf{I}}_{p})$ and let $A$ be a $p\times p$ idempotent matrix with rank $r\leq p$. Then $\boldsymbol{\mathbf{\varepsilon}}^{\top}A\boldsymbol{\mathbf{\varepsilon}}/\sigma^{2}\sim\chi^{2}_{r}$.
Lemma 3 (laurent Lemma 1).
Suppose that $U\sim\chi^{2}_{r}$. For any $x>0$ it holds that
$\operatorname{\operatorname{\mathbb{P}}}(U-r\geq 2\sqrt{rx}+2x)\leq e^{-x}.$
Lemma 4 (vershynin Proposition 5.16).
Let $X_{1},\dotsc,X_{n}$ be independent, mean zero sub-exponential random variables. Let $v_{1}=\max_{i}\lVert X_{i}\rVert_{\psi_{1}}$ where $\lVert\cdot\rVert_{\psi_{1}}$ is the sub-exponential norm. Then there exists a constant $c$ such that for any $t>0$ we have
$\operatorname{\operatorname{\mathbb{P}}}\left(\left\lvert\sum_{i=1}^{n}X_{i}\right\rvert\geq t\right)\leq 2\exp\left\{-c\min\left(\frac{t^{2}}{v_{1}^{2}n},\frac{t}{v_{1}}\right)\right\}.$
Lemma 5 comes from Theorem 4.1 in kuchi applied to marginally sub-exponential random vectors.
Lemma 5 (kuchi Theorem 4.1).
Let $\bm{Z}_{1},\dotsc,\bm{Z}_{n}$ be independent random vectors in $\mathbb{R}^{p}$ and let $\boldsymbol{\mathbf{Z}}=[\bm{Z}_{1},\dotsc,\bm{Z}_{n}]^{\top}$. Assume each element of $\bm{Z}_{i}$ is sub-exponential with bounded sub-exponential norm for all $i\in[n]$; $\lVert Z_{ij}\rVert_{\psi_{1}}=\sup_{d\geq 1}d^{-1}(\operatorname{\operatorname{\mathbb{E}}}\lvert Z_{ij}\rvert^{d})^{1/d}\leq K_{z}$ for all $i\in[n]$ and $j\in[p]$ for some constant $K_{z}>0$. Let $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}=\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n$ and $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}}=\operatorname{\operatorname{\mathbb{E}}}[\boldsymbol{\mathbf{Z}}^{\top}\boldsymbol{\mathbf{Z}}/n]$. Define
$\Upsilon_{n}=\max_{j,k}\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}(Z_{ij}Z_{ik}).$
Then we have
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\lvert\boldsymbol{\mathbf{v}}^{\top}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{Z}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{Z}}})\boldsymbol{\mathbf{v}}\right\rvert\precsim k\sqrt{\frac{\Upsilon_{n}\log p}{n}}+K_{z}^{2}\frac{k(\log n\log p)^{2}}{n}$
with probability at least 1O(1/p)1-O(1/p).
With Assumptions 1 and 2, we may apply Lemma 5 to our design matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ because it is elementwise sub-exponential, each entry being the product of a sub-Gaussian and a bounded random variable. Furthermore, we have by the Cauchy-Schwarz inequality that
$\max_{\ell,k}\frac{1}{n}\sum_{i=1}^{n}\operatorname{Var}\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{i\ell}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{ik}\right)\leq\max_{\ell_{1},\ell_{2},\ell_{3},\ell_{4}}\operatorname{\operatorname{\mathbb{E}}}\left(X_{\ell_{1}}^{(1)^{2}}X_{\ell_{2}}^{(1)^{2}}U_{\ell_{3}}^{(1)^{2}}U_{\ell_{4}}^{(1)^{2}}\right)=O(1)$
since the entries of $\boldsymbol{\mathbf{X}}^{(1)}$ and $\boldsymbol{\mathbf{U}}^{(1)}$ have bounded moments. Thus $\Upsilon_{n}=O(1)$ in our setting.
Lemma 6 (loh Lemma 12).
Let $\boldsymbol{\mathbf{\Sigma}}\in\mathbb{R}^{p\times p}$ be a symmetric matrix such that $\lvert\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}\boldsymbol{\mathbf{v}}\rvert\leq\delta_{1}$ for all $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p}$ with $\lVert\boldsymbol{\mathbf{v}}\rVert_{2}=1$ and $\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq 2s$. It holds for all $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p}$ that
$\lvert\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}\boldsymbol{\mathbf{v}}\rvert\leq 27\delta_{1}\left(\lVert\boldsymbol{\mathbf{v}}\rVert_{2}^{2}+\frac{1}{s}\lVert\boldsymbol{\mathbf{v}}\rVert_{1}^{2}\right).$
Treating $j$ as fixed and suppressing the dependence of $\boldsymbol{\mathbf{W}}_{-j}$ on $j$, recall that the sample and population Gram matrices are defined as
$\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}],\quad\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\operatorname{\operatorname{\mathbb{E}}}\left(\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\right).$
Lemma 7.
For a set of indices $S\subset[p(q+1)-1]$, denote by $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{S}$ the submatrix of $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ with columns indexed by $S$. Under Assumptions 1-4, there exist constants $M_{uw}$ and $C_{0}$ such that with probability at least $1-C_{0}\exp(-\log(pq))$ we have
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]_{S}\right\rVert_{\mathrm{op}}^{2}\leq M_{uw}$
for all $S$ satisfying $\lvert S\rvert\leq\hat{s}^{\mathrm{max}}_{\gamma}+\hat{s}^{\mathrm{max}}_{\beta}$, provided that $(\hat{s}_{\gamma}^{\mathrm{max}}+\hat{s}_{\beta}^{\mathrm{max}})\log(pq)=O(\sqrt{n})$, as assumed in Theorem 1.
Proof.
Letting $k=\lvert S\rvert$, it suffices to show that $\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\boldsymbol{\mathbf{v}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ is bounded with the provided probability. We may write
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\boldsymbol{\mathbf{v}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ $=\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\{\boldsymbol{\mathbf{v}}^{\top}\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}+\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}\right\}.$
The second term is bounded by Assump 3. For the first term, by Lemma 5 we have
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq k,\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}\leq 1}\left\{\boldsymbol{\mathbf{v}}^{\top}\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}\right\}\precsim k\left(\frac{\log(pq)}{n}\right)^{1/2}+\frac{k\log(pq)^{2}}{n/(\log n)^{2}}$
with probability at least $1-C_{0}\exp(-\log(pq))$ and we see that the right-hand side is $o(1)$ by Assumption 4. ∎
S4.2 Theorem 1 proof sketch
The proof of Theorem 1 closely follows that of Theorem 1 in zhangli. We give a sketch of the main proof technique, omitting constants while focusing on the important quantities. As such, the inequalities in this subsection should not be taken literally. Letting $\boldsymbol{\mathbf{\nu}}=\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\Delta}}=\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}$ denote the error terms, we first bound the sample squared error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$ by the complexity term $E_{j}=(\sigma^{2}_{\varepsilon}/n)\cdot\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$. Recall the penalty term is $\mathrm{pen}=\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}$. The proof occurs in three steps.
In the first step, a basic inequality is used to bound the sample squared error by the stochastic term plus the penalty term,
$\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\epsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\mathrm{pen}.$
The stochastic term is further bounded by the sample squared error and a projection $\mathcal{P}$ of error terms onto the support set of $\boldsymbol{\mathbf{\nu}}$ and $\boldsymbol{\mathbf{\Delta}}$:
$\langle\boldsymbol{\mathbf{\epsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle\leq\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lVert\mathcal{P}\boldsymbol{\mathbf{\epsilon}}\rVert_{2}^{2}.$
In the second step, by a counting argument and with $\chi^{2}$ tails, the projection term is shown to satisfy
$\lVert\mathcal{P}\boldsymbol{\mathbf{\epsilon}}\rVert_{2}^{2}\leq E_{j}+\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{r}{n}$
where $r$ is a high probability term with an exponentially small tail. Combining the above yields
$\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\leq E_{j}+\mathrm{pen}$
with high probability.
In the third step, the penalty term is shown to satisfy
$\mathrm{pen}\leq E_{j}+\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}.$
Lemmas 5 and 6 are then used to bound the difference between sample squared error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$ and the population squared error $\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$ by the estimation error. Combined with the above, this yields an upper bound on the population error as
$\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\precsim E_{j}+\frac{1}{L}(E_{j}+\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}).$
for a large constant $L$. Finally Assumption 3 is applied to get $\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\precsim E_{j}$ whence the desired bound.
S4.3 Proof of Theorem 1
For ease of notation, we will drop the dependence of $\boldsymbol{\mathbf{\gamma}}_{j}$, $\boldsymbol{\mathbf{\beta}}_{j}$, $\boldsymbol{\mathbf{\varepsilon}}_{j}$ and $\boldsymbol{\mathbf{W}}_{-j}$ on $j$. Let $S_{\beta},S_{\gamma},\hat{S}_{\beta},\hat{S}_{\gamma}$ be the support sets of $\boldsymbol{\mathbf{\beta}},\boldsymbol{\mathbf{\gamma}},\hat{\boldsymbol{\mathbf{\beta}}},\hat{\boldsymbol{\mathbf{\gamma}}}$, respectively. Let $S_{\beta,g}=\{h\in[q]:\boldsymbol{\mathbf{\beta}}_{h}\neq\boldsymbol{\mathbf{0}}\}$ index the blocks $\boldsymbol{\mathbf{\beta}}_{h}$ of $\boldsymbol{\mathbf{\beta}}$ that are not identically zero and let $\hat{S}_{\beta,g}$ be the corresponding block indices for $\hat{\boldsymbol{\mathbf{\beta}}}$. For any vector $\boldsymbol{\mathbf{v}}$ and set of block indices $S$, let $\boldsymbol{\mathbf{v}}_{(S)}$ denote the sub-vector containing blocks in $S$. Let $s_{\beta},s_{\gamma},s_{\beta,g},\hat{s}_{\beta},\hat{s}_{\gamma},\hat{s}_{\beta,g}$ be the number of elements in $S_{\beta},S_{\gamma},S_{\beta,g},\hat{S}_{\beta},\hat{S}_{\gamma},\hat{S}_{\beta,g}$, respectively. Our proof occurs in three steps.
Step 1
In this step we bound the error $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}/n$ by the stochastic term $\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle/n$, which is then bounded by a projection of $\boldsymbol{\mathbf{\varepsilon}}$ onto the columns of $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$.
Since our penalty function
$g(\boldsymbol{\mathbf{\gamma}},\boldsymbol{\mathbf{\beta}})=\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}$
is convex, by Lemma 1 we have
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}+g(\hat{\boldsymbol{\mathbf{\gamma}}},\hat{\boldsymbol{\mathbf{\beta}}})+\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\leq\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\beta}}\rVert_{2}^{2}+g({\boldsymbol{\mathbf{\gamma}}},{\boldsymbol{\mathbf{\beta}}})$
where $\boldsymbol{\mathbf{\nu}}=\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\Delta}}=\hat{\boldsymbol{\mathbf{\beta}}}-\boldsymbol{\mathbf{\beta}}$. Since $\boldsymbol{\mathbf{\varepsilon}}=\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\beta}}$ we may write
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}$ $=\frac{1}{2n}\lVert\boldsymbol{\mathbf{\varepsilon}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$
$=\frac{1}{2n}\lVert\boldsymbol{\mathbf{\varepsilon}}\rVert_{2}^{2}-\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\frac{1}{2n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}.$
Plugging this into the previous expression and substituting the penalty expression then yields
$\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}.$
Notice that $\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S^{c}_{\beta}}\rVert_{1}$, $\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}}\rVert_{1}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}^{c}}\rVert_{1}$, and $\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}=\lVert\boldsymbol{\mathbf{\beta}}_{S_{\beta}}\rVert_{1}$. Hence we can express the above as
$\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\underbrace{\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}}_{\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\underbrace{\lambda\lVert\boldsymbol{\mathbf{\beta}}_{S_{\beta}}\rVert_{1}}_{\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}.$
Thus we have
$\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda(\lVert\boldsymbol{\mathbf{\beta}}_{S_{\beta}}\rVert_{1}-\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{S_{\beta}}\rVert_{1})+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}$
$\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}$
using the triangle inequality for the $\ell_{1}$ norm. The same development holds for $\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}$. Finally, notice that
$\lVert\boldsymbol{\mathbf{\Delta}}_{(S^{c}_{\beta,g})}\rVert_{1,2}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{(S^{c}_{\beta,g})}\rVert_{1,2},\;\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}=\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{(S_{\beta,g})}\rVert_{1,2}+\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{(S^{c}_{\beta,g})}\rVert_{1,2},\;\text{and}\;\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}=\lVert\boldsymbol{\mathbf{\beta}}_{(S_{\beta,g})}\rVert_{1,2}.$
Hence the previous development holds for the $\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}$ term by the triangle inequality of $\lVert\cdot\rVert_{1,2}$. All in all we have
$\begin{split}&\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g}^{c})}\rVert_{1,2}\\ &\qquad\leq\frac{1}{n}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}.\end{split}$ (S1)
Now let $\mathcal{I}$ and $\mathcal{J}$ be arbitrary index sets of the columns of $\boldsymbol{\mathbf{U}}$ and $\boldsymbol{\mathbf{W}}$ respectively. Denote by $\mathcal{P}_{\mathcal{I},\mathcal{J}}$ the orthogonal projection onto the columns of $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ indexed by $(\mathcal{I},\mathcal{J})$. Let $\mathcal{I}_{0}=S_{\gamma}\cup\hat{S}_{\gamma}$ and $\mathcal{J}_{0}=S_{\beta}\cup\hat{S}_{\beta}$ denote the unions of the true and estimated support sets of $\boldsymbol{\mathbf{\gamma}}$ and $\boldsymbol{\mathbf{\beta}}$. We seek to bound the stochastic term
$\begin{split}\langle\boldsymbol{\mathbf{\varepsilon}},\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle&=\langle\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}}),\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rangle\\ &\leq\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}\leq\frac{1}{2a_{1}}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{a_{1}}{2}\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}.\end{split}$ (S2)
The last inequality follows from the fact that $2xy\leq ax^{2}+y^{2}/a$ holds for any constant $a>0$ and real
Step 2
Following zhangli, we first bound the term $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ with a counting argument. For fixed $s^{\prime}_{\gamma}$, $s^{\prime}_{\beta}$, and $s^{\prime}_{\beta,g}$, we will bound the cardinality of the set
$\mathcal{H}(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})=\{(\mathcal{I},\mathcal{J})\subset[q]\times[(p-1)(q+1)]:\lvert\mathcal{I}\rvert=s^{\prime}_{\gamma},\,\lvert\mathcal{J}\rvert=s^{\prime}_{\beta},\,\lvert g(\mathcal{J})\rvert=s^{\prime}_{\beta,g}\}$
where $g(\mathcal{J})$ is the number of nonzero groups of $\boldsymbol{\mathbf{\beta}}_{\mathcal{J}}$. For ease of notation, we will write $\mathcal{H}$ while keeping in mind its dependence on $s^{\prime}_{\gamma},s^{\prime}_{\beta},$ and $s^{\prime}_{\beta,g}$. We will show that
$\log\lvert\mathcal{H}\rvert\leq s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)$
by considering the two cases $s^{\prime}_{\beta}=s^{\prime}_{\beta,g}$ and $s^{\prime}_{\beta}>s^{\prime}_{\beta,g}$ separately. These are the only cases since we cannot have more nonzero groups than nonzero elements.
  1. 1.
    $s^{\prime}_{\beta}=s^{\prime}_{\beta,g}$: In this case, we have $\lvert\mathcal{H}\rvert\leq\binom{q}{s^{\prime}_{\gamma}}\binom{q}{s^{\prime}_{\beta,g}}(p-1)^{s^{\prime}_{\beta}}$. Hence
    $\log\lvert\mathcal{H}\rvert$ $\leq\log\binom{q}{s^{\prime}_{\gamma}}+\log\binom{q}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(p-1)$
    $\leq s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)$
    where we use $\log\binom{n}{k}\leq k\log(en/k)$ which follows from Stirling’s approximation.
  2. 2.
    $s^{\prime}_{\beta}>s^{\prime}_{\beta,g}$: In this case, the cardinality is bounded by
    $\lvert\mathcal{H}\rvert\leq\binom{q}{s^{\prime}_{\gamma}}\binom{q}{s^{\prime}_{\beta,g}}\binom{(p-1)(s^{\prime}_{\beta,g}+1)}{s^{\prime}_{\beta}}$.
    Since by Stirling’s approximation
    $\log\binom{(p-1)(s^{\prime}_{\beta,g}+1)}{s^{\prime}_{\beta}}\leq s^{\prime}_{\beta}\log\frac{e(p-1)(s^{\prime}_{\beta,g}+1)}{s^{\prime}_{\beta}}\leq s^{\prime}_{\beta}\log(ep),$
    we have
    $\log\lvert\mathcal{H}\rvert\leq s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)$
    as desired.
Define $k_{0}$ to be the exponential of the right hand side of the above inequality, so that $\lvert\mathcal{H}\rvert\leq k_{0}$. For any $(\mathcal{I},\mathcal{J})\in\mathcal{H}$, since $\mathcal{P}_{\mathcal{I},\mathcal{J}}$ is idempotent, Lemma 2 implies
$\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}/\sigma^{2}_{\varepsilon}\sim\chi_{d}^{2}$
where $d\leq\lvert\mathcal{I}\rvert+\lvert\mathcal{J}\rvert=s^{\prime}_{\gamma}+s^{\prime}_{\beta}$ is the rank of $\mathcal{P}_{\mathcal{I},\mathcal{J}}$. By Lemma 3 we have for arbitrary $t^{\prime}>0$ that
$\operatorname{\operatorname{\mathbb{P}}}\left(\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq\sigma^{2}_{\varepsilon}(2\sqrt{dt^{\prime}}+d+2t^{\prime})\right)\leq e^{-t^{\prime}}.$
Since $2\sqrt{dt^{\prime}}\leq d+t^{\prime}$ and $d\leq s^{\prime}_{\gamma}+s^{\prime}_{\beta}\leq\log k_{0}$, we have
$\operatorname{\operatorname{\mathbb{P}}}\left(\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq\sigma^{2}_{\varepsilon}(2\log k_{0}+3t^{\prime})\right)\leq e^{-t^{\prime}}.$
Taking the supremum over $\mathcal{H}$ and applying the union bound yields
$\operatorname{\operatorname{\mathbb{P}}}\left(\sup_{(\mathcal{I},\mathcal{J})\in\mathcal{H}}\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq\sigma^{2}_{\varepsilon}(2\log k_{0}+3t^{\prime})\right)\leq\lvert\mathcal{H}\rvert e^{-t^{\prime}}.$
Now set $t^{\prime}=t/3+\log k_{0}$ for $t>0$. Then $\lvert\mathcal{H}\rvert e^{-t^{\prime}}=\lvert\mathcal{H}\rvert/{k_{0}}\cdot e^{-t/3}\leq e^{-t/3}$ since $\lvert\mathcal{H}\rvert\leq k_{0}$. Substituting these expressions into the previous bound (and recalling that we have fixed $s^{\prime}_{\gamma}$, $s^{\prime}_{\beta}$, and $s^{\prime}_{\beta,g}$) yields
$\operatorname{\operatorname{\mathbb{P}}}\left(\sup_{(\mathcal{I},\mathcal{J})\in\mathcal{H}(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})}\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\geq 5\sigma^{2}_{\varepsilon}{\left[s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)\right]}+\sigma^{2}_{\varepsilon}t\right)\leq e^{-t/3},$ (S3)
recalling that
$\log k_{0}=s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep).$
This gives a concentration bound of $\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ over all possible subsets $\mathcal{I}$ of columns of $\boldsymbol{\mathbf{U}}$ and $\mathcal{J}$ of columns of $\boldsymbol{\mathbf{W}}$ satisfying $\lvert\mathcal{I}\rvert=s^{\prime}_{\gamma}$, $\lvert\mathcal{J}\rvert=s^{\prime}_{\beta}$, and $\lvert g(\mathcal{J})\rvert=s^{\prime}_{\beta,g}$. Recalling that $\mathcal{I}_{0}$ and $\mathcal{J}_{0}$ are the support sets of $\boldsymbol{\mathbf{\nu}}$ and $\boldsymbol{\mathbf{\Delta}}$, we can now bound $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$. Define
$r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})=\left(\sup_{(\mathcal{I},\mathcal{J})\in\mathcal{H}(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})}\lVert\mathcal{P}_{\mathcal{I},\mathcal{J}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}-5\sigma^{2}_{\varepsilon}\left\{s^{\prime}_{\gamma}\log\frac{eq}{s^{\prime}_{\gamma}}+s^{\prime}_{\beta,g}\log\frac{eq}{s^{\prime}_{\beta,g}}+s^{\prime}_{\beta}\log(ep)\right\}\right)_{+}$
and
$r=\sup_{s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}}r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}).$
It is clear that
$\lvert\mathcal{I}_{0}\rvert\leq s_{\gamma}+\hat{s}_{\gamma},\;\lvert\mathcal{J}_{0}\rvert\leq s_{\beta}+\hat{s}_{\beta},\;\text{and}\;\lvert g(\mathcal{J}_{0})\rvert\leq s_{\beta,g}+\hat{s}_{\beta,g}.$
Thus we have
$\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}\leq 5\sigma^{2}_{\varepsilon}\left\{(s_{\gamma}+\hat{s}_{\gamma})\log\frac{eq}{s_{\gamma}}+(s_{\beta}+\hat{s}_{\beta})\log(ep)+(s_{\beta,g}+\hat{s}_{\beta,g})\log\frac{eq}{s_{\beta,g}}\right\}+r.$ (S4)
We also have the following concentration inequality on $r$ for $t>0$, which comes from the definition of $r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})$ along with (S3):
$\operatorname{\operatorname{\mathbb{P}}}(r\geq t\sigma^{2}_{\varepsilon})$ $\leq\sum_{s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}}\operatorname{\operatorname{\mathbb{P}}}(r(s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g})\geq t\sigma^{2}_{\varepsilon})\leq\sum_{s^{\prime}_{\gamma},s^{\prime}_{\beta},s^{\prime}_{\beta,g}}e^{-t/3}\leq q^{3}p\,c_{1}e^{-t/3},$
where $c_{1}>0$ is a constant and the sum is taken over $s^{\prime}_{\gamma}\in[q],\,s^{\prime}_{\beta}\in[(p-1)(q+1)],\,s^{\prime}_{\beta,g}\in[q]$. Then for a sufficiently large constant $\tilde{M}$ and some $c_{2}>0$, by letting
$t=\tilde{M}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))$
we have
$\begin{split}&\operatorname{\operatorname{\mathbb{P}}}\left\{r\geq\tilde{M}\sigma^{2}_{\varepsilon}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}\\ &\qquad\leq c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}.\end{split}$ (S5)
We now use the KKT optimality conditions to bound $\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\rVert_{2}^{2}$ in terms of $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$. This takes care of the stochastic term after plugging back into (S1) and (S2). To ease the notation when describing the conditions, define the vectors
[…]
$(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}$ $=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell})$ for $(\hat{\boldsymbol{\mathbf{\beta}}}_{0})_{\ell}\neq 0$ (S7)
$(\boldsymbol{\mathbf{r}}_{\beta})_{\ell}$ $=\lambda\,\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell})+\lambda_{g}\,\frac{(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}}{\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{h}\rVert_{2}}$ $\text{for $(\hat{\boldsymbol{\mathbf{\beta}}}_{h})_{\ell}\neq 0,\,h\in[q]$}.$ (S8)
Squaring both sides of (S6) and summing over $\ell$ gives
$\lambda^{2}\hat{s}_{\gamma}=\frac{1}{n^{2}}\lVert\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\rVert_{2}^{2}$
and doing the same for (S7) and (S8) gives
$\lambda^{2}\hat{s}_{\beta}+\lambda_{g}^{2}\hat{s}_{\beta,g}\leq\frac{1}{n^{2}}\lVert\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\rVert_{2}^{2}$
since in (S8) the cross term $\operatorname{sign}((\hat{\boldsymbol{\mathbf{\beta}}}_{j,h})_{\ell})\times(\hat{\boldsymbol{\mathbf{\beta}}}_{j,h})_{\ell}$ is nonnegative. We have shown that
$\begin{split}\lambda^{2}\hat{s}_{\gamma}+\lambda^{2}\hat{s}_{\beta}+\lambda_{g}^{2}\hat{s}_{\beta,g}&\leq\frac{1}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})\right\rVert_{2}^{2}\\ &=\frac{1}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}(\boldsymbol{\mathbf{\varepsilon}}-\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}-\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}})\right\rVert_{2}^{2}.\end{split}$ (S9)
Using that $\lVert\boldsymbol{\mathbf{a}}-\boldsymbol{\mathbf{b}}\rVert_{2}^{2}\leq 2\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}+2\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}$ for any vectors $\boldsymbol{\mathbf{a}}$ and $\boldsymbol{\mathbf{b}}$ along with Lemma 7, we have with probability at least $1-C_{0}\exp(-\log(pq))$ that
$\begin{split}\lambda^{2}\hat{s}_{\gamma}+\lambda^{2}\hat{s}_{\beta}+\lambda_{g}^{2}\hat{s}_{\beta,g}&\leq\frac{2}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{2}^{2}+\frac{2}{n^{2}}\left\lVert[\boldsymbol{\mathbf{U}}_{\hat{S}_{\gamma}},\boldsymbol{\mathbf{W}}_{\hat{S}_{\beta}}]^{\top}\left(\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}+\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}}\right)\right\rVert_{2}^{2}\\ &\leq\frac{2}{n}M_{uw}\left\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\right\rVert_{2}^{2}+\frac{2}{n}M_{uw}\left\lVert\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}+\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}}\right\rVert_{2}^{2},\end{split}$ (S10)
where the last inequality uses that $\hat{S}_{\gamma}\subset\mathcal{I}_{0}$ and $\hat{S}_{\beta}\subset\mathcal{J}_{0}$.
Now let
$\lambda$ $=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\gamma}+s_{\beta}}+\frac{s_{\beta,g}}{s_{\gamma}+s_{\beta}}\log(eq/s_{\beta,g})\right)^{1/2},$
$\lambda_{g}$ $=C\frac{\sigma_{\varepsilon}}{\sqrt{n}}\left(\frac{s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)}{s_{\beta,g}}+\log(eq/s_{\beta,g})\right)^{1/2},$
where $C=\sqrt{5M_{uw}a_{2}}$ for some $a_{2}>2$. Combining (S10) and (S4) gives
$\begin{split}&\left(1-\frac{2}{a_{2}}\right)\left\lVert\mathcal{P}_{\mathcal{I}_{0},\mathcal{J}_{0}}(\boldsymbol{\mathbf{\varepsilon}})\right\rVert_{2}^{2}\\ &\qquad\leq 5\sigma^{2}_{\varepsilon}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)+\frac{2}{a_{2}}\left\lVert\boldsymbol{\mathbf{U}}{\boldsymbol{\mathbf{\nu}}}+\boldsymbol{\mathbf{W}}{\boldsymbol{\mathbf{\Delta}}}\right\rVert_{2}^{2}+r\end{split}$ (S11)
It is then straightforward to multiply both sides by $a_{1}a_{2}/(2(a_{2}-2))$ and plug into (S1) and (S2) to get
$\begin{split}&\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S^{c}_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S^{c}_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S^{c}_{\beta,g}}\rVert_{1,2}\\ &\qquad\leq\frac{1}{2a_{1}}\cdot\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{5a_{1}a_{2}}{2(a_{2}-2)}\cdot E_{j}+\frac{a_{1}}{a_{2}-2}\cdot\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\\ &\qquad\quad+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}+\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}\end{split}$ (S12)
where $E_{j}=(\sigma^{2}_{\varepsilon}/n)\cdot\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$.
Step 3
We proceed to bound the penalty term by the quantity $\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert$. Standard $\ell_{1},\ell_{2}$ norm inequalities along with Assumption 3 lead to
$\frac{\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}+\frac{\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}}{\sqrt{s_{\beta,g}}}$ $\leq\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{2}$
$\leq 2\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}\leq\frac{2}{\sqrt{\phi_{0}^{\ast}}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
With our specified values of $\lambda$ and $\lambda_{g}$, we have after multiplying through by $C\sqrt{E_{j}}$ that
$\lambda\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta,g}}\rVert_{1,2}$ $\leq\frac{2C}{\sqrt{\phi_{0}^{\ast}}}\sqrt{E_{j}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}$
$\leq a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$
for any $a_{3}>0$. Plugging this into (LABEL:eqn:s11) gives
$\begin{split}&\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\\ &\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}\right)E_{j}+\frac{1}{a_{3}}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}.\end{split}$ (S13)
We now bound the difference between $\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}/n$ and $\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$. Let $L>0$ be an arbitrarily large constant. By Assumption 4 and Lemma 5, we have with probability at least $1-C^{\prime}\exp(-\log(pq))$ that
$\sup_{\lVert\boldsymbol{\mathbf{v}}\rVert_{0}\leq 2(s_{\gamma}+s_{\beta}),\,\lVert\boldsymbol{\mathbf{v}}\rVert_{2}=1}\left\lvert\boldsymbol{\mathbf{v}}^{\top}\left(\frac{[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]}{n}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\boldsymbol{\mathbf{v}}\right\rvert\leq\frac{1}{27L}$
for sufficiently large $n$. By Lemma 6, it holds with probability at least $1-C^{\prime}\exp(-\log(pq))$ that
$\begin{split}&\left\lvert(\boldsymbol{\mathbf{\nu}}^{\top},\,\boldsymbol{\mathbf{\Delta}}^{\top})\left(\frac{[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]}{n}-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\right)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rvert\\ &\qquad\leq\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right)\end{split}$ (S14)
for sufficiently large $n$. Plugging this into (S13) gives
$\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}-\frac{1}{a_{3}}\right)\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
$\qquad+\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\left(\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}-\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\right)$
$\quad\leq\left(\frac{5a_{1}a_{2}}{2(a_{2}-2)}+a_{3}\frac{C^{2}}{\phi_{0}^{\ast}}\right)E_{j}+\frac{a_{1}a_{2}}{2(a_{2}-2)}\cdot\frac{r}{n}$
$\qquad+\left(1-\frac{1}{2a_{1}}-\frac{a_{1}}{a_{2}-2}\right)\cdot\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right).$
Recalling that the above holds for any $a_{1}>0$, $a_{2}>2$, and $a_{3}>0$, choose $a_{1}=2$, $a_{2}=6$, and $a_{3}=6$ to get
$\frac{1}{2}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right)+\frac{r}{n}.$
By the concentration bound on $r$ in (S5), we in fact have
$\frac{1}{2}\left\lVert\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{1/2}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right)$ (S15)
with probability at least $1-c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$. Applying Assumption 3 to the left-hand side, we have with the same high probability
$\frac{\phi_{0}^{\ast}}{2}(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2})\precsim E_{j}+\frac{1}{L}\left(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}+\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\right).$ (S16)
Next, taking $a_{1}=2-\sqrt{2}$, $a_{2}=6$ in (LABEL:eqn:s11) cancels out the $\lVert\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}/n$ terms. Appealing again to (S5) yields
$\frac{\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}+\frac{\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g}^{c})}\rVert_{1,2}}{\sqrt{s_{\beta,g}}}\leq\sqrt{E_{j}}+\frac{\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}+\frac{\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}}{\sqrt{s_{\beta,g}}}.$
with probability at least $1-c_{1}\exp\left\{-c_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$.
Adding $(\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1})/\sqrt{s_{\gamma}+s_{\beta}}$ to both sides and using
$\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}\leq\sqrt{s_{\gamma}+s_{\beta}}(\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2})\quad\text{and}\quad\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}\leq\sqrt{s_{\beta,g}}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}$
yields
$\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}}{\sqrt{s_{\gamma}+s_{\beta}}}\leq\sqrt{E_{j}}+2\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}+3\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}$
and after squaring both sides we have
$\frac{\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}^{2}}{s_{\gamma}+s_{\beta}}\leq k_{0}E_{j}+k_{1}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+k_{2}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$ (S17)
for some absolute constants $k_{0},k_{1},k_{2}>0$. Plugging (S17) into the right-hand side of (S16) yields
$\frac{\phi_{0}^{\ast}}{2}\left(\left\lVert\boldsymbol{\mathbf{\nu}}\right\rVert_{2}^{2}+\left\lVert\boldsymbol{\mathbf{\Delta}}\right\rVert_{2}^{2}\right)\precsim\left(1+\frac{k_{0}}{L}\right)E_{j}+\frac{k_{1}+1}{L}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\frac{k_{2}+1}{L}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}$
Finally, plugging in the expression for $E_{j}$ and recalling that $L$ is arbitrarily large yields
$\lVert\boldsymbol{\mathbf{\nu}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{2}^{2}\precsim\frac{\sigma^{2}_{\varepsilon}}{n}\left(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g})\right)$
with probability at least $1-C_{1}\exp\left\{-C_{2}(s_{\gamma}\log(eq/s_{\gamma})+s_{\beta}\log(ep)+s_{\beta,g}\log(eq/s_{\beta,g}))\right\}$ for some positive constants $C_{1},C_{2}$ as desired. $\square$
S4.4 Proof of Theorem 2
The proof of Theorem 2 occurs in three steps.
Step 1
Recall that $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]/n$. We wish to show that with high probability
$\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\leq\frac{3}{2}(\lambda+\lambda_{g}).$ (S18)
To ease the notation for this step, define the vectors
[…]
so we have for all $\ell\in[p(q+1)-1]$
$\left\lvert\left([\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}(\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}})/n\right)_{\ell}\right\rvert\leq\lambda+\lambda_{g}.$
Recalling that
$\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}=\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\gamma}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\beta}}=\boldsymbol{\mathbf{\varepsilon}}-\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}-\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},$
it follows by the triangle inequality that
$\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}-\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}$ $\leq\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}-\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}$
[…]
$\leq\lambda+\lambda_{g}.$
Thus to show (S18) it suffices to show that with high probability
$\left\lVert\frac{1}{n}[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\leq\frac{\lambda+\lambda_{g}}{2}.$ (S19)
To see this, recall that each element of $\bm{U}$ is bounded by a constant $M_{u}$ by Assumption 1. Then the products of random variables $U_{\ell}$ and $X_{\ell}$ with the random noise $\varepsilon$ are sub-exponential with norm satisfying
$\max_{\ell}\lVert U_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq M_{u}\sigma_{\varepsilon}$
$\max_{\ell}\lVert X_{\ell}\,\varepsilon\rVert_{\psi_{1}}$ $\leq\max_{\ell}\lVert U_{\ell}\rVert_{\psi_{2}}\max_{\ell}\lVert X_{\ell}\rVert_{\psi_{2}}\lVert\varepsilon\rVert_{\psi_{2}}\leq M_{u}K_{X}\sigma_{\varepsilon}$
where $K_{X}$ is the bound on the sub-Gaussian norm of elements of $\bm{X}$ from Assumption 2. Define the vector $\boldsymbol{\mathbf{v}}=[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}/n$, where each element $v_{i}$ is a sum of centered sub-exponential random variables with bounded norm. By Lemma 4 we have
$\operatorname{\operatorname{\mathbb{P}}}\left\{\lvert v_{i}\rvert>\frac{1}{2}(\lambda+\lambda_{g})\right\}\leq 2\exp\left(-c\min\left(\frac{n(\lambda+\lambda_{g})^{2}}{4M_{u}^{2}K_{X}^{2}\sigma_{\varepsilon}^{2}},\frac{n(\lambda+\lambda_{g})}{2M_{u}K_{X}\sigma_{\varepsilon}}\right)\right).$
where $c>0$ is a constant coming from Lemma 4. Since by (18)
$\lambda+\lambda_{g}\geq C\sigma_{\varepsilon}\sqrt{\frac{\log p}{n}}\iff\frac{(\lambda+\lambda_{g})^{2}n}{C^{2}\sigma_{\varepsilon}^{2}}\geq\log p,$
we have
$\frac{(\lambda+\lambda_{g})^{2}n}{C^{2}\sigma_{\varepsilon}^{2}}\cdot C_{0}^{\prime\prime}\geq C_{0}^{\prime\prime}\log p$
where $C^{\prime\prime}_{0}=C^{2}/(4M_{u}^{2}K_{X}^{2})$. Similarly, the inequality
$\frac{(\lambda+\lambda_{g})n}{C\sigma_{\varepsilon}}\geq\sqrt{n\log p}$
implies
$\frac{(\lambda+\lambda_{g})n}{C\sigma_{\varepsilon}}\cdot\sqrt{C_{0}^{\prime\prime}}\geq\sqrt{C_{0}^{\prime\prime}}\cdot\sqrt{n\log p}\geq\tilde{A}^{-1/2}\sqrt{C_{0}^{\prime\prime}}\log p$
provided that $\log p\leq\tilde{A}n$ for some constant $\tilde{A}>0$, which we assumed in the hypothesis of Theorem 2. The above shows that we can pick either argument of the minimum in Lemma 4 to develop and hence
$\operatorname{\operatorname{\mathbb{P}}}\left\{\lvert v_{i}\rvert>\frac{1}{2}(\lambda+\lambda_{g})\right\}\leq 2\exp\left(-C_{0}^{\prime}\log p\right).$
for some constant $C^{\prime}_{0}$ that increases with $C$, recalling that $C$ comes from (18). Then by the union bound we have
$\operatorname{\operatorname{\mathbb{P}}}\left(\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\geq\frac{\lambda+\lambda_{g}}{2}\right)$ $=\operatorname{\operatorname{\mathbb{P}}}\left(\max_{i}\lvert v_{i}\rvert\geq\frac{\lambda+\lambda_{g}}{2}\right)$
$\leq(p(q+1))\cdot 2\exp\left(-C^{\prime}_{0}\log p\right)$
$\leq 2\exp\left(-C^{\prime}_{0}\log p+\log p+\log(q+1)\right)$
$\leq 2\exp(-c_{3}\log p)$
for some $c_{3}>0$. In the last step we used $\log p\asymp\log q$. This shows (S19) holds with high probability.
Next we will show the cone condition
$\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}\leq 4\tau_{j}(\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1})$ (S20)
[…]
$\frac{1}{2n}\lVert\boldsymbol{\mathbf{x}}_{j}-\boldsymbol{\mathbf{U}}\hat{\boldsymbol{\mathbf{\gamma}}}-\boldsymbol{\mathbf{W}}\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{2}^{2}+\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq\frac{1}{2n}\lVert\boldsymbol{\mathbf{\varepsilon}}_{j}\rVert_{2}^{2}+\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}.$
Rearranging the above leads to
$\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}+\frac{1}{n}\langle\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},\boldsymbol{\mathbf{\varepsilon}}\rangle.$ (S21)
By the same argument leading to (S19), the events
$\mathcal{A}_{1}=\left\{\frac{1}{n}\left\lVert\boldsymbol{\mathbf{U}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\leq\frac{\lambda}{2}\right\}\quad\text{and}\quad\mathcal{A}_{2}=\left\{\frac{1}{n}\left\lVert\boldsymbol{\mathbf{W}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\right\rVert_{\infty}\leq\frac{\lambda}{2}\right\}$
hold with probability at least $1-2\exp(-c_{4}\log p)$ and $1-2\exp(-c_{5}\log p)$ respectively for some constants $c_{4}>0$ and $c_{5}>0$.
Conditional on $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, we have by Hölder’s inequality
$\frac{1}{n}\langle\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}}+\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},\boldsymbol{\mathbf{\varepsilon}}\rangle$ $=\frac{1}{n}\langle\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{\nu}},\boldsymbol{\mathbf{\varepsilon}}\rangle+\frac{1}{n}\langle\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{\Delta}},\boldsymbol{\mathbf{\varepsilon}}\rangle$
$\leq\frac{1}{n}\lVert\boldsymbol{\mathbf{U}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\rVert_{\infty}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\frac{1}{n}\lVert\boldsymbol{\mathbf{W}}^{\top}\boldsymbol{\mathbf{\varepsilon}}\rVert_{\infty}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}$
$\leq\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}$
so that (S21) becomes
$\lambda\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lambda\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lambda_{g}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq\lambda\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+\lambda\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+\lambda_{g}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}+\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\frac{\lambda}{2}\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}.$
Since $\lambda_{g}/\lambda=\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$, we have
$2\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+2\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+2\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}\leq 2\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}+2\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}+2\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}+\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}.$
Now subtract by the left-hand side and add $\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}$, $\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}$, and $\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}\lVert\boldsymbol{\mathbf{\Delta}}_{-0}\rVert_{1,2}$ to find
$\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}+\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\boldsymbol{\mathbf{\Delta}}_{-0}\rVert_{1,2}$ $\leq 2\left(\lVert\boldsymbol{\mathbf{\gamma}}\rVert_{1}-\lVert\hat{\boldsymbol{\mathbf{\gamma}}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}\right)$
$\quad+2\left(\lVert\boldsymbol{\mathbf{\beta}}\rVert_{1}-\lVert\hat{\boldsymbol{\mathbf{\beta}}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}\right)$
$\quad+2\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\left(\lVert\boldsymbol{\mathbf{\beta}}_{-0}\rVert_{1,2}-\lVert\hat{\boldsymbol{\mathbf{\beta}}}_{-0}\rVert_{1,2}+\lVert\boldsymbol{\mathbf{\Delta}}_{-0}\rVert_{1,2}\right),$
Then using the triangle inequality and the fact that the terms in parentheses vanish outside of the respective support sets $S_{\gamma},S_{\beta},$ and $S_{\beta,g}$, we have
$\lVert\boldsymbol{\mathbf{\nu}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}\rVert_{1}\leq 4\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+4\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}+4\sqrt{\frac{s_{\gamma}+s_{\beta}}{s_{\beta,g}}}\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}.$
Since $\lVert\boldsymbol{\mathbf{\Delta}}_{(S_{\beta,g})}\rVert_{1,2}\leq\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1}$, the above yields
$\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}^{c}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}^{c}}\rVert_{1}\leq 4\tau_{j}(\lVert\boldsymbol{\mathbf{\nu}}_{S_{\gamma}}\rVert_{1}+\lVert\boldsymbol{\mathbf{\Delta}}_{S_{\beta}}\rVert_{1})$
as desired, recalling again that $\tau_{j}=1+\sqrt{(s_{\gamma}+s_{\beta})/s_{\beta,g}}$ as defined in Assumption 5
Step 2
In this step, we bound the diagonal difference $\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert$. Notice that $\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ is a sum of sub-exponential random variables with sub-exponential norm bounded by $M_{u}^{2}K_{X}^{2}$, where $M_{u}$ is from Assumption 1 and $K_{X}$ from Assumption 2. We have by Lemma 4
$\operatorname{\operatorname{\mathbb{P}}}\left(\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert>\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\right)$ $\leq 2\exp\left(-cn\min\left(\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)^{2}}{4M_{u}^{2}K_{X}^{2}},\frac{\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)}{2M_{u}K_{X}}\right)\right)$
$\leq 2\exp\left(-cn\min\left(\frac{(\phi^{\ast}_{0})^{2}}{4M_{u}^{2}K_{X}^{2}},\frac{\phi^{\ast}_{0}}{2M_{u}K_{X}}\right)\right)$
$\leq 2\exp(-c_{6}n)$
where we used $\phi^{\ast}_{0}\leq\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$ by Assumption 3. Since
$\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\rvert\leq\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$
implies
$\frac{1}{2}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)$
which by Assumption 3 implies
$\frac{\phi^{\ast}_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\phi^{\ast}_{1},$
it follows from the above concentration bound that
$\operatorname{\operatorname{\mathbb{P}}}\left(\frac{\phi^{\ast}_{0}}{2}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\phi^{\ast}_{1}\right)\geq 1-2\exp(-c_{6}n).$ (S22)
Next we bound the off-diagonal terms. Again by Lemmas 4 and Assumption 3,
$\operatorname{\operatorname{\mathbb{P}}}\left(\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)\rvert\geq 2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)\right)$
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)^{2}}{M^{2}K_{X}^{2}},\frac{2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,k)}{MK_{X}}\right)\right)$
$\qquad\leq 2\exp\left(-cn\min\left(\frac{4(\phi^{\ast}_{1})^{2}}{M^{2}K_{X}^{2}},\frac{4\phi^{\ast}_{1}}{MK_{X}}\right)\right)$
$\qquad\leq 2\exp(-c_{7}n).$
By Assumption 5, we have
$-2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq 2\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$
$\implies-\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq 3\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)$
$\implies-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}$
Combining this with the above concentration bound yields
$\operatorname{\operatorname{\mathbb{P}}}\left(-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\right)\geq 1-2\exp(-c_{7}n).$ (S23)
Step 3
Define the index set $\tilde{S}=S_{\gamma}\cup\{q+i\mid i\in S_{\beta}\}$. The set $\{q+i\mid i\in S_{\beta}\}$ indexes the columns of the submatrix $\boldsymbol{\mathbf{W}}$ of the matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ corresponding to the support set of $\boldsymbol{\mathbf{\beta}}$. With slight abuse of notation, for a vector $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}$ let $\boldsymbol{\mathbf{v}}_{\tilde{S}}\in\mathbb{R}^{p(q+1)-1}$ be the vector that equals $\boldsymbol{\mathbf{v}}$ on the set $\tilde{S}$ and is zero on $\tilde{S}^{c}$, so that it has $s_{\gamma}+s_{\beta}$ nonzero elements. Define $\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}$ analogously.
We will show that conditional on $\mathcal{A}_{1}$, $\mathcal{A}_{2}$, and the bounds
$\begin{split}\frac{\phi^{\ast}_{0}}{2}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\leq 2\phi^{\ast}_{1},\\ -\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}&\leq\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})},\end{split}$ (S24)
we have
$\inf_{\boldsymbol{\mathbf{v}}\in\mathcal{V}}\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}\rVert_{2}}{\sqrt{n}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}}\geq\sqrt{\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}}}>0,$
where $\mathcal{V}=\{\boldsymbol{\mathbf{v}}\in\mathbb{R}^{p(q+1)-1}\mid\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}\leq 4\tau_{j}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\}$ and $c_{0}>2/\phi^{\ast}_{0}$ in Assumption 5.
First we have that
$\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$ $=\frac{\boldsymbol{\mathbf{v}}_{\tilde{S}}^{\top}\operatorname{diag}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}})\boldsymbol{\mathbf{v}}_{\tilde{S}}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}+\frac{\boldsymbol{\mathbf{v}}_{\tilde{S}}^{\top}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}-\operatorname{diag}(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}))\boldsymbol{\mathbf{v}}_{\tilde{S}}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}.$
Then by $\lVert\boldsymbol{\mathbf{a}}+\boldsymbol{\mathbf{b}}\rVert_{2}^{2}=\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}+2\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{b}}+\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}$ we have
$\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}\rVert_{2}^{2}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$ $\geq\frac{\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}+2\frac{\boldsymbol{\mathbf{v}}_{\tilde{S}}^{\top}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}}{n\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{2}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}^{c}}\rVert_{1}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}-\frac{8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{1+8\tau_{j}}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}\frac{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{1}^{2}}{\lVert\boldsymbol{\mathbf{v}}_{\tilde{S}}\rVert_{2}^{2}}$
$\geq\frac{\phi^{\ast}_{0}}{2}-\frac{(1+8\tau_{j})(s_{\beta}+s_{\gamma})}{c_{0}(1+8\tau_{j})(s_{\beta}+s_{\gamma})}=\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}}>0.$
We have shown that
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}\geq\left(\frac{\phi^{\ast}_{0}}{2}-\frac{1}{c_{0}}\right)\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}$ (S25)
Final step
It is true that for $\ell\in[p(q+1)-1]$ we have
$\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}=\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}+\sum_{k\neq\ell}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{k}.$
Then by (S24) and the triangle inequality we have
$\left\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}\right\rvert-\left\lvert\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert$ $\leq\left\lvert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(\ell,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}-\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert$
$\leq\left\lvert\sum_{k\neq\ell}\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}(k,\ell)\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{k}\right\rvert$
$\leq\frac{3}{c_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\sum_{k\neq\ell}\left\lvert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{k}\right\rvert$.
Rearranging terms and applying (S24) yields
$\frac{\phi^{\ast}_{0}}{2}\left\lvert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\ell}\right\rvert\leq\left\lvert\left(\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right)_{\ell}\right\rvert+\frac{3}{c_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$
Since this holds for all $\ell\in[p(q+1)-1]$, we have shown that
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\leq\frac{2}{\phi^{\ast}_{0}}\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}+\frac{6}{c_{0}\phi^{\ast}_{0}(1+8\tau_{j})(s_{\gamma}+s_{\beta})}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}.$ (S26)
Combining (S18) and (S20) we have
$\frac{1}{n}\left\lVert[\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{2}^{2}$ $\leq\left\lVert\hat{\boldsymbol{\mathbf{\Sigma}}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}$
$\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{1}$
$\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\sqrt{s_{\gamma}+s_{\beta}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
and combining this with (S25) gives us
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}^{2}\leq\frac{3}{2}(\lambda+\lambda_{g})\cdot(1+4\tau_{j})\left(\frac{2c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
and therefore
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})\left(\frac{c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)\sqrt{s_{\gamma}+s_{\beta}}.$
On the other hand, by (S20) we have
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}\leq(1+4\tau_{j})\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{1}\leq(1+4\tau_{j})\sqrt{s_{\beta}+s_{\gamma}}\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}_{\tilde{S}}\right\rVert_{2}$
so combining this with the above yields
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{1}\leq 3(\lambda+\lambda_{g})(1+4\tau_{j})^{2}\left(\frac{c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)(s_{\gamma}+s_{\beta})$
Finally, plugging this into (S26) and with (S18) yields
$\left\lVert\begin{pmatrix}\boldsymbol{\mathbf{\nu}}\\ \boldsymbol{\mathbf{\Delta}}\end{pmatrix}\right\rVert_{\infty}$ $\leq\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})+\frac{18(\lambda+\lambda_{g})(1+4\tau_{j})^{2}}{c_{0}\phi^{\ast}_{0}(1+8\tau_{j})}\left(\frac{c_{0}}{c_{0}\phi^{\ast}_{0}-2}\right)$
$=\frac{3}{\phi^{\ast}_{0}}(\lambda+\lambda_{g})\left(1+\frac{6(1+4\tau_{j})^{2}}{(1+8\tau_{j})(c_{0}\phi^{\ast}_{0}-2)}\right)$
as desired. $\square$
S5 Sufficient conditions for theoretical results
While Assumptions 1-4 are enough to prove Theorem 1 and the addition of Assumption 5 for Theorem 2, it is desirable to present a set of interpretable sufficient conditions. The key difference in the theoretical analysis of our proposed method versus that of zhangli is that, due to the stage 1 centering in (20), the columns of the design matrix in zhangli are centered. Indeed a typical column of their design matrix $\boldsymbol{\mathbf{W}}_{\mathrm{ZL}}$ consists of the interaction term $z^{(i)}_{j}u^{(i)}_{h}$ and hence $\operatorname{\operatorname{\mathbb{E}}}(z^{(i)}_{j}u^{(i)}_{h})=\operatorname{\operatorname{\mathbb{E}}}(u^{(i)}_{h}\operatorname{\operatorname{\mathbb{E}}}(z^{(i)}_{j}\>|\>\boldsymbol{\mathbf{u}}^{(i)}))=0$. The trade-off of this simplification is that in the non-oracle case, the difference between $\boldsymbol{\mathbf{W}}_{\mathrm{ZL}}$ and the noisy design matrix $\hat{\boldsymbol{\mathbf{W}}}_{\mathrm{ZL}}$ consisting of the estimated terms $\hat{z}^{(i)}_{j}u^{(i)}_{h}$ must be bounded separately. By contrast, the design matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$ of our proposed method constains the terms $x^{(i)}_{j}u^{(i)}_{h}$ where it does not necessarily hold that $\operatorname{\operatorname{\mathbb{E}}}(x^{(i)}_{j}\>|\>\boldsymbol{\mathbf{u}}^{(i)})=0$, but does obviate the need to bound a noisy design matrix.
One consequence of the above is that the matrix $\operatorname{\operatorname{\mathbb{E}}}(\boldsymbol{\mathbf{W}}_{\mathrm{ZL}}^{\top}\boldsymbol{\mathbf{W}}_{\mathrm{ZL}}/n)$ is a covariance matrix whose spectrum is easily bounded via bounds on the eigenvalues of $\operatorname{Cov}(\bm{U})$ and $\operatorname{Cov}(\bm{X}\>|\>\bm{U})$; these are Assumptions 1 and 2 in zhangli. Meanwhile, our analogous matrix $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$ is an (uncentered) Gram matrix. Nevertheless, Theorems 1 and 2 follows from interpretable bounds on the eigenvalues of $\operatorname{Cov}(\bm{U})$ and $\operatorname{Cov}(\bm{X}\>|\>\bm{U})$, plus additional assumptions that control the mean term.
S5.1 Sufficient conditions for Theorem 1
Assumption 6.
The covariates $\{\boldsymbol{\mathbf{u}}^{(i)}\}_{i=1}^{n}$ are i.i.d. mean zero random vectors with covariance matrix satisfying
$\phi_{0}\leq\lambda_{\mathrm{min}}(\operatorname{Cov}(\boldsymbol{\mathbf{u}}^{(i)}))\leq\lambda_{\mathrm{max}}(\operatorname{Cov}(\boldsymbol{\mathbf{u}}^{(i)}))\leq\phi_{1}$
for some constants $0<\phi_{0}\leq\phi_{1}<\infty$.
Assumption 7.
Recall that $\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}})=\operatorname{Cov}(\bm{X}\>|\>\bm{U}=\boldsymbol{\mathbf{u}})$. Suppose for constants $\psi_{0},\psi_{1}>0$ we have
$\psi_{0}\leq\lambda_{\mathrm{min}}(\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}}^{(i)}))\leq\lambda_{\mathrm{max}}(\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}}^{(i)}))\leq\psi_{1}.$
Assumptions 6 and 7 are analogous to Assumptions 1 and 2 in zhangli respectively, and are standard for results in the high-dimensional setting.
Next, as our design matrix is the concatenated $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$, we must restrict the cross-covariance between $\boldsymbol{\mathbf{U}}$ and $\boldsymbol{\mathbf{W}}$.
Assumption 8.
For any vectors $\boldsymbol{\mathbf{a}}\in\mathbb{R}^{q}$, $\boldsymbol{\mathbf{b}}\in\mathbb{R}^{(p-1)(q+1)}$ define the correlation of projections
$\rho^{(i)}(\boldsymbol{\mathbf{a}},\boldsymbol{\mathbf{b}})=\operatorname{Corr}\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{u}}^{(i)},\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{w}}^{(i)}\right)$
and let $\rho^{(i)}_{\min}=\min_{\lVert\boldsymbol{\mathbf{a}}\rVert_{2}=1,\lVert\boldsymbol{\mathbf{b}}\rVert_{2}=1}\rho^{(i)}(\boldsymbol{\mathbf{a}},\boldsymbol{\mathbf{b}})$ denote the minimal canonical correlation between $\boldsymbol{\mathbf{u}}^{(i)}$ and $\boldsymbol{\mathbf{w}}^{(i)}$. We assume there exists a constant $\rho_{0}$ satisfying $0\leq\rho_{0}<1$ such that $\rho^{(i)}_{\min}>-\rho_{0}$ for all $i\in[n]$. Equivalently, all canonical correlations between $\boldsymbol{\mathbf{u}}^{(i)}$ and $\boldsymbol{\mathbf{w}}^{(i)}$ are uniformly bounded below by $\rho_{0}$.
Assumptions 9 and 10 bound the mean term and control the uncentered nature of our design matrix $\boldsymbol{\mathbf{U}},\boldsymbol{\mathbf{W}}]$. In the context of eQTL studies, Assumption 9 and 10 are met under routine normalization and variance–stabilization steps in standard data-processing pipelines.
Assumption 9.
There exist constants $M_{\mu},M_{u,1}>0$ such that $\operatorname{\operatorname{\mathbb{E}}}\lVert\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})\rVert_{2}^{4}\leq M_{\mu}$ and $\operatorname{\operatorname{\mathbb{E}}}\lVert\boldsymbol{\mathbf{u}}\rVert_{2}^{4}\leq M_{u,1}$.
Assumption 10.
There exists a constant $M_{\infty}>0$ such that $\lVert\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})\rVert_{\infty}\leq M_{\infty}$.
Together, Assumptions 6 through 10 are sufficient to imply Assumptions 3 and 2. This is summarized in the following propositions, which are proved in Appendix S5.2.
Proposition 1.
Assumptions 1 and 6-9 imply Assumption 3.
Proposition 2.
Assumptions 1, 7, and 10 imply Assumption 2.
S5.2 Sufficient conditions for Theorem 2
While Assumption 5 controls the mutual coherence of the Gram matrix $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}$, it is desirable to relate this bound to the covariance of responses $\bm{X}$ and covariates $\bm{U}$. In fact with Assumption 10 we have
[…]
Therefore, Assumption 5 is satisfied when the covariances of $\bm{X}$ and $\bm{U}$ are small and the mean of $\bm{X}$ is small; for instance it holds when $\operatorname{Cov}(\bm{U})=\boldsymbol{\mathbf{I}}$.
S5.3 Proof of Proposition 1
We wish to prove that Assumptions 1 and 6-9 imply Assumption 3. It suffices to upper and lower bound $\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ for all unit vectors $\boldsymbol{\mathbf{v}}$. Express the block matrix form as
$\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}=\begin{pmatrix}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}&\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{(12)}\\ \boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{(12)\top}&\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\end{pmatrix}$
and let $\boldsymbol{\mathbf{w}}=(\boldsymbol{\mathbf{x}},u_{1}\boldsymbol{\mathbf{x}},\dotsc,u_{q}\boldsymbol{\mathbf{x}})\in\mathbb{R}^{(p-1)(q+1)}$ be a row of the data matrix $\boldsymbol{\mathbf{W}}$. Express an arbitrary unit vector $\boldsymbol{\mathbf{v}}$ as $(\boldsymbol{\mathbf{a}},\boldsymbol{\mathbf{b}})$ where $\boldsymbol{\mathbf{a}}\in\mathbb{R}^{q}$ and $\boldsymbol{\mathbf{b}}\in\mathbb{R}^{(p-1)(q+1)}$. Further express $\boldsymbol{\mathbf{b}}$ as $(\boldsymbol{\mathbf{b}}_{0},\boldsymbol{\mathbf{b}}_{1},\dotsc,\boldsymbol{\mathbf{b}}_{q})$ where $\boldsymbol{\mathbf{b}}_{h}\in\mathbb{R}^{p-1}$, $h=0,\dotsc,q$. Let $\boldsymbol{\mathbf{B}}=[\boldsymbol{\mathbf{b}}_{1},\dotsc,\boldsymbol{\mathbf{b}}_{q}]\in\mathbb{R}^{(p-1)\times q}$ and define $c(\boldsymbol{\mathbf{u}})=\boldsymbol{\mathbf{b}}_{0}+\boldsymbol{\mathbf{B}}\boldsymbol{\mathbf{u}}$. This allows us to write $\boldsymbol{\mathbf{w}}^{\top}\boldsymbol{\mathbf{b}}=\boldsymbol{\mathbf{x}}^{\top}c(\boldsymbol{\mathbf{u}})$. Recall the notation $\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})=\operatorname{\operatorname{\mathbb{E}}}(\boldsymbol{\mathbf{x}}\>|\>\boldsymbol{\mathbf{u}})$ and $\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}})=\operatorname{Cov}(\boldsymbol{\mathbf{x}}\>|\>\boldsymbol{\mathbf{u}})$.
Lower bound: Since $\boldsymbol{\mathbf{u}}$ is centered we have $\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\geq\phi_{0}\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}$ by Assumption 6. On the other hand, by Assumptions 6 and 7 and the law of total variance we have
$\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}=\operatorname{\operatorname{\mathbb{E}}}(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{x}})^{2}$ $\geq\operatorname{Var}(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{x}})$
$\geq\operatorname{\operatorname{\mathbb{E}}}\left(\operatorname{Var}(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{x}}\>|\>\boldsymbol{\mathbf{u}})\right)$
$=\operatorname{\operatorname{\mathbb{E}}}\left(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}})c(\boldsymbol{\mathbf{u}})\right)$
$\geq\psi_{0}\operatorname{\operatorname{\mathbb{E}}}\lVert c(\boldsymbol{\mathbf{u}})\rVert_{2}^{2}$
$=\psi_{0}\left\{\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{2}+\operatorname{tr}\left(\boldsymbol{\mathbf{B}}\operatorname{Cov}(\boldsymbol{\mathbf{u}})\boldsymbol{\mathbf{B}}^{\top}\right)\right\}$
$\geq\psi_{0}\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{2}+\psi_{0}\phi_{0}\lVert\boldsymbol{\mathbf{B}}\rVert_{F}^{2}\geq\phi_{0}^{\prime}\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}$
where $\phi_{0}^{\prime}=\min(\psi_{0},\psi_{0}\phi_{0})$. Now let
$\rho=\frac{\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{(12)}\boldsymbol{\mathbf{b}}}{\sqrt{\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\right)\left(\boldsymbol{\mathbf{b}}^{\top}\operatorname{Cov}(\boldsymbol{\mathbf{w}})\boldsymbol{\mathbf{b}}\right)}}.$
Since $\boldsymbol{\mathbf{u}}$ is centered, $\rho$ is the correlation between the projections $\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{u}}$ and $\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{w}}$. By Assumption 8, it holds for some $0\leq\rho_{0}<1$ that
$\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{(12)}\boldsymbol{\mathbf{b}}$ $=\rho\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\right)^{1/2}\left(\boldsymbol{\mathbf{b}}^{\top}\operatorname{Cov}(\boldsymbol{\mathbf{w}})\boldsymbol{\mathbf{b}}\right)^{1/2}$
$\geq-\rho_{0}\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\right)^{1/2}\left(\boldsymbol{\mathbf{b}}^{\top}\operatorname{Cov}(\boldsymbol{\mathbf{w}})\boldsymbol{\mathbf{b}}\right)^{1/2}$
$\geq-\rho_{0}\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\right)^{1/2}\left(\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}\right)^{1/2},$
since $\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\succsim\operatorname{Cov}(\boldsymbol{\mathbf{w}})$. All in all we have
$\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ $=\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}+\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}+2\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{(12)}\boldsymbol{\mathbf{b}}$
$\geq\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}+\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}-2\rho_{0}\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\right)^{1/2}\left(\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}\right)^{1/2}.$
By completing the square in both $\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}$ and $\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}$ this yields
$\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ $\geq(1-\rho_{0}^{2})\max\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}},\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}\right)$
$\geq(1-\rho_{0}^{2})\max(\phi_{0}\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2},\phi_{0}^{\prime}\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2})$
$\geq(1-\rho_{0}^{2})\frac{\phi_{0}\phi_{0}^{\prime}}{\phi_{0}+\phi_{0}^{\prime}}=:\phi_{0}^{\ast},$
in the last step recalling that $\lVert\boldsymbol{\mathbf{v}}\rVert_{2}^{2}=\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}=1$.
Upper bound: By Assumption 6 we have $\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\leq\phi_{1}\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}$. We also have
$\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}$ $=\operatorname{\operatorname{\mathbb{E}}}(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{x}})^{2}$
$=\operatorname{\operatorname{\mathbb{E}}}\left(c(\boldsymbol{\mathbf{u}})^{\top}\operatorname{\operatorname{\mathbb{E}}}(\boldsymbol{\mathbf{x}}\boldsymbol{\mathbf{x}}^{\top}\>|\>\boldsymbol{\mathbf{u}})c(\boldsymbol{\mathbf{u}})\right)$
$=\operatorname{\operatorname{\mathbb{E}}}\left(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}})c(\boldsymbol{\mathbf{u}})\right)+\operatorname{\operatorname{\mathbb{E}}}\left(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})^{\top}c(\boldsymbol{\mathbf{u}})\right).$ (S27)
By Assumptions 6 and 7 we upper bound the first term as
$\operatorname{\operatorname{\mathbb{E}}}\left(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}})c(\boldsymbol{\mathbf{u}})\right)$ $\leq\psi_{1}\operatorname{\operatorname{\mathbb{E}}}\lVert c(\boldsymbol{\mathbf{u}})\rVert_{2}^{2}$
$=\psi_{1}\left\{\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{2}+\operatorname{tr}\left(\boldsymbol{\mathbf{B}}\operatorname{Cov}(\boldsymbol{\mathbf{u}})\boldsymbol{\mathbf{B}}^{\top}\right)\right\}$
$\leq\psi_{1}\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{2}+\psi_{1}\phi_{1}\lVert\boldsymbol{\mathbf{B}}\rVert_{F}^{2}\leq\phi_{1}^{\prime}\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2},$ (S28)
where $\phi_{1}^{\prime}=\max(\psi_{1},\psi_{1}\phi_{1})$.
For the second term in (S5.3), applying Cauchy-Schwarz twice yields
$\operatorname{\operatorname{\mathbb{E}}}(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}}))^{2}\leq\operatorname{\operatorname{\mathbb{E}}}\left(\lVert c(\boldsymbol{\mathbf{u}})\rVert_{2}^{2}\lVert\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})\rVert_{2}^{2}\right)\leq\left(\operatorname{\operatorname{\mathbb{E}}}\lVert c(\boldsymbol{\mathbf{u}})\rVert_{2}^{4}\right)^{1/2}\left(\operatorname{\operatorname{\mathbb{E}}}\lVert\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}})\rVert_{2}^{4}\right)^{1/2}.$
Since $\lVert c(\boldsymbol{\mathbf{u}})\rVert_{2}^{4}\leq 8\left(\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{4}+\lVert\boldsymbol{\mathbf{B}}\rVert_{F}^{4}\lVert\boldsymbol{\mathbf{u}}\rVert_{2}^{4}\right)$, by Assumption 9, we have
$(\operatorname{\operatorname{\mathbb{E}}}\lVert c(\boldsymbol{\mathbf{u}})\rVert_{2}^{4})^{1/2}\leq\sqrt{8}\left(\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{B}}\rVert_{F}^{2}(\operatorname{\operatorname{\mathbb{E}}}\lVert\boldsymbol{\mathbf{u}}\rVert_{2}^{4})^{1/2}\right)\leq\sqrt{8}\left(\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{B}}\rVert_{F}^{2}M_{u,1}^{1/2}\right).$ (S29)
Therefore,
$\operatorname{\operatorname{\mathbb{E}}}(c(\boldsymbol{\mathbf{u}})^{\top}\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}}))^{2}\leq\sqrt{8}\left(\lVert\boldsymbol{\mathbf{b}}_{0}\rVert_{2}^{2}+\lVert\boldsymbol{\mathbf{B}}\rVert_{F}^{2}M_{u,1}^{1/2}\right)M_{\mu}^{1/2}\leq\sqrt{8}M_{\mu}^{1/2}(1+M_{u,1}^{1/2})\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}=:\phi_{2}^{\prime}\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}.$ (S30)
by Assumption 9. Plugging (S28) and (S30) into (S5.3) yields
$\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}\leq(\phi_{1}^{\prime}+\phi_{2}^{\prime})\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}.$
For cross term, we have by the AM-GM inequality
$\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{(12)}\boldsymbol{\mathbf{b}}$ $=\rho\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\right)^{1/2}\left(\boldsymbol{\mathbf{b}}^{\top}\operatorname{Cov}(\boldsymbol{\mathbf{w}})\boldsymbol{\mathbf{b}}\right)^{1/2}$
$\leq\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}\right)^{1/2}\left(\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}\right)^{1/2}$
$\leq\frac{1}{2}\left({\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}}+\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}\right).$
All in all we have
$\boldsymbol{\mathbf{v}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{v}}$ $=\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}+\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}+2\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{W}}}^{(12)}\boldsymbol{\mathbf{b}}$
$\leq\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}+\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}+\left(\boldsymbol{\mathbf{a}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{U}}\boldsymbol{\mathbf{U}}}\boldsymbol{\mathbf{a}}+\boldsymbol{\mathbf{b}}^{\top}\boldsymbol{\mathbf{\Sigma}}_{\boldsymbol{\mathbf{W}}\boldsymbol{\mathbf{W}}}\boldsymbol{\mathbf{b}}\right)$
$=2\left(\phi_{1}\lVert\boldsymbol{\mathbf{a}}\rVert_{2}^{2}+(\phi_{1}^{\prime}+\phi_{2}^{\prime})\lVert\boldsymbol{\mathbf{b}}\rVert_{2}^{2}\right)$
$\leq 2\max(\phi_{1},\phi_{1}^{\prime}+\phi_{2}^{\prime})=:\phi^{\ast}_{1}.\qed$
S5.4 Proof of Proposition 2
We show that Assumptions 1, 7, and 10 imply Assumption 2. Recall that $\boldsymbol{\mathbf{X}}\>|\>\boldsymbol{\mathbf{U}}=\boldsymbol{\mathbf{u}}\sim\mathcal{N}(\boldsymbol{\mathbf{\mu}}(\boldsymbol{\mathbf{u}}),\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}}))$. Write
$X_{i}-\operatorname{\operatorname{\mathbb{E}}}X_{i}=(X_{i}-\mu_{i}(\boldsymbol{\mathbf{u}}))+(\mu_{i}(\boldsymbol{\mathbf{u}})-\operatorname{\operatorname{\mathbb{E}}}[\mu_{i}(\boldsymbol{\mathbf{u}})]).$
It suffices to show that both components are sub-Gaussian with bounded sub-Gaussian norm. Conditionally on $\boldsymbol{\mathbf{u}}$ we have
$\operatorname{\operatorname{\mathbb{E}}}(X_{i}-\mu_{i}(\boldsymbol{\mathbf{u}})\>|\>\boldsymbol{\mathbf{U}}=\boldsymbol{\mathbf{u}})=\epsilon_{i}(\boldsymbol{\mathbf{u}})\sim\mathcal{N}(0,[\boldsymbol{\mathbf{\Sigma}}(\boldsymbol{\mathbf{u}})]_{ii}).$
By Assumption 7 we have $\operatorname{Var}(\epsilon_{i}(\boldsymbol{\mathbf{u}}))\leq\psi_{1}$ and hence $\epsilon_{i}(\boldsymbol{\mathbf{u}})$ is sub-Gaussian with $\lVert\epsilon_{i}(\boldsymbol{\mathbf{u}})\rVert_{\psi_{2}}\leq K_{1}\sqrt{\psi_{1}}$ for some absolute constant $K_{1}>0$. Since this holds uniformly in $\boldsymbol{\mathbf{u}}$, it follows that $X_{i}-\mu_{i}(\boldsymbol{\mathbf{u}})$ is (marginally) sub-Gaussian with $\lVert X_{i}-\mu_{i}(\boldsymbol{\mathbf{u}})\rVert_{\psi_{2}}\leq K_{1}\sqrt{\psi_{1}}$. On the other hand, by Assumption 10 we have $\lvert\mu_{i}(\boldsymbol{\mathbf{u}})-\operatorname{\operatorname{\mathbb{E}}}[\mu_{i}(\boldsymbol{\mathbf{u}})]\rvert\leq 2M_{\infty}$. Therefore, $\mu_{i}(\boldsymbol{\mathbf{u}})-\operatorname{\operatorname{\mathbb{E}}}[\mu_{i}(\boldsymbol{\mathbf{u}})]$ is sub-Gaussian with bounded sub-Gaussian norm. ∎